A Complete Bibliography of Publications in Monte Carlo Methods and Applications

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

05 April 2022 Version 1.18

Title word cross-reference

(0, m, 2) [Xia02]. (n, k) [Gol03]. (t, m, 2) [DK06]. (t, m, s) [WLD21]. 1 [BJ01, CA12, NPM^{+}06]. 2 [FM01, NPM^{+}06]. 2^{p} [NS09]. 3 [KS05, Kur95a]. b ≥ 2 [Xia02]. C [PS98]. C([0, T]) [KPV18]. G_{I}^{0} [CRGF18]. Δ^{2} [Mis07]. ϵ [CJV16, GG05]. GF(2) [Tak96b]. k [ASTY19]. L^{2} [Ego97]. L_{p}(T) [KM15]. m [Tak96a, Tak97]. M/M/r [CS96]. M^{X}/G/1 [SC96]. p [FKM08]. ±1 [EM03]. R [Tor20]. t [Nad08a, Nad08b, Sak10]. Θ [Buc04]. Z [ONZ99].

-adic [FKM08]. -copula [Sak10]. -distribution [Mis07]. -isomorphic [Ego97]. -Maruyama [Buc04]. -nets [DK06, WLD21, Xia02]. -optimal [GG05]. -particle [Gol03]. -perfect [CJV16]. -sequences [Tak96a, Tak97]. -space [PS98]. -wise [ASTY19].

1 [BOTAZ19, TOTAI18]. 11 [Hal05a]. 17 [LP13].

2 [Oga97, Tuf98]. 2000 [Ano00g, Ano00h]. 2003 [Ano02e].

3D [RBB21]. 3rd [Ano00a].

97j [Oga97]. 97m [Tuf98].

a.s [FP02]. abrupt [KLR^{+}03]. Absorbing [GZ01]. absorption [Hal21]. Abstracts [Ano00h]. accelerators [ONZ99]. acceptance [Nk16]. accuracies [Gri17]. Accuracy [SSS06, KM15, RL18, Tur19]. across [MR04]. Adaptative [Aro04]. Adaptive [BST10, DM10, Kaw07, mMSD04, BFP09, HvSST14, KS16, LL11, ZZA21]. Additive [DT01]. Adiabatic [DK98b]. adic [FKM08]. Adjoint [BP97, BP98a, BP98b, KRSV99, mMSD04]. adjusted [MDMS20]. Admissible [NS07]. Aerosol [ZPK02, SRKL96, SK00]. aerosols [LT08]. Affected [SS07]. aggregate [Tor20]. aggregation [SZKS21]. Algebra [AAD04, Hal08b]. Algebraic [Lik98, Ant11, ER06, MM12]. Algorithm [Ave04, BG01, DSGZ01, GHT00, HPY07, Sim95, SVH^{+}04, UŠ96, BCR11, BN15, CA12, CRS14, FN09, FGD13, Gri10, Gri17, HvSST14, JML20, KLP14, LL13, MP12, MDMS20, NB19, Raj19, SS03, Sab17, Sab19a, Sab19b, SP20, SS21a, SSDT21, Sha10, SS14a, SS21b, SAKG15]. Algorithms [CL02b, Hei04, HMG01, KRSV99, NP04b, Pap04, SM09, SM04, WK05, BGSR08, EW01, ER06, FP02, KS95a, LOR18, NÖ09b, RV99, SRKL96, Sab16d, Sim18, SK05, Spa21, VMS08]. allowing [BN15]. almost [SD96]. American [BCZ05, BKS06]. among [Nad08a]. Amplitude [AD01]. analog [KS16, Smi98]. Analysis [BBBR19, BAO^{+}04, CS96, KOSY01, DTS22, Hei08, HvSST14, Kol18, Kol21, MZB04, OO03, PPN20, SSDM21, SK05, WENG09, ZCC04]. angular [BS18]. anisotropic [Sab19b, SS18a, SS19a]. Announcement [Ano99e, Ano00a]. anonymous [Eğe09]. antiferromagnets [HK14]. antithetic [AJC16]. appearing [MPC03]. Appl [Hal05a, LP13, Oga97, Tuf98]. Application [BGSR08, CRS14, KS00, AJC16, FIN02, KK09, MH12, MM00, OO03, PP05, SZKS21, FGM17, Lej03]. application-based [MH12]. Applications [FP99, LM05, Ökt96, UŠ96, BM19, DM10, Har19, KD99, KD04, ÖG09, PR19, PRS05, SRKL96, Sab16c, SK18, TTEA01, Cos01]. applied [Aze12, BFP97, IOR21, LP12, MK06, Nin03, NÖ09b]. Approach [DMZ03, BCZ05, DK98a, Gui97, KLP14, LLM16, LL14, Lej04, MZ98, Min01, NMH04, PP19, Sha10]. approaches [PS05]. Approximate [EM03, ES10, Kan95, Ego20, Khi00]. Approximated [GHT00]. approximately [Zhe13]. Approximating [LN04, BCR11, Hab11]. Approximation [BEH16, GA99, Hid20, KS00, Kaw07, KP02, LS97, PRS05, Tuf04, BFP09, BST10, BBG15, BH18, Cap01, CP02, DM10, Gob01, KLW21, KT11b, KW02, LP12, Mal07, NY19b, Nin03, OO03, OY19, Raj19, SH22, Voy97, Wel06, Wih01, YY18, Yam21]. Approximations [BLNSP06, CL02b, DMZ03, EZ04, Ego07, New01, FM01, MT13, NY19a, Rot07]. aquifers [KS04b]. arbitrary [JWK19]. area [ES17, YY18]. Arithmetic [LPT03, JS07, JS10]. Arrays [Lik98]. article [Oga97]. Asian [JS10, BK14, JS07]. aspect [WN19]. aspects [dBDD01]. assisted [MDMS20]. Assumptions [FGM^{+}01]. Asymptotic [FGM^{+}01, NT97, NZ09, FT00, HS22, OY19, Shv03]. Asymptotical [GN05]. Asymptotically [NS07]. asymptotics [SS20b]. Atmosphere [SA96]. attachment [Gui08]. attack [Vid07]. autocorrelation [Man03, Nak97]. autologistic [ZNS10]. Automatic [ECLR21, KT11a]. availability [MZB04]. average [JS07, JS10, LPT03, SS15, SS18b]. averages [Ari15]. averaging [LP12].

Backward [SS01, BMO01, IOR21, MS16, NY19a, PG19]. Balance [GN99, Gol03, NP04a]. Balanced [KS06a, FP99]. Balking [CS96]. ball [KM22]. Band [BAO^{+}04, LL14, NMH04]. bandwidth [BZ20]. barrier [CK18, FHS13, Sag11]. Barriers [MPZP04, CP02, DMR16]. base [Xia02]. Based [LT04, SUZ04, BGSR08, CPSH07, CP15, DK98a, Ego20, FGD13, FGM17, FS12, Gri10, IIO14, Kaw06, KS15, KPSZ96, LM05, MC20, MH12, NK06, New01, Raj19, Sab16d, SS14a, SAKG15, SH22, Tur19, WENG09, YK08, Zhe13]. basket [DM10]. Baxter [KKS13]. Bayesian [BZ20, BBBR19, CCG15, LWC18, PP19, PPN20, PP21, Row02, WENG09]. be [Hal04, Hal05a, Hal05b]. Becker [Gui08]. Behavior [Oll01, GN05]. behaviour [FP02, Shv03]. benchmark [CA12, PWY99, SNDS14]. Berlin [Ano00h, Ano02e]. Bermudan [KS04a, LL21]. Bernstein [SH22]. Berry [Bis09]. Beryllium [HKHV98]. Bessel [Alf05, MG10]. best [DK06]. beta [CCG15, Voy98]. beta-distribution [Voy98]. Between [DT01, Nao95, Khi00]. beyond [NMH04]. bias [IP17, NT97, TOTAI18]. biased [PÖ20]. biasing [MZB04]. Biharmonic [AS95, SS03]. Binary [Nek20, Nek16, PMW10]. biomedical [TTEA01]. bit [Nek20]. bivariate [KRSJ17]. Black [Sin14]. Blodgett [SZKS21]. Board [Ano96a, Ano96b, Ano96c, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d, Ano00e, Ano01a, Ano01b, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano04, Ano05a, Ano05b, Ano05c, Ano06a, Ano06b, Ano06c]. bodies [MR04]. Body [ZZA21]. Boltzmann [BQA03, CRS14, FM01, KS95a, Khi00, Nek03, PW01, PS05, Rog99, TM20, Wag08]. bond [AHT04]. Bootstrapping [Man03]. bound [DK06]. Boundaries [DK98b, SL14]. Boundary [Hau00c, KMS04, MKL01, ST95, SA96, SM09, Sim95, CA12, GP19, Hal16, HBA16, NVDA07, Pöt12, Rog99, Sab08, SM12, Sab16a, Sim18, Wel06]. bounded [BJ01]. Bounds [KS04a, AH12, BDGZ20, FL10, KS06b, Rud10, Sha10]. Branching [RKM04, ES11]. Breusch [Man03]. bridge [Bee21, JWK19]. Brownian [AG03, CP02, CL01b, DMR16, GA99, GP19, JWK19, KS00, KPV18, Oga01, Osa01]. BSDE [Lej01]. BSDEs [KLW21, LL13]. Budapest [HKHV98]. building [Zal00]. Burgers [BFP97, JLH10]. burst [SE18]. Bursts [KPSZ96].

cadmium [SZKS21]. calculating [AD99, Ego20]. calculation [EM17, KM15, PWY99, Zhe13]. Calculations [BP98a, BP98b, KLR^{+}03]. calculus [BCZ05, NY19a]. Calibration [ELZ11]. called [Oga01]. can [Hal04, Hal05a, Hal05b]. Capacities [Com01]. Carlo [Ano99e, Ano00a, Ano00g, Ano02e, Ano03b, DS10, Hal05a, JS10, LP13, Oga97, ÖG09, Sab04b, Tuf98, ZC19, ATBM14, AAD04, Ant96, AE15, Ant15, Aro04, Ars98, Ars07, AD99, Aze12, Bab99, Bal08, BHA18, BCZ05, BQA03, BK14, Ben16, BP02, BP97, BP98a, BP98b, BS18, BOTAZ19, BDGZ20, BAO^{+}04, BG01, But03, CL01a, CL02a, CCMZ08, CA12, CRS14, CP01, DL14, DK98a, DMZ03, ELZ11, ES17, EUW98, EW02, ER06, Erm11, FVK16, FVK17, FM01, GM04, Gri10, Gri14, Gri17, Gui97, Hal04, Hal05b, Hal06, Hal08a, Hau00b, Hau00a, Hei95, HvSST14, Hor02, HPY07, HMG01, JS07, KSPZ20, Kaw07, KD99, KD04, KS95a, Khi00, KPSZ96, KM15, Kra01, KRSV99, LL11, LCRK18, LOR18, LT04, Lej04, Leo06, LM05, Lik98]. Carlo [LK02, MT13, MZ98, MZB04, Mar10, MKL01, MH12, MR04, McL11, MM00, MP02, MWMS18, NEBW20, NT21, NXÖ18, NPM^{+}06, NMH04, NÖ09b, Ökt96, ONZ99, Pan15, Pap04, PW01, PG19, PWY99, Ple00, PGS09, PS98, PIR04, Pöt12, RS21, RST96, Raj19, RBB21, Rog99, Row03, Rud10, SA96, SK97a, Sab16d, SE18, SP20, SD96, SNDS14, Sen01, SAKG15, Sin14, Smi98, SK05, SS07, SM08, SS14b, SS19b, Sta95, Sug04, TOTAI18, TM20, TTEA01, Tuf96, Tuf04, UV00, VAYT20, VA04, VDM00, Wag08, War18, YJH21, ZPK02, ZCC04, mMSD04]. Carlo-Based [LT04]. cascades [KK09]. case [EUW98, Erm11, PP03, PW01, RJG13]. CAT [AHT04]. cathodoluminescence [SK18]. cavity [HBBA15]. Cellular [BAO^{+}04]. censored [LL14]. central [NO09a, Gol03]. centres [Gol04]. certain [Tak96b, Tur19]. CFTP [BN15, FN09]. Chain [FVK16, FVK17, LK02, FN09, NB19, Rud10, YY18, MWMS18]. Chains [LT04, Mat99, Ari15, Hal21, Smi98]. change [Ave04, KLR^{+}03]. changing [Erm11]. chaos [Ego20, NR02, SS14a, SS17, YK08]. characteristics [EM17]. charge [YJH21]. chemistry [KW02]. choice [Eğe09]. chord [MR04]. CIGS [RBB21]. cipher [FVK16, FVK17]. CIR [Alf05, Hal15a, Hal15b]. circular [BZ20, SL14]. circular-shaped [SL14]. class [BJ22, EM03, KKS13, Lin06, Oga01, Wag15, Yan13]. classes [Tur19, Zal00]. Classification [LTD01]. Clinical [Nad07]. clustering [BN15]. clusters [LCRK18]. coagulating [KS01]. Coagulation [DT01, GZ01, Gui99, SK00, SLP07, WK05, Bab99, EW01, FG04, KS03, SRKL96, SK97a]. coefficient [BMO01, CL01a, KSNS15]. Coefficients [Pap04, Row02, NZ09, SK97a, SS14a, Spa21]. coherent [ATBM14]. Coin [NP04b]. collision [KS95a]. collocation [SM12]. COM [KRSJ17]. COM-Poisson [KRSJ17]. combustion [BC11, MK06, SH08]. Comments [Tuf98]. Communication [Wih01]. Comparative [Nao95, Raj19]. Comparing [BOTAZ19, LL20]. Comparison [Bea09, BFP97, Har19, Ima13, LT04, Nad07, KSPZ20, Lin06, RST96, SD96]. Competency [Sin14]. Complexity [Pag07, Lev16, Nek16]. complicated [ST00]. Component [DGKP08, Gri17]. Component-by-component [DGKP08]. components [ONZ99]. Compression [SUZ04]. Computable [KT11a]. Computation [Hei04, Cap01, Hei95, JS07, JS10, MS16, PR19]. computational [BM19]. Computations [Nao95, BDGZ20, DK98a, FGM17]. compute [CL01a]. computed [TTEA01]. Computer [KS00, PGS09, Ben16]. Computing [BFP09, BL15, KRSJ17, Nad08a, Ari15, ES17, ES20, MQH14, RS19, Rud10]. concentration [SL14]. Concept [BP98b]. Concrete [MPZP04]. condition [SS01]. conditions [CA12, HBA16, Sab16a, Wel06]. conducting [YJH21]. cone [Ste00]. Conference [Ano99e, Ano00g]. confidence [LL14, MM20, RS21]. Congruence [Ant95]. Congruential [NS07, AM17, EUW98, GN05]. Connection [DT01, Khi00]. Consensus [SK15]. conservation [BJ01]. Constant [CP01, YJH21]. constraint [KLW21]. Constructing [Hal15a, MM20]. Construction [Mor02, Mor05, Yag02, DGKP08, Hal16, Mor08]. constructions [Bee21]. Contaminant [SVH^{+}04]. continuous [BMO01, BJ22, IP17, Khi00]. continuous-time [BJ22]. contribution [BS18]. Control [GHT00, NHD06, Pag07, PGB98, BG13, ECLR21, IOR21, KT11a, KLP14, OY19, Pöt12]. Convective [SA96]. Convergence [BF01, BK95, Gol04, KW97, KP02, Rey17, AJC16, BH18, IK00, Kab05, KHO97, KS16, Wel06, Zhe13, BT96]. convex [MR04]. copula [Sak10]. copulas [KRSJ17]. Core [HKHV98]. Corput [FIN02, IM04]. Correction [ÖG09, IP17]. correlated [ABKT18, SM03, ZPK02]. Correlation [Nak98]. correlations [Rog96]. Corrigenda [Hal05a]. cost [RV99]. Couette [BHA18, HBA16]. Coulomb [VA04]. Coupled [BP98b]. Coupling [BP02, NB19]. Covariance [Row00, KS14]. covariate [IN17]. crack [JLH10]. credit [DL14, Sak10]. criterion [KS14]. Critical [ES11, HK14]. Cross [ZPK02, Ant15]. Cross-correlated [ZPK02]. cross-section [Ant15]. crossing [GP19, Pöt12]. cryptanalysis [FVK16, FVK17]. cubatures [AE15]. cube [PC04]. cuckoo [EHE18]. curves [LL20]. cusum [Hab11]. cusumsq [Hab11]. cutoff [FM01]. CVA [ATBM14]. CVaR [BFP09]. cylinder [SKL09]. cylinders [Sab16a]. cylindrical [PGS09]. CZTS [RBB21].

D [BJ01, CA12, FM01, KS05, Kur95a, NPM^{+}06]. Dagger [ZCC04]. Dagger-sampling [ZCC04]. Darcy [SKL09, SS17]. Data [Nad07, LL14, OO03, PPN20, Tor20, ZZA21]. de-biased [PÖ20]. debiasing [McL11]. decomposition [Ant95]. decompositions [Nek20]. Decreasing [FP02]. Deep [BP97, PWY99]. degenerate [Wih01]. densities [DMR16, Nek20]. Density [BT96, LH04, Nao95, BZ20, CLP17, ES17, ES20, Kaw06, YJH21, ZZA21]. Dependence [Nak98, WLD21]. dependent [CP02, KNS04, NÖ09b, PP19]. depending [KM15]. Depositing [NPM^{+}06]. depth [MM00]. derivative [MH12, MH13]. Derivatives [KS04a, CCMZ08, EBSY18, KSC11]. Describing [Tor20]. descriptive [Bea09]. Design [Ano96d, NPM^{+}06, FGM17, WN19]. design-based [FGM17]. detector [MM00]. Determination [NK06]. deterministic [BFP97, Hei95, Wag10]. Deviational [Wag08]. Deviations [Com01, KM11b, KS06b]. Devices [BAO^{+}04, VA04, NVDA07]. diaphony [PS10]. difference [EW02]. different [KRSJ17, RST96]. Differential [Ano99e, Ano00g, BT96, BF01, Hau00b, Hau00c, Kan95, KM95, LN04, BMO01, BEH16, BH18, Buc04, EZ04, Ego07, ES10, EM17, EP19, ÉM13, FP99, GR08, Hab12, HS22, Hid20, KM02, LOR18, LWC18, MPC03, NY19b, NT21, NP04a, PG19, Pri01, RJG13, Rot07, WENG09, Xia96, Yan13, Zhe13]. diffusing [KS01]. Diffusion [CP01, ELV10, HMG01, KT11b, KP02, NPM^{+}06, CLP17, FHS13, Hau00a, Lej03, MS14, Raj19, Rey17, SL14, SLK15, Sab16a, Sab16b, Sab17, SK18, Sab19b, SP20, SS18a, SS19a, SS21b, Wih01, YJH21, ZC19]. diffusion-reaction [SK18]. diffusion-recombination [SS21b]. Diffusions [BLNSP06, BST10, Bis09, Gob01, MG10, Oga01]. Diffusive [Oll01]. Digital [LTD01]. Digitized [SM04]. dilute [BHA18]. dimensional [BEH16, BH18, CRS14, CJV16, ÉM13, Hid20, HBBA15, Kol20, Mor02, Mor05, Mor08, Pan15, PS10, Rey17, Sim95, SS14b, War18]. Dimensions [ELRU04, LW10, SS15, SS18b]. Direct [Gui99, KRSV99, WK05, Khi00, MZB04, Rog96, SN13]. Dirichlet [AS95, Bou05, NÖ09b, SS95, Sab16a]. Discrepancy [GP12, IM04, Mor99, Mor04, Ökt96, AH12, DK06, DGKP08, FL10, Mor98, Mor02, Mor05, Mor08, MM12, Nk16, ÖG09, PC04, RST96, Sha10, Tuf96, Tuf98, Xia96]. Discrete [SSL04, Hal21, HS22, KM11a, OO03, PS05, Voy97]. discrete-stochastic [PS05, Voy97]. discretely [Bis09]. Discretization [KLW21, Alf05, NY19a, OY19, Pri01]. discretized [Wih01]. Dispersion [Kur95b, KS95b, Kur95a, Kur97, SA96, SK98, CCG15, KOSY01]. distance [NS09, Rey17]. Distributed [PGB98, Row02, Ave04, Buc04, FKM08]. Distribution [HPY07, SUZ04, BS18, CRGF18, FN09, Hab11, Kol20, MP12, Mak15, MM20, MR04, Mis07, NZ09, SK18, SSG99, Tor20, Voy98]. distributions [Ego97, FT00, Nad08b, PR19]. DNS [KOSY01]. domain [CL02a]. domains [NÖ09b]. Döring [Gui08]. Double [FHS13, CL01a, Kol21]. Double-barrier [FHS13]. doubly [MS16]. draws [Rei20]. Drift [KSPZ20, DMR16, Sab16b, Sab17, SK18, Sab19b, SP20, SS18a, Spa21]. drift-diffusion [SP20, SS18a]. drift-diffusion-reaction [Sab17]. drifts [Osa01]. driven [AG03, BHA18, GR08, Hau00b, Mar10]. DSMC [HBBA15]. dual [NB19]. duration [But03]. Dynamic [HMG01, PO04, Ave04, Hei08, Man03, MZ98]. Dynamical [MM12, Mor04]. Dynamics [Sei04, ZPK02, LLLP12, EW01, LT08].

ECDLP [Vid07]. Edgeworth [KM02]. Editorial [DS10, Ano96a, Ano96b, Ano96c, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d, Ano00e, Ano01a, Ano01b, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano04, Ano05a, Ano05b, Ano05c, Ano06a, Ano06b, Ano06c]. Effect [IN17, ZPK02]. Effective [SM04]. effects [WENG09]. Efficient [Gob01, HPY07, KSC11, ABKT18, AM15, DTS22, Gri14, JML20, JLH10, KW02]. eigenvalues [DK98a]. Elasticity [CP01, SS02]. elastostatics [KS15]. Electron [BP98b]. electronic [Ben16]. Element [BP02]. ellipsoids [RS21, SS20a]. Elliptic [MKL01, Sab08, SS21a, SSDT21, SS20a, Sim18]. emission [EN20]. emissions [eZN22]. Empirical [SSS06, BG13, FP02]. energy [BS18, KK09]. engineering [KD99, Lej03]. entrapment [HTKM19]. Entropy [CL02b]. environment [ES11]. Equation [BQA03, DT01, KNS04, NAKS04, WK05, AG03, Aze12, Bab99, BFP97, CA12, CRS14, CJV16, EW02, GM04, GA99, GR08, JLH10, KSNS15, KS95a, Khi00, KW97, KS01, KS03, KW02, LT08, LWC18, Man03, Oga01, PW01, PS05, Rog99, Rot07, RJ20, SRKL96, SK97a, SS03, SL14, Sab19a, SS17, TM20, Wag08, Wag15]. Equations [Ano99e, Ano00g, Ars07, BT96, BF01, GN99, GZ01, Hau00b, Hau00c, Kan95, KM95, LN04, LS97, Lik98, NP04b, Sim95, Ant11, Aze12, BMO01, BEH16, BH18, Buc04, DKS^{+}98, EZ04, Ego07, ES10, EM17, ER06, EP19, ÉM13, FP99, FM01, Gol03, Gui97, Gui08, Hab12, HS22, Hid20, IK00, KLP14, KS15, KM02, LOR18, MPC03, NY19b, NT21, Nek03, NP04a, PG19, PS98, Pri01, Rie99, RJG13, SSL06, SM09, SLK15, SS21a, SSDT21, SS20a, Sim18, SS19b, WENG09, Xia96, Yan13, Zhe13, dBDD01, Gui99]. equilibrium [Ari15]. Equity [JWK19, MBK06]. Equity-linked [JWK19]. Erratum [JS10, LP13, Oga97]. Error [Kan95, PS98, Rud10, Tuf04, AH12, AP04, KT11a, KS03, NZ09, OY19, Owe06, RJG13, SS03, TOTAI18]. Errors [GN99, SSS06, SS07, Hal04, Hal05a, Hal05b, SS20b]. escape [SP20]. Esseen [Bis09]. estimate [Sha10]. estimated [Hal04, Hal05a, Hal05b]. Estimates [CP01, SS07, CP02]. Estimating [Rei20, SM04, Spa21, LL14]. Estimation [AD01, CRT02, Nao95, NHD06, Pap98, Tuf04, eZN22, AN12, BJ22, BZ20, CLP17, KSC11, LWC18, MP12, MM00, NO09a, OW07, Oga08, Pit06, PS98, Pöt12, RJ20, SS03, SH22, ZZA21]. Estimations [Kan95, KS03, Smi98]. estimator [CK18, McL11]. Estimators [SSS06, AJC16, BOTAZ19, Erm11, GLP17, NT97, PÖ20, SD96, SM08, SS18b, TOTAI18]. Euclidean [Ant95]. Euler [BT96, BEH16, BH18, CLP17, DKS^{+}98, Hid20, Kan95, KHO97, KM02, KP02, NP04a, NZ09]. Eulerian [DK98b, KS04b, Nak98, SK03]. evaluation [AP04, EM03, MT08, Mis07]. Evaporation [Ple00, SZKS21]. Event [Nad07, FGM17, MS14, PPN20]. evolution [AG03, Gui08, Rog96]. Exact [ÉM13, FG04, JS07, KM11a, MG10, Nak97, Zhe13, JS10]. Examining [TM20]. examples [PR19]. exchange [CL01a]. excitations [Sab08]. Excursion [Hau00c]. existence [PP21]. exit [BL15]. exit-time [BL15]. Expansion [Sab08, Ego20, KT11a, OY19, SS17]. expansions [KM02, NT97]. expectation [Rud10]. expectations [Ego07, ES10, Ego20, Zhe13]. experiment [SS14b]. Experimental [Ano96d, KSPZ20]. Explicit [MK06, DMR16]. Exploitation [CCMZ08]. Exponential [KS06b, KK09, NK06]. exponential-normal [KK09]. exponents [Wih01]. expression [Nak97]. extensible [Har16]. extension [BMS09]. extensions [Sab19a]. Exterior [SS95]. Extrapolation [Pag07]. extreme [AN12].

factor [Hal15a]. Factorization [Row00]. Fallout [KPSZ96]. Fast [CPSH07, CL18, LP11, LP13, SLP07]. feedback [MC20]. Feistel [AM17]. Feistel-inspired [AM17]. Feller [PR19]. Feynman [LOR18, MT08]. Fibonacci [AM15]. fictitious [KS95a]. Field [Hor02, HK14, HBBA15, KM22]. Field-induced [HK14]. Fields [KS06c, BK95, CL18, KKS13, KS06b, LP11, LP13, Lev16, PMW10, PO04]. films [BS18, RBB21]. filter [PRS05]. filtering [FP99]. Filters [New01]. Filtration [KS04c]. Finance [LP12, KT11b, MQH14, Cos01]. financial [ELZ11, Har19, KSC11]. Finite [Ars07, BP02, BFP97, BL15, KM11a]. finite-range [BL15]. firefly [EHE18]. First [Ano99e, Ano00a, BLNSP06, Ben16, FHS13, MPC03, Rot07]. first- [MPC03]. first-passage [FHS13]. fissured [Lej04]. fitness [Gui08]. fitting [TTEA01]. Fix [Voy97]. Fixed [SSL06]. floating [Nek16]. Flow [WK05, BHA18, HBBA15, HBA16, KS04b, KS05, Kol18, MPC03, SK03, SKL09]. Flows [KS95b, KSK97, SK97b, BP02, Min01, SK00]. fluctuation [SLK15]. fluctuation-induced [SLK15]. Fluctuations [ZPK02, SL14]. Fluid [HMG01, KS95b]. flux [SL14]. fly [FGD13]. FMRI [Row03]. Footprint [KRSV99, KLR^{+}03]. forced [MZ98]. Foreword [Sab04a, Sab04b]. form [KK09, NB19]. formalization [LLLP12]. Forms [Bou05]. formula [Ego20, ES20]. formulas [ES10, Zhe13]. Formulation [ST95]. Forward [SS01, LOR18, NY19a]. forward-backward [NY19a]. Fourier [Ima13, KS06c]. Fourier-Wavelet [KS06c]. Fourth [Ano00f]. fractal [Kol20]. fractional [AG03, GA99, GR08, KPV18]. fractured [CL02a]. Fragmentation [Gui99, Wag10]. framework [LL11]. Fredholm [SS19b]. free [Nek03]. Freivalds [JML20]. Frequency [BAO^{+}04]. Frobenius [Mor08]. frog [EUW98]. frontier [SSDM21]. Frontmatter [Ano14a, Ano14b, Ano14c, Ano14d, Ano15d, Ano15a, Ano15b, Ano15c, Ano16d, Ano16a, Ano16b, Ano16c, Ano17a, Ano17b, Ano17c, Ano18a, Ano18b, Ano18c, Ano18d, Ano19d, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22]. Full [BAO^{+}04, NMH04]. Full-Band [BAO^{+}04]. fully [IOR21, KLP14]. function [CA12, CRS14, KS14, MR04, Nak97, Xia96]. Functional [CP15, PP05, SS03, Buc04, EM03, EZ04, Mal07, NO09a, Sag11, Zhe13]. functionals [Cap01, Ego07, ES10, Ego20, Yam21, Zhe13]. functions [AD99, CDGG21, EM03, FT00, Gri17, Hab11, ST00, Zal00].

G [BOTAZ19, TOTAI18]. Gains [KLW21]. Gains-process [KLW21]. Gamma [BP97, BP98b, BBG15, SAKG15]. gamma-rays [SAKG15]. GaN [KSPZ20]. gas [BHA18, BC11]. gas-phase [BC11]. Gaussian [AP04, BK95, CL18, Ego97, FGD13, Gri10, Gri17, JML20, KM22, KKS13, KS14, KS06c, LP11, LP13, Lev16, PP03, PMW10, PP04, Tur11]. gelation [EW01]. general [LT08, McL11]. generalization [DT01]. Generalized [BP98b, FGM^{+}01, Gui08, FIN02, KS16, KPV18]. Generalizing [LW10]. generated [EZ04, IM04, Mor98, Mor99, Mor04, MM12, Nad08b, SSL04, eZN22]. Generating [Ste00, Gri10, Yag00]. Generation [ASTY19, Chi13, UŠ96, CL18, Eğe09, FGD13, Nek16, Tak00]. Generator [Sug95, Ant95, BOTAZ19, MQH14, Sug04, Yag02, YK08]. Generators [GGP06, NS07, AM17, AM15, EUW98, GN05, Ima13, MH12, MH13, NS09]. generic [BMO01]. Genetic [LK02, Sha10]. geometric [ES20, KS16, RS21, Rei20, Xia02]. Geometrical [VDM00]. Geometry [HTKM19, Lev16]. getLHS [BOTAZ19]. getRDS [BOTAZ19]. Gibbs [Row00, Spa21]. Gillespie [Raj19]. given [Kol20, RL18, Tur19]. Global [Kol18, Kol21, SS07, SVH^{+}04, KT11a, ME09, Sab19a, SS21a, SSDT21, SBH04, SK05, ZYD19]. Godfrey [Man03]. Good [Pap04, PS10, VAYT20]. governed [SK97a, SLK15]. governing [KS01]. GPU [AM15, CPSH07, LCRK18]. GPU-based [CPSH07]. gradient [BJ22, BGSR08]. graph [Lej03]. gravity [BHA18, HBBA15]. Greeks [JWK19]. Green [CRS14]. Green’s [CA12]. Grid [LM05, CL02a, SS21a, SSDT21]. Grid-based [LM05]. gridless [Lej04]. grids [SSL04]. Growth [NPM^{+}06, Hei14, SRKL96]. GWAS [KS16].

Halton [BM19, FL10, MC04, NEBW20, Owe06]. Hammerstein [GA99]. Hamming [Tak96a]. hazard [PP19]. heat [Sab19a]. Heath [CK18]. heavy [ZZA21]. heavy-tailed [ZZA21]. hedging [BCZ05, IIO14]. Height [BP98a, KLR^{+}03]. Helmholtz [CA12]. Hermite [PG19]. Heston [BBG15, CK18, MH13]. Hidden [EN20, Cap01]. High [BQA03, ELRU04, Kur97, MQH14, Yam21, KK09, MK06, ONZ99, SS14b, War18]. high- [ONZ99]. high-dimensional [SS14b, War18]. High-Reynolds [Kur97]. high-temperature [MK06]. higher [GP12]. Highly [Pap98, DTS22]. hitting [CP02]. HJB [KLP14]. HMM [eZN22]. Hole [NPM^{+}06]. Homogeneous [GN99, KSSV03, Nak98, SK98, BC11, BK95, FM01, Nak97, NP04a, SE18]. homogenization [LLM16, Lej01]. Horizontally [SK98]. Horner [Yag00]. Hybrid [BC11, DL14, Ökt96, Tak00, EHE18, ÖG09, SE18]. Hybrid-Monte [Ökt96, ÖG09]. hydrometeorological [PO04]. hyper [DM10, RS19, RS21]. hyper-ellipsoids [RS21]. hyper-rectangular [DM10]. hyper-volumes [RS19]. hyperbolic [LPT03, Rot07]. hyperspheres [AW10]. hypersurface [ES17]. hypothesis [KS14].

ICM [Row02]. identifiabilities [MWMS18]. identification [HKN12]. identity [Rei20]. II [BT96]. IID [ES11]. illumination [SBH04]. illustration [Mis07]. IMACS [Ano00a, Ano02e, Ano03b, DS10, Sab04b].

Image [DSGZ01, SUZ04]. imaging [SK18]. Implementation [HvSST14, BMS09, LCRK18, NXÖ18]. Implementing [PÖ20]. Importance [BP97, Sta95, BFP09, CRT02, FS12, Kaw06, MS14, ME09, Shv03, UV00]. improve [BG13]. Improved [FVK16, FVK17, FL10]. Improvement [CP01]. improves [AM17]. Improving [Pöt12]. imputation [ZNS10]. incoming [SL14]. Incorporation [VA04]. Increasing [Sak10]. independent [ASTY19]. index [MC20, Sin14]. Indices [SS07, Lin06, SM08]. induced [HK14, SLK15]. inequalities [Bis09]. inference [PPN20]. Infinite [Pan15]. Infinite-dimensional [Pan15]. inflated [IN17]. Influence [HKHV98, NPM^{+}06]. Information [Lev16, LNO15]. inhomogeneous [KS01, Yam21]. initial [BDGZ20, NVDA07]. initial-boundary [NVDA07]. inner [Sak10]. innovation [LP12, Raj19]. input [GA99, RL18]. inspired [AM17]. integers [FKM08]. Integral [Ars07, Mis07, NP04b, ST95, GA99, IK00, PS98]. integrals [AD99, EM03, EZ04, KM15, Mal07]. Integration [LK02, Mat99, DM10, Dic06, FIN02, Kab05, Pan15, RST96, SS20b, ST00]. integro [RJG13]. integro-differential [RJG13]. Interacting [CL01b, Oll01, CDGG21, Osa01]. Interaction [VA04]. interactions [Wel06]. Interest [CP01, LCRK18]. intermediate [JWK19]. intermittent [SK00]. International [Ano96d, Ano00g, Ano99e]. intersection [RS21]. interval [BJ01, MM20]. Invariant [CLP17, PR19]. inverse [KSNS15]. inversion [SN13]. Investigation [Kab05, BM19, KK09]. Investors [HR02]. involving [BEH16, BH18, ÉM13, Hid20]. Irrational [Sug95]. irregular [Yam21]. isomorphic [Ego97]. Isotropic [Kur95a, Nak98, CL18, KS15, Nak97, SL14]. issues [Min01]. Itô [NY19b]. iteration [BKS06]. Iterative [DKS^{+}98, PS98, SS02, SL10]. IV [Ano02e, Sab04b]. IVth [Ano03b].

Jackknife [IP17]. joint [CCG15, PPN20]. Jointly [Row02]. July [Ano99e, Ano00g]. Jump [BLNSP06, BL15, FHS13, HBA16]. jump-diffusion [FHS13]. Jumps [KP02, HKN12]. just [VAYT20].

Kac [CJV16, LOR18, MT08, NR02, Nek03]. Kac-type [LOR18]. Kernel [Nao95, BGSR08, BZ20, Nad08b, ZZA21]. kernel-based [BGSR08]. Kernels [Gui99]. killed [CP02, Hau00a]. kind [NB19]. kinetic [SE18]. kinetic-thermodynamic [SE18]. Kinetics [HKHV98, SLK15]. Korobov [Pap04]. Kosterlitz [HK14]. Kou [Bal08]. Kronecker [Chi13]. Kusuoka [Nin03].

lagged [AM15]. lagged-Fibonacci [AM15]. Lagrangian [CK04, KS01, Kur95b, KS95b, Kur97, KSK97, KOSY01, KSSV03, KLR^{+}03, MPC03, Nak97, Pit06, PGB98, SK97b, SK98, SS01]. Lamé [SSL06]. Land [KPSZ96]. Land-based [KPSZ96]. Langevin [MDMS20]. Langmuir [SZKS21]. Laplace [SSL06]. Large [Com01, HR02, SVH^{+}04, KM11b, SSL04, SM09, SL10, Sab16d]. largest [Nad08a]. latent [Lin06]. Latitudes [BQA03]. Lattice [Pap04, ELV10, GP12, HK14, RJG13]. Law [BT96, BJ01, HR02]. Layer [SA96]. layers [BS16, CRS14]. leap [EUW98, KT11a]. leap-frog [EUW98]. learning [KLW21]. left [ABKT18]. left-tail [ABKT18]. length [MR04]. Levy [KKS13, Kaw06, KT11b, Leo06, Mar10, YY18]. Libor [BMS09]. life [SK18]. like [SC96]. likelihood [BJ22, LWC18]. likelihoods [Rei20]. Limit [GLP17, Gol03, BK14, NO09a, SS15]. Linear [AAD04, DMZ03, EUW98, Hal06, Lik98, NS07, PGB98, AM17, Ant11, ER06, GN05, Hal08a, IM04, Lej01, Mor98, NT21, ONZ99, PP19, Rie99, RL18, SM09, SL10, Sab16d, War18, Zal00]. linearized [PS05, TM20]. lines [ES17, ES20]. linked [JWK19]. Lipschitz [NZ09]. loads [IOR21]. Local [Hau00c, Kur95a, BEH16, BH18, ÉM13, LWC18, NY19a]. Local-Isotropic [Kur95a]. Log [KS04c, ABKT18, BJ22, LL20]. log-likelihood [BJ22]. log-normal [ABKT18]. log-rank [LL20]. Log-Stable [KS04c]. logistic [PP21]. Lomax [NK06]. long [IP17, Yag02]. long-period [Yag02]. longitudinal [PPN20]. lot [AW10]. Lottery [BG01]. Low [Mor98, DGKP08, Har16, Mor02, Mor05, Mor08, MM12, Nk16, PC04, RST96, Tuf96, Tuf98, Xia96]. low-discrepancy [DGKP08, Nk16, PC04, Xia96]. low-WAFOM [Har16]. lower [BDGZ20, Sha10]. LSMC [WN19]. Lyapunov [Wih01].

M [BOTAZ19, TOTAI18]. M/G/1 [BOTAZ19, TOTAI18]. machine [KLW21]. maintenance [SC96]. Malaysian [MBK06]. Malliavin [BCZ05, NY19a]. management [DL14]. Manhattan [Ben16]. mappings [YK08]. Maps [Mor98, Mor99]. margin [BDGZ20]. marginal [Rei20, Sag11]. marginalized [IN17]. Market [MBK06, BMS09]. Markov [Ari15, CRT02, DMZ03, EN20, FN09, FVK16, FVK17, Hal21, LT04, LK02, Mat99, MWMS18, NB19, Rud10, YY18]. Markov-Chain [MWMS18]. Markovian [AK02, BBBR19, Cap01, CHK01, EN20, Pap98]. Marsaglia [AW10]. martingale [LL14]. martingales [PG19]. Maruyama [BEH16, BH18, Buc04, Hid20, Kan95]. Mass [WK05]. Masthead [Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14e]. material [BS16]. Mathematical [Ano96d, Ant15, LLLP12]. Matlab [SAKG15]. Matlab-based [SAKG15]. Matrices [Row00, Gri14]. Matrix [LS97, JML20, Mal07, Sab16d, SZKS21]. matrix-based [Sab16d]. matter [SAKG15]. maximum [Ant15, Hid20]. Maxwell [FM01]. Maxwellian [PW01]. MCM2001 [Ano00a]. MCM2003 [Ano02e]. MCMC [BBBR19, FGM17, LTD01, Row02, ZNS10, ZYD19]. Mean [CP01, Hei04, Row02, CCG15, Hal21, IP17, PP21, RS19]. Mean-Reverting [CP01]. Means [Sug95, Hei95, SSL04]. Measurement [MPZP04, SSS06]. measurements [KSPZ20, MM00, TTEA01]. Measures [GN99, Hau00b, SUZ04, EZ04, ES20, FP02]. Media [KSSV03, SM04, BCR11, KS04b, KS05, Lej04, MR04, SK03, Smi98]. median [AN12]. Medium [KS04c, ELV10]. Memory [AM15, Buc04]. mesh [Kas17].

Mesoscopic [BP02]. Method [Aro04, Gui99, KS00, Lik98, MKL01, MP02, Nao95, NAKS04, Oga96, PGB98, ST95, AP04, Ari15, AD99, BPP01, Ben16, BFP97, BJ01, CL01a, CL02a, CDGG21, EM17, ES17, FVK16, FVK17, FP99, GM04, Hab12, HS22, HBBA15, KT11a, Kas17, Kaw06, KW97, Kol21, KS16, KM15, Mar10, McL11, MM12, NP04a, NZ09, Nk16, NXÖ18, Nin03, Oga97, OY19, PW01, PG19, PGS09, PO04, RS19, RBB21, Rog99, RJG13, SM09, SL14, Sab16b, SN13, SNDS14, SS17, SS18a, SS19a, SS20a, Shv03, SH22, SS19b, Sug04, TM20, VDM00, Yag00, YJH21, Zhe13, Cos01]. Methods [AAD04, Ano96d, Ano99e, Ano00g, Ant96, BP02, KS06a, Kra01, KS06c, LP13, LT04, LTD01, Oga97, Tuf04, AE15, Aze12, Bal08, BCZ05, Bee21, BG13, BDGZ20, BBBR19, CCMZ08, CJV16, CP15, DL14, Hal04, Hal05a, Hal05b, Hei95, IK00, JLH10, KSNS15, Kab05, KD99, Khi00, KS03, KOSY01, Lej04, MK06, MWMS18, RS21, RST96, Row03, SS02, SL10, SM12, Sab16c, Sen01, Tuf96, Tuf98, UV00, Voy98, ZYD19, Ano00a, Ano03b, DS10, Sab04b]. Metropolis [MDMS20]. microchannel [HBA16]. microelectronic [NVDA07]. microstructure [Oga08]. Milstein [KS06a]. Minimal [CL02b]. minimization [GK08]. minorization [Spa21]. misspecifications [IN17]. Mixed [NVDA07, AH12, CA12, PÖ20, SS01, Sab16a, WENG09]. mixed-effects [WENG09]. Mixing [Row02]. Model [CS96, EN20, Hor02, KNS04, Kur95b, Kur95a, KSSV03, Oga01, SK98, Bal08, BBG15, BMS09, BBBR19, CL01a, CL02a, CK18, ES17, EBSY18, Hal15a, Hei14, IN17, KRSJ17, KS01, KS04b, LPT03, LCRK18, Lin06, Man03, MH13, SK03, SE18, Sak10, SZKS21, ZNS10]. Modeling [KPSZ96, KS04c, SVH^{+}04, BC11, CCG15, CRS14, Gui08, Kol20, MPC03, NVDA07, PGS09, PO04, RL18, SH08, VMS08]. Modelling [SM03, Min01, Shv03, Voy98]. Models [Ano00h, BP02, CK04, KS95b, Kur97, KSK97, SK97b, SS01, BJ22, BK95, CRT02, CCG15, Ego97, ELZ11, Hei08, IP17, Kol21, KOSY01, Lin06, LWC18, NK06, PP19, PPN20, Pit06, PMW10, SSDM21, Wag10, Wag15, WENG09]. modes [LWC18]. modification [Ant95]. modifications [VDM00]. Modified [PGB98, Chi13]. Modulated [AD01]. Modulations [LTD01]. moduli [NS09]. Molecular [Sei04]. molecules [FM01]. Monaco [Ano00g, Ano99e]. monotone [BN15, Mor99]. Monte [Ano99e, Ano00a, Ano00g, Ano02e, Ano03b, DS10, Hal05a, JS10, LP13, Oga97, ÖG09, Sab04b, Tuf98, ZC19, ATBM14, AAD04, Ano00g, Ant96, AE15, Ant15, Aro04, Ars98, Ars07, AD99, Aze12, Bab99, Bal08, BHA18, BCZ05, BQA03, BK14, Ben16, BP02, BP97, BP98a, BP98b, BS18, BOTAZ19, BDGZ20, BAO^{+}04, BG01, But03, CL01a, CL02a, CCMZ08, CA12, CRS14, CP01, DL14, DK98a, DMZ03, ELZ11, ES17, EUW98, EW02, ER06, Erm11, FVK16, FVK17, FM01, GM04, Gri10, Gri14, Gri17, Gui97, Hal04, Hal05b, Hal06, Hal08a, Hau00b, Hau00a, Hei95, HvSST14, Hor02, HPY07, HMG01, JS07, KSPZ20, Kaw07, KD99, KD04, KS95a, Khi00, KPSZ96, KM15, Kra01, KRSV99, LL11, LCRK18, LOR18, LT04, Lej04, Leo06, LM05]. Monte [Lik98, LK02, MT13, MZ98, MZB04, Mar10, MKL01, MH12, MR04, McL11, MM00, MP02, MWMS18, NEBW20, NT21, NXÖ18, NPM^{+}06, NMH04, NÖ09b, Ökt96, ONZ99, Pan15, Pap04, PW01, PG19, PWY99, Ple00, PGS09, PS98, PIR04, Pöt12, RS21, RST96, Raj19, RBB21, Rog99, Row03, Rud10, SA96, SK97a, Sab16d, SE18, SP20, SD96, SNDS14, Sen01, SAKG15, Sin14, Smi98, SK05, SS07, SM08, SS14b, SS19b, Sta95, Sug04, TOTAI18, TM20, TTEA01, Tuf96, Tuf04, UV00, VAYT20, VA04, VDM00, Wag08, War18, YJH21, ZPK02, ZCC04, mMSD04]. Monte-Carlo [FM01, LOR18, MR04, MWMS18, Pan15, RBB21]. Morgenstern [Mak15]. morphology [BS18]. Motion [KS00, KSK97, SK97b, AG03, CP02, DMR16, GA99, GP19, KPV18, Nek03, CP02]. motions [Osa01]. Moving [DK98b]. MR1414863 [Oga97]. MR1434423 [Tuf98]. MR2338086 [JS10]. MTTF [CRT02, Pap98]. Multi [Pag07, LCRK18, PP19, Pit06]. multi-GPU [LCRK18]. multi-stage [PP19]. Multi-step [Pag07]. Multidimensional [Ars07, Bea09, DTS22, DKS^{+}98, NY19b, PO04]. multifactor [Sak10]. Multilevel [BK14, BDGZ20, Mar10, NEBW20, AJC16, GLP17, HvSST14, LCRK18]. Multiple [BMS09, GZ01, LWC18]. multiples [Tak96b]. Multiplicative [DT01, Gui99, NS09]. Multiscale [KS04c]. Multivalued [LN04]. multivariate [Dic06, NK06].

Nanbu [KW97, NT97]. nanocrystals [SZKS21]. nanosystems [PGS09]. Narrow [VA04, SP20]. Narrow-Width [VA04]. Natural [UŠ96]. Navier [Sim95]. negative [Ant11]. nested [FGM17, Lin06]. Nesting [War18]. nets [DK06, WLD21, Xia02]. network [LL21, MDMS20]. networks [ECLR21]. Neumann [CA12, MT13]. Neural [LL21, ECLR21, MDMS20]. neurologic [Row03]. neutral [Eğe09]. neutron [ONZ99, Sen01]. Newton [Hab12]. Nifty [Sin14]. Ninomiya [AJC16]. no [Hal05a, Oga97, Tuf98]. noise [GR08, Oga08, PP04]. Non [Ant11, CHK01, Hal06, Nao95, AP04, BBBR19, Ego97, FGD13, FGM17, FM01, KKS13, MR04, MWMS18, NZ09, SN13, Smi98, WENG09, War18, YJH21]. non-analog [Smi98]. non-Bayesian [WENG09]. non-constant [YJH21]. non-convex [MR04]. non-Gaussian [AP04, Ego97, KKS13]. non-identifiabilities [MWMS18]. Non-Linear [Hal06, War18]. non-Lipschitz [NZ09]. Non-Markovian [CHK01, BBBR19]. Non-negative [Ant11]. Non-parametric [Nao95, FGM17]. non-stationary [FGD13]. non-uniform [SN13]. nonalgebraic [Yag02]. Noncommutative [Com01]. nonconservative [LOR18]. nonhomogeneous [ELV10]. Nonlinear [New01, BHA18, BPP01, CRS14, FG04, KLP14, KS95a, KHO97, Oga01, PS98, dBDD01]. nonnegative [ZZA21]. nonrecursive [Yag02, YK08]. nonstationary [Gri10]. Normal [Tuf04, ABKT18, KK09, MM20]. Normalization [ELRU04]. note [Hab11, Hab12, HS22, KD99]. Nuclear [KPSZ96, MPZP04]. nucleation [SE18]. Number [GGP06, Kur97, Sug95, AM15, Ima13, MH12, MH13, MQH14, Sak10, Tak96b, Tak00, Yag02, YK08]. Numbers [Ant96, UŠ96, Ant95, EL18, Yag00]. Numerical [AS95, BF01, Hau00c, KSNS15, Mat99, MS16, SVH^{+}04, FIN02, Hal15a, Hal15b, Hal16, Hei95, IIO14, Kab05, KLP14, Min01, MPC03, OY19, PMW10, PO04, RST96, ST00, Voy97, VMS08, Xia96, Yan13, dBDD01, KSK97]. numerics [PP03, PP05]. Nyström [RJG13].

Object [DSGZ01]. observation [PRS05]. observed [Bis09]. Oceanic [CK04]. October [Ano00h]. ODE [MK06]. on-the-fly [FGD13]. One [SK98, BEH16, BH18, CJV16, ÉM13, Hid20, KKS13, PS10, Rey17]. one-dimensional [BEH16, BH18, CJV16, ÉM13, Hid20, PS10, Rey17]. One-Particle [SK98]. open [PGS09]. Operator [NAKS04, Ant95, Mor08]. Operator-Split [NAKS04]. Operators [DMZ03, LK02, NÖ09b]. optical [TTEA01]. Optimal [AD01, CHK01, CDGG21, CL02b, GHT00, LNO15, NS07, NHD06, PP03, Pap04, PGB98, Sei04, AD99, BM19, GG05, Kab05, Kas17, PRS05, WN19]. Optimising [Bee21]. optimization [EHE18, ME09, PS98, SS03, ZYD19]. Optimizing [Ars98]. option [BGSR08, DM10, ECLR21, LL21, PP05]. options [BCZ05, BK14, BKS06, CK18, GK08, JS07, JS10, LPT03, Sag11, Sin14]. Order [BLNSP06, MPC03, NY19a, NY19b, Rey17, Rot07, SS21a, SSDT21, VMS08, YY18, Yam21]. Ordinary [KS06a, LWC18, PP04]. Orlicz [KM11b]. Ornstein [KM11a]. outline [Hal04]. output [But03]. overview [BKS06]. oxygenation [MM00].

packing [AW10]. Pair [KS95b, Rog96]. Pairs [Kur95a]. papers [DS10, Sab04b]. parabolic [NÖ09b, Pri01, SS20a]. paradigm [PIR04]. Parallel [AAD04, DK98a, KMS04, MH12, MH13, PGB98, ZYD19, Ari15, LLLP12, Chi13, EUW98, LL13]. parallelepipeds [Sab19b]. Parameter [NHD06, Pit06, KM15]. parameters [IP17]. Parametric [Ars07, FGM17, Nao95]. Pareto [HPY07]. parking [AW10]. Partial [Ano99e, Ano00g, GR08, LOR18, LNO15, NT21, Nin03, PRS05, Pri01, Rot07, Xia96]. Particle [BP98b, KNS04, Kur95b, Kur95a, Kur97, KSK97, NPM^{+}06, Oga96, Oll01, SA96, SK97b, SK98, BFP97, BJ01, Cap01, CDGG21, Gol03, KS03, NVDA07, NR02, Nek03, Oga97, Oga01, Pit06, Rog96, SKL09, SS18b, Wag08, Wel06, ZC19]. Particles [KS95b, CL01b, KS01, KOSY01, Osa01, PGS09, SK00, SL14]. passage [FHS13]. past [NB19]. paths [CPSH07, SBH04]. Patterned [Row00]. PD [WENG09]. PDE [BCR11, Lej01]. PDEs [IOR21, SSL04, Sab08, SM12, SS14a, War18]. PDF [Hei14, KW02, SH08]. PDMC [ZC19]. Pearson [Tor20]. penalized [PPN20]. Penalty [KS00]. Penetration [BP97, MPZP04, PWY99]. percentage [Nad08a]. Perfect [CJV16]. perfectly [Wel06]. performance [MC20, MQH14, TOTAI18]. period [Yag02]. periodic [But03]. Permeability [HMG01, KS04c]. permutation [MY09, PS10]. Perron [Mor08]. perspective [MH13, MQH14]. perturbation [KS15]. Petersburg [Ano00f]. Phase [AD01, KD04, NPM^{+}06, BC11, Min01, MPC03]. phenomena [EW01]. photo [ONZ99]. photo-neutron [ONZ99]. photon [Sen01]. Photoneutrons [HKHV98]. Piecewise [DMZ03, IM04, Mor98, Mor99, Zal00]. PK [WENG09]. PK/PD [WENG09]. planar [HBA16]. Plasma [BQA03, BS16, CRS14]. Platen [CK18]. Point [GHT00, Smi98, Bea09, BH18, DGKP08, GP12, Har16, Kol20, Nek16, SN13]. Points [Pap04, Nad08a, Ste00]. Poiseuille [BHA18]. Poisson [Bee21, CRS14, GM04, Hau00b, IN17, KRSJ17, TM20]. Policy [BKS06]. Pollard [Vid07]. polynomial [GP12, SS14a, SS17]. polynomials [SH22, Tak96b, Zhe13]. population [AN12, Hei14]. porosity [CL01a]. Porous [KSSV03, KS04c, SM04, BCR11, CL02a, KS04b, KS05, Lej04, SK03, SKL09, Smi98]. Portfolio [MBK06, GG05, Sak10]. positivity [Hal15a]. possible [DK06]. posterior [PP21, Rei20]. potential [CDGG21]. potentials [YJH21]. Power [Hei04, IOR21]. practical [Bou95, Hal08a]. pre [TTEA01]. pre-computed [TTEA01]. Prediction [CHK01, Sei04, But03]. preference [Eğe09]. preferential [Gui08]. Preliminary [BG13]. Prelims [Ano11]. premixed [SH08]. Presence [SSS06, Oga08]. presented [Sab04b]. preserving [Hal15a, Hal16]. preset [Gri17]. Preventive [SC96]. price [HKN12, KSC11]. Pricing [BCZ05, CK18, Sag11, AHT04, BGSR08, DM10, ECLR21, EBSY18, GK08, JWK19, LL21, MH12, MH13, PP05, Sin14]. primitive [Tak96b]. prior [PPN20]. Probabilistic [AH12, Ano99e, Ano00g, Min01, Ökt96, BCR11, ÖG09, PO04, Wag15]. probabilities [AK02, GP19, Hal21, Pöt12, Sab16c]. Probability [KM11b, SK18, CP02, Hal04, Hal05b, Kol20, KS06b, NB19, Nek20]. probability. [Hal05a]. Problem [AS95, BP97, BG01, GHT00, KRSV99, PGB98, SS95, Sim95, CA12, KSNS15, Kol18, MT13, ME09, NVDA07]. Problems [AAD04, MKL01, ST95, BPP01, Gri14, Kas17, LNO15, NÖ09b, PRS05, Rog99, SS02, Sab16a, Sab16b, Sab17, Sab19b, SP20, Sen01, SS18a, SS19a, SS21b]. procedure [BZ20, DKS^{+}98]. procedures [LL11, Voy97]. Process [Ple00, SS95, BEH16, BS16, BL15, ÉM13, Gui08, Hal15b, Hid20, KLW21, KS14, PRS05, SK18]. Processes [Ano00h, DSGZ01, GZ01, KP02, SLP07, AK02, Alf05, Cap01, Ego20, FHS13, FGD13, FG04, Gol03, Gol04, Gri10, Kaw06, KM11a, KT11b, KM11b, Leo06, MS14, NT97, PR19, PP04, Rey17, Rie99, RL18, SK97a, Tur11, Tur19, Wih01]. Processing [DSGZ01]. product [JML20, Xia96]. production [ONZ99, SC96]. Profiles [NPM^{+}06, Eğe09]. Profit [CS96]. Project [But03, Ben16]. projection [IK00, KSNS15, SL10]. projection-statistical [IK00]. proof [KS16, ÖG09]. Propagation [NR02, JLH10]. Properties [SM04, BMO01, Bou95, Xia02]. proposed [BOTAZ19]. Proximal [GHT00]. PSA [MZ98]. Pseudo [GGP06, Sug95, UŠ96, Ant95, MH13, MQH14, Sug04, Tak00]. Pseudo-Random [GGP06, UŠ96, Sug95, Ant95, MH13, MQH14, Sug04, Tak00]. pseudorandom [FT00, Nek16, Yag02, YK08]. Pulse [BP98a]. Pulse-Height-Spectrum [BP98a]. puzzles [MP12].

QMC [AHT04, BM19, SS20b]. quadratic [PP03]. quadrature [VAYT20]. quality [AM17, WLD21]. quantification [Hei14]. quantify [JLH10]. quantiles [MM20]. Quantisation [New01]. quantitative [MQH14]. quantities [eZN22]. Quantization [FS12, BPP01, CP15, PP03, PP05, PRS05, Sag11]. quantization-based [CP15]. Quantum [FGM^{+}01, Hei04]. Quasi [AAD04, Aze12, Bal08, DMZ03, ER06, Hal05b, HPY07, LT04, LM05, MKL01, Pap04, RST96, SS14b, SS19b, Tuf04, AE15, CCMZ08, ELZ11, ELV10, EL18, Hal04, Hal05a, LT08, Leo06, NXÖ18, Owe06, SN13, SK05, SS20b, Hal05a]. quasi-asymptotics [SS20b]. Quasi-Monte [AAD04, Bal08, ER06, HPY07, LT04, LM05, MKL01, RST96, SS14b, SS19b, Tuf04, Hal05a, AE15, CCMZ08, ELZ11, Hal05b, Leo06, NXÖ18, SK05]. quasi-Monte-Carlo [Hal04]. Quasi-probability [Hal05b, Hal04]. Quasi-probability. [Hal05a]. quasi-random [ELV10, EL18, SN13]. quasi-standard [Owe06]. quasi-stochastic [LT08]. quasilinear [GR08]. Quasirandom [KMS04, RKM04, LLM16]. queue [BOTAZ19, SC96]. queue-like [SC96]. Queueing [CS96, BBBR19, Cos01]. queues [TOTAI18].

Radioactive [KPSZ96]. radionuclide [Smi98]. radiosity [CPSH07, SBH04]. radiotherapy [ONZ99]. Raikov [Fuk96]. Random [AW10, AE15, GGP06, Hau00b, Hor02, KS06c, Oga96, RKM04, ST95, SS95, SS02, SS03, SSL06, SM09, Sab16a, Sab16b, Sab17, Sab19b, SS18a, SS20a, SS21b, SM04, Sim18, SS07, SVH^{+}04, ST00, Tak97, UŠ96, Wag10, ASTY19, AM15, Ant95, BK95, CL18, Ego20, ELV10, ES11, EL18, Gri14, Ima13, KM22, KM11b, KKS13, KS06b, LP11, LP13, Lev16, Mak15, MH12, MH13, MQH14, MS16, MR04, Nad08a, Nek20, Oga97, PMW10, Rie99, RV99, SK97a, SSL04, Sab08, SKL09, SL14, Sab16c, Sab19a, SS21a, SSDT21, SN13, SS14a, SS17, SS19a, SSG99, SM03, Ste00, Sug95, Sug04, Tak00, Tur19, Yag00]. random-bit [Nek20]. Randomization [SM09, Tuf04, EL18, KLP14, Kol21]. Randomized [HPY07, BK95, CCMZ08]. Randomizers [FGM^{+}01]. Randomness [Yag00, ASTY19]. Range [VA04, BL15]. ranges [SSG99]. rank [GP12, LL20]. ranked [AN12]. Rapid [HMG01]. Rare [MS14, FGM17]. rarely [eZN22]. Rate [BT96, CP01, KP02, BH18, Gol04, KHO97, LCRK18, PP19]. Ratio [SSS06, MM20, SD96]. Ray [BP97, BP98b]. rays [SAKG15]. reaction [SLK15, Sab17, SK18, Sab19b]. reaction-diffusion [SLK15]. reactions [BC11]. Reactor [HKHV98]. Real [Oga08, TTEA01, OW07]. Real-time [Oga08, OW07]. Realizability [Hei08]. reciprocal [Tak97]. recombination [SS21b]. recommendations [Bou95]. recovering [KSNS15]. rectangles [Sab19b]. rectangular [DM10]. Recursive [Cap01, PR19, SH22, FS12, PW01]. Reduction [Aro04, Kaw07, NAKS04, Bee21, BOTAZ19, Hei95, KD99, KD04, KS03, MP02, TOTAI18, ZCC04, Cos01]. Reflected [Hau00b, HKHV98, BST10, CLP17, Gob01, Yan13]. Reflecting [KS00, Wel06]. Reflection [Hau00c]. Reflections [DK98b]. regime [Aze12, EBSY18]. regions [DM10]. Regression [SSS06, BG13, CCG15, FGM17, LL21, PP21, SH22, WN19, Zal00]. regular [GLP17]. regularization [Ant11]. Rejection [LH04, Voy98, Nk16]. Relative [Kur95b, KS95b, Kur95a, Kur97, KOSY01, TOTAI18]. relaxation [Zal00]. Reliability [KD04, KM15, KD99, MZB04, NK06, RL18, Tur19, ZCC04]. Reliable [Pap98, JML20]. Remarks [EL18, Pag07]. Reneging [CS96]. Repetition [GGP06]. replica [Ari15, LLLP12]. Replication [Kel04]. replications [Sak10]. representation [DMR16, IOR21, LOR18]. representations [MT08]. repulsion [CL01b]. Resampling [MBK06]. Research [HKHV98]. reservoir [Lej03]. resources [But03]. respect [EZ04]. restart [TM20]. restarted [MP12]. Restricted [Kel04, Man03]. Result [Ökt96, ÖG09]. Results [KSK97, AP04, TTEA01]. retrial [TOTAI18]. retrospective [JS07, JS10]. reuse [CPSH07]. Reusing [SBH04]. reversion [IP17]. Reverting [CP01]. Review [Kra01, MQH14]. Revisited [PR19]. Reynolds [Kur97]. rho [Vid07]. Richardson [Pag07]. Riesz [Fuk96]. ring [FKM08]. risk [AK02, DL14, FGM17, Sak10]. RJMCMC [DSGZ01]. Robin [Sab16a]. robust [AN12, ST00]. Robustness [Oga96, Oga97]. Romberg [Pag07]. Rotation [Sug95, MP02]. roughness [KLR^{+}03]. ROW [KM95]. ROW-Type [KM95]. ruin [AK02]. rule [HR02, Yag00]. rules [Eğe09].

Sample [SS97, NB19, RS19, UV00]. sample-mean [RS19]. samples [FGD13, Gri10]. Sampling [CRGF18, LS97, Row00, SLP07, Sta95, AN12, BFP09, CRT02, CP15, FGD13, FS12, JLH10, Kaw06, KM11a, KS16, Leo06, MS14, ME09, Nin03, Shv03, Spa21, ST00, ZCC04]. Santalo [ES20]. scalar [BJ01]. Scale [SVH^{+}04, Hei08, Kaw07, SH22]. Scheme [BT96, Hau00c, KM95, AJC16, Bab99, BBG15, Buc04, CLP17, Hal15b, IIO14, KHO97, NT21, OW07, Oga08, Rey17, Rie99, Wel06, Yan13]. Schemes [BF01, Vid07, Alf05, EW02, Gob01, Hal15a, Hal16, KM02, MT08, MPC03]. Scholes [Sin14]. Schrödinger [Wag15, dBDD01]. science [SK15]. Scrambled [MC04, MY09, WLD21]. scrambling [AM17]. SDE [KHO97, Mar10]. SDEs [KS06a, MS16, NY19a, NZ09, OY19, YY18, Yam21]. search [EHE18, Har16]. Second [Ano96d, MPC03, NY19a, SS21a, SSDT21, YY18]. second-order [MPC03, NY19a, YY18]. section [Ant15]. Security [Sug04, JWK19]. Selected [DS10]. Selection [Sab04b, BZ20, LLM16, Lin06, RV99]. self [Hei14]. self-similar [Hei14]. semi [HS22, IIO14, IK00, Lej01, Sab16a]. semi-cylinders [Sab16a]. semi-discrete [HS22]. semi-linear [Lej01]. semi-static [IIO14]. semi-statistical [IK00]. semiclassical [NMH04]. Semiconductor [BAO^{+}04]. semiconductors [KSPZ20]. semilinear [IOR21, LOR18]. Seminar [Ano00a, Ano02e, Ano03b, DS10, Sab04b]. semipermeable [DMR16]. sensitivities [PWY99]. Sensitivity [GP19, SSDM21, SS07, CCMZ08, DTS22, KSC11, Kol18, Kol21, MM00, PPN20, SK05, SM08]. Separable [Row00]. Separation [Row02]. September [Ano02e]. Sequence [MC04, Ökt96, BM19, FIN02, FKM08, NEBW20, ÖG09]. Sequences [Ant96, RKM04, AH12, Chi13, DTS22, FL10, Har19, IM04, MY09, Mor98, Mor99, Mor02, Mor04, Mor05, Mor08, MM12, Nk16, PC04, PÖ20, RST96, SN13, Tak96a, Tak97, Tuf96, Tuf98, Xia96]. Sequential [Hal06, Hal08a, LS97]. set [AN12]. sets [Bea09, DGKP08, GP12, Har16, Kol20]. setting [NÖ09b]. Seventh [DS10]. shaped [SL14]. Sharp [CP02, TM20]. sheath [CRS14]. shift [Bou95]. shifted [Gol04]. Shock [DK98b]. Short [VA04]. Si [RBB21]. Sigma [Hal08b]. Sigma-algebra [Hal08b]. sign [Erm11]. sign-changing [Erm11]. Signals [AD01]. Significant [Row03]. similar [Hei14]. simple [VAYT20, Cos01]. simplest [Erm11]. simplex [PC04]. Simulating [BBG15, Hau00c, Lej03, LN04]. Simulation [AK02, Ano96d, BQA03, BP97, Bou05, Gui99, Hau00b, Hor02, KS00, KM22, KPV18, Kra01, LT04, Mak15, ONZ99, Ple00, PMW10, SA96, SLP07, Tur11, VA04, WK05, ATBM14, AP04, ABKT18, Ave04, BHA18, BS16, BS18, BOTAZ19, But03, CJV16, ÉM13, FN09, FG04, Hau00a, KSPZ20, Khi00, KS04b, KS05, KS15, LP11, LP13, LCRK18, LT08, Leo06, Lev16, MG10, MR04, MS14, Min01, Nek20, NMH04, PIR04, PP04, Raj19, Rog96, SRKL96, SK97a, SK03, SKL09, SLK15, SE18, Sak10, Smi98, SH08, SZKS21, TOTAI18, Tur19, YY18, mMSD04, Ano00f, Mis07]. Simulations [BAO^{+}04, NPM^{+}06, ZPK02, MT08, RBB21]. single [Man03]. singular [BCR11]. singularities [Sim18]. Sintering [WK05]. six [SD96]. size [DGKP08, ES20]. skew [DMR16, Osa01]. Skewed [Nad08b]. skin [MM00]. slip [HBA16]. small [ASTY19, DGKP08, KS15, NT97, SS19a, SM08]. Smolouchovsky [SRKL96]. Smoluchowski [Bab99, DT01, GZ01, Gui97, KW97, KS01, KS03, SK97a, SLK15]. smooth [AD99]. Smoothed [LH04, Cap01]. Sobol [Har19]. Sobol′ [MY09]. social [Eğe09]. software [NK06]. SOI [VA04]. sojourn [Tak96a]. Solid [NPM^{+}06]. Solutes [SVH^{+}04]. Solution [AS95, GN99, KNS04, Lik98, Rog99, BJ01, CRS14, EM17, Gri14, KSNS15, Lej01, MK06, PS98, RJ20, SS02, Xia96]. Solutions [DT01, Kan95, NAKS04, BCR11, EZ04, Ego07, ES10, Rot07, Zhe13]. Solve [WK05]. Solving [Hal06, ER06, EP19, Gol03, KS15, LL13, MP12, PS05, Rie99, SRKL96, SM12, Sab16b, Sab17, Sab19b, SP20, SS14a, SS18a, SS19a, SS20a, SS21b, SS19b]. Some [AP04, BMO01, Kra01, MT08, Nao95, Xia02, Khi00, NT21, Sab19a, Xia96, eZN22, Zal00]. Source [Row02]. Space [BQA03, KD04, KNS04, BJ22, Dic06, EM03, KM11b, KM15, PS98]. Space-dependent [KNS04]. Sparsified [SM09]. spatial [Kol20]. spatially [KS01]. SPDEs [Oga01]. Special [LLM16]. Spectra [Mor08]. Spectral [ELRU04, KS06c, NS09, BK95, GM04, Gri10, SM12, SL14]. spectral-based [Gri10]. Spectrum [BP98a, Nak98]. Speed [LK02, Kab05]. sphere [CL18, SK18]. Spheres [ST95, SS95, SS02, SS03, SSL06, Sab16b, Sab17, Sab19a, SS18a, SS19a, SS21b]. spherical [Gol04, SSL04]. spline [PPN20]. Split [NAKS04]. Splitting [Kel04, KD04, Sab16c, Sta95]. spot [NO09a]. sputtering [BS16, RBB21]. Square [NPM^{+}06, HBBA15]. Square-Wave [NPM^{+}06]. squared [Alf05]. St [Ano00f]. stability [HS22]. Stable [KM95, KS04c, KM11a]. stage [MS14, PP19]. standard [Owe06, PIR04]. star [DK06, Sha10]. state [BJ22, FN09, NB19, PIR04, SS21b, eZN22]. state-space [BJ22]. States [GZ01]. static [IIO14]. stationary [FGD13, PGS09, Rog99]. Statistical [Kol20, Ave04, IK00, Kol21, Rog96]. Statistically [KSSV03, Hal04, Hal05a, Hal05b]. Statistics [FGM^{+}01, Bea09, BBBR19, VMS08]. steady [FN09, NB19, PIR04, SS21b, eZN22]. steady-state [FN09, NB19, PIR04, SS21b, eZN22]. step [FP02, Pag07]. Stochastic [AS95, Ano96d, BT96, BF01, CK04, EW01, FP02, GN99, GHT00, Hau00b, Hau00c, Kan95, Kas17, Kaw07, KS01, KS03, KS04b, KS05, KS15, KM95, KS95b, Kur95a, Kur97, KSK97, KSSV03, KLR^{+}03, KS06c, LP12, LN04, NAKS04, NHD06, PGB98, SRKL96, SK97b, SK98, SS01, SK03, SKL09, SL10, SM12, SLK15, SS14a, SS17, Sim95, WK05, Zal00, AG03, BMO01, BPP01, BFP09, BGSR08, BMS09, BEH16, BH18, BFP97, BJ01, Buc04, DTS22, EZ04, Ego07, ES10, EM17, EP19, ÉM13, FP99, GG05, GA99, GR08, Hab12, HS22, Hei08, Hid20, KSNS15, Kol18, KM02, KS14, KW02, KOSY01, LCRK18, LLM16, LT08, MH13, MPC03, MK06, NY19b, NT21, NP04a, OO03, PG19, Pit06, PS05, Pri01]. stochastic [Rot07, RL18, Sab16c, Sab16d, SSDM21, SH22, SZKS21, Voy97, Wel06, Yan13, Zhe13, dBDD01, Ano00h]. Stokes [Sim95]. stopped [BST10]. stopping [Kas17, PRS05]. Strang [Voy97]. Strategies [SS97]. strategy [IIO14]. Stratified [Leo06, SLP07, CP15]. stress [Hei08]. Strong [AJC16, BH18, BLNSP06, KS00, CL01b]. structure [Ave04, WLD21, Wih01]. structures [LLM16]. student [MC20, Nad08a, Nad08b]. Study [BS16, SSS06, DTS22, JLH10, Raj19, SNDS14, Sin14]. studying [EW01]. Style [KS04a]. sub [Tur11]. sub-Gaussian [Tur11]. Subdiffusion [CK04]. Subdomains [HTKM19]. Subgrid [KS04c, Hei08]. subgrid-scale [Hei08]. subject [CA12]. substitution [FVK16, FVK17]. substitution-transposition [FVK17]. Substrates [NPM^{+}06]. sudoku [MP12, LW10]. sulfide [SZKS21]. sum [ABKT18]. summary [Hal08a]. sums [Fuk96, KM11b, KS06b]. supercomputing [AM15]. Superdiffusion [CK04]. Surface [NPM^{+}06, KLR^{+}03, Smi98, YJH21]. Survey [Tuf04]. Surveys [SS97]. survival [LL20, Sab16c]. switching [EBSY18, LNO15]. symmetric [BL15, Osa01]. synchrony [Row03]. System [MZB04, PGB98, CDGG21, Hab12, Mor04, MM12, RL18, SC96]. systematic [JLH10]. Systems [Hal06, KD04, Kra01, Lik98, NR02, Oll01, Pap98, Ant11, Ave04, Hal08a, IOR21, KD99, Nek03, SM09, SL10, Sab16d]. Systolic [Lik98].

Tagged [Osa01]. tail [ABKT18, ZZA21]. tailed [ZZA21]. takeovers [HR02]. taking [EM03]. tangent [ES17]. tau [KT11a]. tau-leap [KT11a]. Taylor [Dic06]. teaching [MC20]. Technique [Aro04, MPZP04, Pap98, Ant15, KS15, MM00, MP02]. Techniques [Ars98, Ars07, Hal06, AHT04, BN15]. temperature [MK06]. tempered [KM11a]. Tensor [Nak98]. term [Buc04, IP17]. terminal [MS16]. Test [ELRU04, GGP06, AW10, LL20, Man03, MH12, NS09, Tak96a]. tested [BOTAZ19]. Testing [FGM^{+}01, IP17, KS14, TOTAI18]. tests [Hab11, Tak97]. their [Hal04, Hal05a, Hal05b]. Theis [Aze12]. theorem [FGD13, NO09a, SS15, Wel06, Gol03]. theorems [BK14, GLP17, Hal08b, KKS13]. Theoretical [dBDD01, Min01, PC04]. Theory [Hau00c, Com01, Cos01]. thermodynamic [SE18]. thermostatic [IOR21]. thin [BS18, RBB21]. third [NY19b, Rey17]. third-order [NY19b]. Thouless [HK14]. three [CRS14, Kol20, LW10, Mor05, SS97]. three-dimensional [CRS14, Kol20, Mor05]. Threshold [Vid07]. Time [Hau00c, Nad07, Nak98, BJ22, BH18, BL15, CP02, ÉM13, Gui08, Hal21, IP17, Kaw07, Khi00, MS16, NÖ09b, OW07, Oga08, PP19, PPN20, Pri01, SK18, Shv03, SH22, Tak96a, TTEA01, Yam21]. time-dependent [CP02, NÖ09b, PP19]. time-inhomogeneous [Yam21]. Time-to-Event [Nad07, PPN20]. times [BEH16, FHS13, JWK19]. Tossing [NP04b]. total [Rey17]. Touching [Rie99]. tracking [SP20]. Tractability [NP04b]. trajectories [SP20]. Trajectory [Kel04, MP02]. transform [Fuk96, Ima13]. transformation [Kaw06, TTEA01]. transformations [IM04]. Transformed [LH04]. transforming [PC04]. transient [Aze12, Sab17, SK18, Sab19a, Sab19b, SS19a, SS21b]. transition [DMR16, HK14]. Transport [Ano00h, BP98b, CK04, Hor02, KSSV03, LS97, SVH^{+}04, KW02, PGS09, PIR04, SS01, SKL09, Sen01, SAKG15, Smi98, SS18b]. transposition [FVK16, FVK17]. Trials [Nad07]. triangular [HK14]. triangular-lattice [HK14]. Turbulence [Kur95a, Kur97, Nak98, SK98, Nak97]. Turbulent [Ano00h, Kra01, Kur95b, KS95b, KSK97, SK97b, SK98, Min01, SK00, SS01, SH08]. Two [Kaw07, Kur95b, Kur97, KSK97, NP04a, SK97b, Sim95, DMR16, Hal15a, HBBA15, KOSY01, LL20, MS14, Min01, MPC03, Mor02, Mor08, RS21, SH22]. Two-dimensional [Sim95, HBBA15]. two-factor [Hal15a]. Two-Particle [Kur95b, Kur97, KSK97, SK97b]. two-phase [Min01, MPC03]. two-stage [MS14]. Two-time-scale [Kaw07, SH22]. Type [KM95, AK02, BCR11, KW97, KM02, LOR18, Nek03].

Uhlenbeck [KM11a]. ultra [KK09]. Unbiased [BJ22, RJ20, SS97, SD96]. Uncertainty [Hei14, JLH10, mMSD04]. Unconfined [KS04b]. unconstrained [BFP09]. Understanding [BS18]. Uniform [Eğe09, SUZ04, SN13, Ste00]. uniformly [FKM08]. unknown [BEH16, ÉM13]. Unrestricted [Man03]. updating [MZ98]. upon [SZKS21]. Upper [KS04a, BDGZ20, DK06]. Usage [UV00]. use [Bou95, BN15, IP17, TOTAI18, Tuf96, Tuf98, VMS08]. used [Mar10]. Using [BAO^{+}04, KS00, KNS04, LTD01, LK02, SVH^{+}04, Voy97, AN12, BCZ05, BFP09, BS18, BOTAZ19, But03, CRT02, Cap01, CLP17, CK18, ECLR21, Ego20, ELZ11, EBSY18, FN09, Hau00c, JWK19, KD04, LCRK18, Mat99, Mis07, NEBW20, NY19a, PÖ20, RBB21, Row02, Row03, Tur11, YY18].

Vacuum [Ple00]. valid [Hal04, Hal05a, Hal05b]. Validation [CA12]. Value [MKL01, ST95, Sim95, NVDA07, Rog99]. Valued [Hei04, Gri17, Mal07]. values [EM03]. VaR [BFP09]. variables [ASTY19, Nad08a, Pöt12, SM03]. Variance [Aro04, CP01, Erm11, Hei95, Kaw07, NAKS04, Pag07, Bee21, BOTAZ19, GK08, KD99, KD04, KS03, MP02, Rie99, TOTAI18, ZCC04, Cos01]. variant [JML20]. variants [NP04a]. variate [OY19]. variates [ABKT18, BG13, ECLR21, Mak15]. Variation [Xia96, KM11a, Rey17]. Vector [Hei04, KS06c, Sab16d, Gri17]. vector-valued [Gri17]. Velocity [Nak98, KSPZ20, Nak97]. verification [Ant15, JML20]. version [AJC16, SM09]. versus [IP17]. via [CCMZ08, Hei08, Ima13, MY09, MS14, OY19]. Victoir [AJC16]. view [Com01, KT11b, PC04]. vis [Sin14]. vis-à-vis [Sin14]. Viscoelastic [BP02]. viscous [BJ01]. volatility [BMS09, GG05, LCRK18, MH13, NO09a, OW07, Oga08]. volume [BFP97]. volumes [RS19, RS21]. vs [Man03].

WAFOM [Har16]. Walk [HTKM19, ST95, SS95, SS02, SS03, SSL06, SM09, SM04, SVH^{+}04, ELV10, ES11, Rie99, SSL04, Sab16a, Sab16b, Sab16c, Sab17, Sab19a, Sab19b, SS21a, SSDT21, SS18a, SS19a, SS20a, SS21b, Sim18, Tak97]. Walk-on-Subdomains [HTKM19]. Walks [KMS04, RKM04]. Warnock [Owe06]. Water [MPZP04]. Wave [NPM^{+}06, EW02, KSNS15]. Wavelet [KS06c, Nao95, SUZ04, Tur19]. Wavelet-based [Tur19]. wavelets [Tur11]. Waves [DK98b]. Weak [KHO97, KM95, KP02, Lej01, MPC03, Rot07, BST10, Gob01, KT11a, KSC11, NY19b, OY19, YY18, Yam21, CP02]. Weather [EBSY18]. Weibull [NK06]. Weight [MZ98, Tak96a]. Weighted [PIR04, FP02, GLP17, GK08, KS16, LL20]. well [SS01]. well-mixed [SS01]. Weyl [Fuk96, ST00]. White [PP04, GR08]. WIAS [Ano00h, Ano02e]. Widening [BN15]. Width [VA04]. Wiener [Ego20]. Wigner [KNS04, NAKS04, SNDS14]. wise [ASTY19]. within [PIR04]. without [CL02a, FGM^{+}01, FM01, YY18]. Workshop [Ano96d, Ano00f, Ano00h]. world [Hei14]. Worst [RJG13].

Zakai [RJ20]. zero [BH18, ÉM13, IN17, Rie99]. zero-inflated [IN17]. zerovariance [Rie99]. ziggurat [NXÖ18].

References

Alexandrov:2004:PQM

[AAD04] V. Alexandrov, E. Atanassov, and I. Dimov. Parallel quasi-Monte Carlo methods for linear algebra problems. Monte Carlo Methods and Applications, 10(3–4):213–219, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.213/mcma.2004.10.3-4.213.xml.

Alouini:2018:ESL

[ABKT18] Mohamed-Slim Alouini, Nadhir Ben Rached, Abla Kammoun, and Raul Tempone. On the efficient simulation of the lefttail of the sum of correlated log-normal variates. Monte Carlo Methods and Applications, 24(2):101–115, June 2018. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-2/mcma-2018-0009/mcma-2018-0009.xml.

Atanassov:1999:NOM

[AD99] Emanouil I. Atanassov and Ivan T. Dimov. A new optimal Monte Carlo method for calculating integrals of smooth functions. Monte Carlo Methods and Applications, 5(2):149–167, ???? 1999. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
2/mcma.1999.5.2.149/mcma.1999.5.2.149.xml.

Andrieu:2001:OEA

[AD01] Christophe Andrieu and Arnaud Doucet. Optimal estimation of amplitude and phase modulated signals. Monte Carlo Methods and Applications, 7(1–2):1–14, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.1/mcma.2001.7.1-2.1.xml. Monte Carlo

and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Antonov:2015:RCQ

[AE15] Anton A. Antonov and Sergej M. Ermakov. Random cubatures and quasi-Monte Carlo methods. Monte Carlo Methods and Applications, 21(3):179–??, September 2015. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-3/mcma-2015-0102/mcma-2015-0102.xml.

Anh:2003:FSE

[AG03] V. V. Anh and W. Grecksch. A fractional stochastic evolution equation driven by fractional Brownian motion. Monte Carlo Methods and Applications, 9(3):189–199, ???? 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
3/156939603322728969/156939603322728969.xml.

Aistleitner:2012:PEB

[AH12] Christoph Aistleitner and Markus Hofer. Probabilistic error bounds for the discrepancy of mixed sequences. Monte Carlo Methods and Applications, 18(2):181–??, June 2012. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-2/mcma-2012-0006/mcma-2012-0006.xml.

Albrecher:2004:QTC

[AHT04] Hansjörg Albrecher, Jürgen Hartinger, and Robert F. Tichy. QMC techniques for CAT bond pricing. Monte Carlo Methods and Applications, 10(3–4):197–211, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.197/mcma.2004.10.3-4.197.xml.

AlGerbi:2016:NVS

[AJC16] Anis Al Gerbi, Benjamin Jourdain, and Emmanuelle Clément. Ninomiya–Victoir scheme: Strong convergence, antithetic version and application to multilevel estimators. Monte Carlo Methods and Applications, 22(3):197–??, September 2016. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-3/mcma-2016-0109/mcma-2016-0109.xml.

Albrecher:2002:SRP

[AK02] Hansjörg Albrecher and Josef Kantor. Simulation of ruin probabilities for risk processes of Markovian type. Monte Carlo Methods and Applications, 8(2):111–127, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
2/mcma.2002.8.2.111/mcma.2002.8.2.111.xml.

Alfonsi:2005:DSC

[Alf05] Aurélien Alfonsi. On the discretization schemes for the CIR (and Bessel squared) processes. Monte Carlo Methods and Applications, 11(4):355–384, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2005.11.issue-4/156939605777438569/156939605777438569.
xml.

Andersen:2015:MEL

[AM15] Timothy D. Andersen and Michael Mascagni. Memory efficient lagged-Fibonacci random number generators for GPU supercomputing. Monte Carlo Methods and Applications, 21(2):163–174, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2015.21.issue-2/mcma-2014-0017/mcma-2014-0017.xml.

Aljahdali:2017:FIS

[AM17] Asia Aljahdali and Michael Mascagni. Feistel-inspired scrambling improves the quality of linear congruential generators. Monte Carlo Methods and Applications, 23(2):89–??, June 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-2/mcma-2017-0105/mcma-2017-0105.xml.

Al-Nasser:2012:PME

[AN12] Amer Ibrahim Al-Omari Amjad D. Al-Nasser. On the population median estimation using robust extreme ranked set sampling. Monte Carlo Methods and Applications, 18(2):109–??, June 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-2/mcma-2012-0002/mcma-2012-0002.xml.

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 2(2):172–??, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1996.2.issue-2/mcma.1996.2.2.172/mcma.1996.
2.2.172.xml.

Anonymous:1996:EBb (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1996.2.issue-3/mcma.1996.2.3.251/mcma.1996.
2.3.251.xml.

Anonymous:1996:EBc

[Ano96c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 2(4):343–??, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1996.2.issue-4/mcma.1996.2.4.343/mcma.1996.
2.4.343.xml.

Anonymous:1996:SIW

[Ano96d] Anonymous. Second international workshop on mathematical methods in stochastic simulation and experimental design. Monte Carlo Methods and Applications, 2(1):89–??, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
1/mcma.1996.2.1.89/mcma.1996.2.1.89.xml.

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 3(1):85–??, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1997.3.issue-1/mcma.1997.3.1.
85/mcma.1997.3.1.85.xml.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 3(2):167–??, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1997.3.issue-2/mcma.1997.3.2.167/mcma.1997.
3.2.167.xml.

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 3(3):251–??, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1997.3.issue-3/mcma.1997.3.3.251/mcma.1997.
3.3.251.xml.

Anonymous:1997:EBd

[Ano97d] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 3(4):347–??, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1997.3.issue-4/mcma.1997.3.4.347/mcma.1997.
3.4.347.xml.

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 4(1):91–??, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1998.4.issue-1/mcma.1998.4.1.
91/mcma.1998.4.1.91.xml.

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 4(2):177–??, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1998.4.issue-2/mcma.1998.4.2.177/mcma.1998.
4.2.177.xml.

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 4(3):285–??, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1998.4.issue-3/mcma.1998.4.3.285/mcma.1998.
4.3.285.xml.

Anonymous:1998:EBd

[Ano98d] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 4(4):375–??, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1998.4.issue-4/mcma.1998.4.4.375/mcma.1998.
4.4.375.xml.

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 5(1):81–??, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1999.5.issue-1/mcma.1999.5.1.
81/mcma.1999.5.1.81.xml.

Anonymous:1999:EBb (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1999.5.issue-2/mcma.1999.5.2.189/mcma.1999.
5.2.189.xml.

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 5(3):283–??, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1999.5.issue-3/mcma.1999.5.3.283/mcma.1999.
5.3.283.xml.

Anonymous:1999:EBd

[Ano99d] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 5(4):375–??, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1999.5.issue-4/mcma.1999.5.4.375/mcma.1999.
5.4.375.xml.

Anonymous:1999:ICM

[Ano99e] Anonymous. International Conference on Monte Carlo and Probabilistic Methods for Partial Differential Equations July, 3– 5, 2000 / Monaco, first announcement. Monte Carlo Methods and Applications, 5(3):281–282, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
3/mcma.1999.5.3.281/mcma.1999.5.3.281.xml.

Anonymous:2000:ISM

[Ano00a] Anonymous. 3rd IMACS Seminar on Monte Carlo Methods (MCM2001: First announcement). Monte Carlo Methods and Applications, 6(3):255–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2000.6.issue-3/mcma.2000.6.3.
255/mcma.2000.6.3.255.xml.

Anonymous:2000:EBa

[Ano00b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 6(1):77–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2000.6.issue-1/mcma.2000.6.1.
77/mcma.2000.6.1.77.xml.

Anonymous:2000:EBb

[Ano00c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 6(2):163–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2000.6.issue-2/mcma.2000.6.2.163/mcma.2000.
6.2.163.xml.

Anonymous:2000:EBc

[Ano00d] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 6(3):259–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2000.6.issue-3/mcma.2000.6.3.259/mcma.2000.
6.3.259.xml.

Anonymous:2000:EBd

[Ano00e] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 6(4):361–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2000.6.issue-4/mcma.2000.6.4.361/mcma.2000.
6.4.361.xml.

Anonymous:2000:FSP

[Ano00f] Anonymous. Fourth St. Petersburg Workshop on Simulation. Monte Carlo Methods and Applications, 6(3):257–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2000.6.
issue-3/mcma.2000.6.3.257/mcma.2000.6.3.257.xml.

Anonymous:2000:MCI

[Ano00g] Anonymous. Monte Carlo 2000: International Conference on Monte Carlo and Probabilistic Methods for Partial Differential Equations, Monte Carlo (Monaco): July 3–5, 2000. Monte Carlo Methods and Applications, 6(2):159–??, ???? 2000. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
2/mcma.2000.6.2.159/mcma.2000.6.2.159.xml.

Anonymous:2000:WSM

[Ano00h] Anonymous. Workshop: Stochastic Models for Turbulent Transport Processes, WIAS, Berlin, 23–25 October 2000. The abstracts. Monte Carlo Methods and Applications, 6(4):349–??, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2000.6.
issue-4/mcma.2000.6.4.349/mcma.2000.6.4.349.xml.

Anonymous:2001:EBa

[Ano01a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 7(1–2):213–??, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.213/mcma.2001.7.1-2.213.xml.

Anonymous:2001:EBb

[Ano01b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 7(3–4):421–??, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-3-4/mcma.2001.7.
3-4.421/mcma.2001.7.3-4.421.xml.

Anonymous:2002:EBa

[Ano02a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 8(1):107–??, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-1/mcma.2002.8.1.107/mcma.2002.
8.1.107.xml.

Anonymous:2002:EBb

[Ano02b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 8(2):217–??, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-2/mcma.2002.8.2.217/mcma.2002.
8.2.217.xml.

Anonymous:2002:EBc

[Ano02c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 8(3):317–??, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-3/mcma.2002.8.3.317/mcma.2002.
8.3.317.xml.

Anonymous:2002:EBd

[Ano02d] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 8(4):421–??, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-4/mcma.2002.8.4.421/mcma.2002.
8.4.421.xml.

Anonymous:2002:MII

[Ano02e] Anonymous. MCM2003: IV IMACS Monte Carlo Seminar WIAS, Berlin, 15–19 September 2003. Monte Carlo Methods and Applications, 8(2):215–??, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-2/mcma.2002.8.2.215/mcma.2002.
8.2.215.xml.

Anonymous:2003:EB

[Ano03a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 9(3):291–??, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2003.9.issue-3/156939603322729030/156939603322729030.
xml.

Anonymous:2003:IIS

[Ano03b] Anonymous. IVth IMACS Seminar on Monte Carlo Methods. Monte Carlo Methods and Applications, 9(1):87–??, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2003.9.
issue-1/156939603322587498/156939603322587498.xml.

Anonymous:2004:EB

[Ano04] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 10(2):179–??, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2004.10.issue-2/156939604777303262/
156939604777303262.xml.

Anonymous:2005:EBa

[Ano05a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 11(1):93–??, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2005.11.issue-1/1569396054027247/1569396054027247.
xml.

Anonymous:2005:EBb

[Ano05b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 11(2):199–??, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2005.11.issue-2/156939605777585953/
156939605777585953.xml.

Anonymous:2005:EBc

[Ano05c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 11(4):465–??, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2005.11.issue-4/156939605777438523/
156939605777438523.xml.

Anonymous:2006:EBa

[Ano06a] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 12(1):95–??, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2006.12.issue-1/156939606776886661/156939606776886661.
xml.

Anonymous:2006:EBb

[Ano06b] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 12(2):187–??, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2006.12.issue-2/156939606777488851/
156939606777488851.xml.

Anonymous:2006:EBc

[Ano06c] Anonymous. Editorial Board. Monte Carlo Methods and Applications, 12(3–4):343–??, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2006.12.issue-3/156939606778705173/
156939606778705173.xml.

Anonymous:2011:P

[Ano11] Anonymous. Prelims. Monte Carlo Methods and Applications, 17(4):i–??, December 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2011.17.issue-4/mcma.2011.prelims4/mcma.2011.
prelims4.xml.

Anonymous:2012:Ma

[Ano12a] Anonymous. Masthead. Monte Carlo Methods and Applications, 18 (1):i–??, March 2012. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2012.18.issue-1/mcma-2012-masthead1/mcma-2012-masthead1.
xml.

Anonymous:2012:Mb

[Ano12b] Anonymous. Masthead. Monte Carlo Methods and Applications, 18 (2):i–??, June 2012. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2012.18.issue-2/mcma-2012-masthead2/mcma-2012-masthead2.
xml.

Anonymous:2012:Mc

[Ano12c] Anonymous. Masthead. Monte Carlo Methods and Applications, 18(3):i–??, September 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2012.18.issue-3/mcma-2012-masthead3/mcma-2012-
masthead3.xml.

Anonymous:2012:Md

[Ano12d] Anonymous. Masthead. Monte Carlo Methods and Applications, 18(4):i–??, December 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2012.18.issue-4/mcma-2012-masthead4/mcma-2012-
masthead4.xml.

Anonymous:2013:Ma

[Ano13a] Anonymous. Masthead. Monte Carlo Methods and Applications, 19 (1):i–??, March 2013. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2013.19.issue-1/mcma-2013-masthead1/mcma-2013-masthead1.
xml.

Anonymous:2013:Mb

[Ano13b] Anonymous. Masthead. Monte Carlo Methods and Applications, 19(2):i–??, July 2013. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2013.19.issue-2/mcma-2013-masthead2/mcma-2013-masthead2.
xml.

Anonymous:2013:Mc

[Ano13c] Anonymous. Masthead. Monte Carlo Methods and Applications, 19(3):i–??, September 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2013.19.issue-3/mcma-2013-masthead3/mcma-2013-
masthead3.xml.

Anonymous:2013:Md

[Ano13d] Anonymous. Masthead. Monte Carlo Methods and Applications, 19(4):i–??, December 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2013.19.issue-2/mcma-2013-masthead2/mcma-2013-
masthead2.xml.

Anonymous:2014:Fa

[Ano14a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 20(1):i–??, March 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2014.20.issue-1/mcma-2014-frontmatter1/
mcma-2014-frontmatter1.xml.

Anonymous:2014:Fb

[Ano14b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 20(2):i–??, June 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2014.20.issue-2/mcma-2014-frontmatter2/
mcma-2014-frontmatter2.xml.

Anonymous:2014:Fc

[Ano14c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 20(3):i–??, September 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2014.20.issue-3/mcma-2014-frontmatter3/
mcma-2014-frontmatter3.xml.

Anonymous:2014:Fd

[Ano14d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 20(4):i–??, December 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2014.20.issue-4/mcma-2014-frontmatter4/
mcma-2014-frontmatter4.xml.

Anonymous:2014:M

[Ano14e] Anonymous. Masthead. Monte Carlo Methods and Applications, 20(1):i–??, January 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2013.19.issue-3/mcma-2013-masthead3/mcma-2013-
masthead3.xml.

Anonymous:2015:Fa

[Ano15a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 21(1):i–??, March 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2015.21.issue-1/mcma-2015-frontmatter1/
mcma-2015-frontmatter1.xml.

Anonymous:2015:Fb

[Ano15b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 21(2):i–??, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2015.21.issue-2/mcma-2015-frontmatter2/
mcma-2015-frontmatter2.xml.

Anonymous:2015:Fc

[Ano15c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 21(3):i–??, September 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2015.21.issue-3/mcma-2015-frontmatter3/
mcma-2015-frontmatter3.xml.

Anonymous:2015:F

[Ano15d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 21(4):i–??, December 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2015.21.issue-4/mcma-2015-frontmatter4/
mcma-2015-frontmatter4.xml.

Anonymous:2016:Fa

[Ano16a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 22(1):i–??, March 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2016.22.issue-1/mcma-2016-frontmatter1/
mcma-2016-frontmatter1.xml.

Anonymous:2016:Fb

[Ano16b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 22(2):i–??, June 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2016.22.issue-2/mcma-2016-frontmatter2/
mcma-2016-frontmatter2.xml.

Anonymous:2016:Fc

[Ano16c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 22(3):i–??, September 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2016.22.issue-3/mcma-2016-frontmatter3/
mcma-2016-frontmatter3.xml.

Anonymous:2016:F

[Ano16d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 22(4):i–??, December 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2016.22.issue-4/mcma-2016-frontmatter4/
mcma-2016-frontmatter4.xml.

Anonymous:2017:Fa

[Ano17a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 23(1):i–??, March 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2017.23.issue-1/mcma-2017-frontmatter1/
mcma-2017-frontmatter1.xml.

Anonymous:2017:Fb

[Ano17b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 23(2):i–??, June 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2017.23.issue-2/mcma-2017-frontmatter2/
mcma-2017-frontmatter2.xml.

Anonymous:2017:Fc

[Ano17c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 23(3):i–??, September 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2017.23.issue-3/mcma-2017-frontmatter3/
mcma-2017-frontmatter3.xml.

Anonymous:2018:Fa

[Ano18a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 24(1):i–??, March 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2018.24.issue-1/mcma-2018-frontmatter1/
mcma-2018-frontmatter1.xml.

Anonymous:2018:Fb

[Ano18b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 24(2):i–??, June 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2018.24.issue-2/mcma-2018-frontmatter2/
mcma-2018-frontmatter2.xml.

Anonymous:2018:Fc

[Ano18c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 24(3):i–??, September 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2018.24.issue-3/mcma-2018-frontmatter3/
mcma-2018-frontmatter3.xml.

Anonymous:2018:Fd

[Ano18d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 24(4):i–??, December 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2018.24.issue-4/mcma-2018-frontmatter4/
mcma-2018-frontmatter4.xml.

Anonymous:2019:Fa

[Ano19a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 25(1):i–??, March 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-1/mcma-2019-frontmatter1/
mcma-2019-frontmatter1.xml.

Anonymous:2019:Fb

[Ano19b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 25(2):i–??, June 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-2/mcma-2019-frontmatter2/
mcma-2019-frontmatter2.xml.

Anonymous:2019:Fc

[Ano19c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 25(3):i–??, September 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-3/mcma-2019-frontmatter3/
mcma-2019-frontmatter3.xml.

Anonymous:2019:F

[Ano19d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 25(4):i–??, December 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-4/mcma-2019-frontmatter4/
mcma-2019-frontmatter4.xml.

Anonymous:2020:Fa

[Ano20a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 26(1):i–??, March 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2020.26.issue-1/mcma-2020-frontmatter1/
mcma-2020-frontmatter1.xml.

Anonymous:2020:Fb

[Ano20b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 26(2):i–iv, June 1, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-frontmatter2/html.

Anonymous:2020:Fc

[Ano20c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 26(3):i–iv, September 1, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-frontmatter3/html.

Anonymous:2020:Fd

[Ano20d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 26(4):i–iv, December 1, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-frontmatter4/html.

Anonymous:2021:Fa

[Ano21a] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 27(1):i–iv, March 1, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-frontmatter1/html.

Anonymous:2021:Fb

[Ano21b] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 27(2):i–iv, June 1, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-frontmatter2/html.

Anonymous:2021:Fc

[Ano21c] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 27(3):i–iv, September 1, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-frontmatter3/html.

Anonymous:2021:Fd

[Ano21d] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 27(4):i–iv, December 1, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-frontmatter4/html.

Anonymous:2022:Fa

[Ano22] Anonymous. Frontmatter. Monte Carlo Methods and Applications, 28(1):i–iv, March 1, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2022-frontmatter1/html.

Antipov:1995:COP

[Ant95] M. V. Antipov. Congruence operator of the pseudo-random numbers generator and a modification of Euclidean decomposition. Monte Carlo Methods and Applications, 1(3):203–219, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1995.1.
issue-3/mcma.1995.1.3.203/mcma.1995.1.3.203.xml.

Antipov:1996:SNM

[Ant96] M. V. Antipov. Sequences of numbers for Monte Carlo methods. Monte Carlo Methods and Applications, 2(3):219–235, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1996.2.
issue-3/mcma.1996.2.3.219/mcma.1996.2.3.219.xml.

Antyufeev:2011:NNR

[Ant11] Victor S. Antyufeev. Non-negative regularization for systems of linear algebraic equations. Monte Carlo Methods and Applications, 17 (4):399–410, December 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2011.17.issue-4/mcma.2011.016/mcma.2011.016.
xml.

Antyufeev:2015:MVM

[Ant15] Victor S. Antyufeev. Mathematical verification of the Monte Carlo maximum cross-section technique. Monte Carlo Methods and Applications, 21(4):275–??, December 2015. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2015.21.issue-4/
mcma-2015-0106/mcma-2015-0106.xml.

Akian:2004:SRE

[AP04] Jean-Luc Akian and Bénédicte Puig. Some results of error evaluation for a non-Gaussian simulation method. Monte Carlo Methods and Applications, 10(1):51–68, ???? 2004. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-1/156939604323091207/156939604323091207.xml.

Aristoff:2015:PRM

[Ari15] David Aristoff. The parallel replica method for computing equilibrium averages of Markov chains. Monte Carlo Methods and Applications, 21(4):255–??, December 2015. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2015.21.issue-4/
mcma-2015-0110/mcma-2015-0110.xml.

Arouna:2004:AMC

[Aro04] Bouhari Arouna. Adaptative Monte Carlo method, a variance reduction technique. Monte Carlo Methods and Applications, 10(1): 1–24, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2004.10.issue-1/156939604323091180/156939604323091180.
xml.

Arsham:1998:TMC

[Ars98] H. Arsham. Techniques for Monte Carlo optimizing. Monte Carlo Methods and Applications, 4(3):181–229, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
3/mcma.1998.4.3.181/mcma.1998.4.3.181.xml.

Arsham:2007:MCT

[Ars07] Hossein Arsham. Monte Carlo techniques for parametric finite multidimensional integral equations. Monte Carlo Methods and Applications, 13(3):173–195, August 2007. CODEN MCMAC6.

ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2007.13.issue-3/mcma.2007.009/
mcma.2007.009.xml.

Amano:1995:SNS

[AS95] Kazuo Amano and Tomoaki Saito. Stochastic numerical solution of biharmonic Dirichlet problem. Monte Carlo Methods and Applications, 1(1):71–82, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1995.1.issue-1/mcma.1995.1.1.
71/mcma.1995.1.1.71.xml.

Achiha:2019:GKW

[ASTY19] Taku Achiha, Hiroshi Sugita, Kenta Tonohiro, and Yuto Yamamoto. Generation of k-wise independent random variables with small randomness. Monte Carlo Methods and Applications, 25(3):259–??, September 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-3/
mcma-2019-2046/mcma-2019-2046.xml.

Abbas-Turki:2014:TCM

[ATBM14] Lokman A. Abbas-Turki, Aych I. Bouselmi, and Mohammed A. Mikou. Toward a coherent Monte Carlo simulation of CVA. Monte Carlo Methods and Applications, 20(3):195–??, September 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-3/mcma-2013-0026/mcma-2013-0026.xml.

Averina:2004:ASS

[Ave04] T. A. Averina. Algorithm of statistical simulation of dynamic systems with distributed change of structure. Monte Carlo Methods and Applications, 10(3–4):221–226, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.221/mcma.2004.10.3-4.221.xml.

Agapie:2010:RPH

[AW10] Stefan C. Agapie and Paula A. Whitlock. Random packing of hyperspheres and Marsaglia’s parking lot test. Monte Carlo Methods and Applications, 16(3–4):197–209, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.019/mcma.2010.019.xml.

Azevedo:2012:QMC

[Aze12] Juarez S. Azevedo. Quasi Monte Carlo methods applied to equations in transient regime on the theis equation. Monte Carlo Methods and Applications, 18(3):201–??, September 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-3/mcma-2012-0007/mcma-2012-0007.xml.

Babovsky:1999:MCS

[Bab99] Hans Babovsky. On a Monte Carlo scheme for Smoluchowski’s coagulation equation. Monte Carlo Methods and Applications, 5(1):1–18, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1999.5.issue-1/mcma.1999.5.1.1/mcma.1999.5.
1.1.xml.

Baldeaux:2008:QMC

[Bal08] Jan Baldeaux. Quasi-Monte Carlo methods for the Kou model. Monte Carlo Methods and Applications, 14(4):281–302, November 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-4/mcma.2008.012/mcma.2008.012.xml.

Branlard:2004:FAS

[BAO^{+}04] J. Branlard, S. Aboud, P. Osuch, S. Goodnick, and M. Saraniti. Frequency analysis of semiconductor devices using full-band cellular Monte Carlo simulations. Monte Carlo Methods and Applications, 10(3–4):227–233, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.227/mcma.2004.10.3-4.227.xml.

Braham:2019:ANM

[BBBR19] Hayette Braham, Louiza Berdjoudj, Mohamed Boualem, and Nadji Rahmania. Analysis of a non-Markovian queueing model: Bayesian statistics and MCMC methods. Monte Carlo Methods and Applications, 25(2):147–??, June 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-2/
mcma-2019-2035/mcma-2019-2035.xml.

Begin:2015:SHM

[BBG15] Jean-François Bégin, Mylène Bédard, and Patrice Gaillardetz. Simulating from the Heston model: A gamma approximation scheme. Monte Carlo Methods and Applications, 21(3):205–??, September 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-3/mcma-2015-0105/mcma-2015-0105.xml.

Brumback:2011:HMH

[BC11] Terry E. Brumback, Jr. and Chien-Pin Chen. Hybrid modeling of homogeneous gas-phase combustion reactions. Monte Carlo Methods and Applications, 17(2):133–154, June 2011. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-2/mcma.2011.006/mcma.2011.006.xml.

Belaribi:2011:PAA

[BCR11] Nadia Belaribi, François Cuvelier, and Francesco Russo. A probabilistic algorithm approximating solutions of a singular PDE of porous media type. Monte Carlo Methods and Applications, 17 (4):317–369, December 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2011.17.issue-4/mcma.2011.014/mcma.2011.014.
xml.

Bally:2005:PHA

[BCZ05] Vlad Bally, Lucia Caramellino, and Antonino Zanette. Pricing and hedging American options by Monte Carlo methods using a Malliavin calculus approach. Monte Carlo Methods and Applications, 11(2):97–133, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2005.11.issue-2/156939605777585944/156939605777585944.
xml.

Bourgey:2020:MMC

[BDGZ20] Florian Bourgey, Stefano De Marco, Emmanuel Gobet, and Alexandre Zhou. Multilevel Monte Carlo methods and lower– upper bounds in initial margin computations. Monte Carlo Methods and Applications, 26(2):131–161, April 15, 2020. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2020-2062/html.

Beachkofski:2009:CDS

[Bea09] Brian Beachkofski. Comparison of descriptive statistics for multidimensional point sets. Monte Carlo Methods and Applications, 15 (3):211–228, November 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2009.15.issue-3/mcma.2009.012/mcma.2009.012.
xml.

Beentjes:2021:OPB

[Bee21] Casper H. L. Beentjes. Optimising Poisson bridge constructions for variance reduction methods. Monte Carlo Methods and Applications, 27(3):249–275, June 1, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2021-2090/html.

Benabdallah:2016:AEM

[BEH16] Mohsine Benabdallah, Youssfi Elkettani, and Kamal Hiderah. Approximation of Euler–Maruyama for one-dimensional stochastic differential equations involving the local times of the unknown process. Monte Carlo Methods and Applications, 22(4):307–??, December 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/view/j/mcma.
2016.22.issue-4/mcma-2016-0115/mcma-2016-0115.xml.

Benov:2016:MPF

[Ben16] Dobriyan M. Benov. The Manhattan Project, the first electronic computer and the Monte Carlo method. Monte Carlo Methods and Applications, 22(1):73–??, March 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-1/
mcma-2016-0102/mcma-2016-0102.xml.

Bernard:2001:CNS

[BF01] Pierre Bernard and Gérard Fleury. Convergence of numerical schemes for stochastic differential equations. Monte Carlo Methods and Applications, 7(1–2):35–44, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.35/mcma.2001.7.1-2.35.xml. Monte Carlo and probabilistic

methods for partial differential equations (Monte Carlo, 2000).

Bossy:1997:CSP

[BFP97] Mireille Bossy, Loula Fezoui, and Serge Piperno. Comparison of a stochastic particle method and a finite volume deterministic method applied to Burgers’ equation. Monte Carlo Methods and Applications, 3(2):113–140, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
2/mcma.1997.3.2.113/mcma.1997.3.2.113.xml.

Bardou:2009:CVC

[BFP09] O. Bardou, N. Frikha, and G. Pagès. Computing VaR and CVaR using stochastic approximation and adaptive unconstrained importance sampling. Monte Carlo Methods and Applications, 15(3):173– 210, November 2009. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2009.15.issue-3/mcma.2009.011/mcma.2009.011.xml.

Braverman:2001:MCA

[BG01] Mark Braverman and Shay Gueron. A Monte Carlo algorithm for a lottery problem. Monte Carlo Methods and Applications, 7(1–2): 73–79, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-1-2/mcma.2001.7.1-2.73/mcma.2001.7.

1-2.73.xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

BenZineb:2013:PCV

[BG13] Tarik Ben Zineb and Emmanuel Gobet. Preliminary control variates to improve empirical regression methods. Monte Carlo Methods and Applications, 19(4):331–??, December 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-4/mcma-2013-0015/mcma-2013-0015.xml.

Barty:2008:AKB

[BGSR08] Kengy Barty, Pierre Girardeau, Cyrille Strugarek, and Jean-Sébastien Roy. Application of kernel-based stochastic gradient algorithms to option pricing. Monte Carlo Methods and Applications, 14(2):99–127, July 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2008.14.issue-2/mcma.2008.006/mcma.2008.006.
xml.

Benabdallah:2018:SRC

[BH18] Mohsine Benabdallah and Kamal Hiderah. Strong rate of convergence for the Euler–Maruyama approximation of one-dimensional stochastic differential equations involving the local time at point zero. Monte Carlo Methods and Applications, 24(4):249–262, December 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2018.24.issue-4/mcma-2018-2021/mcma-2018-2021.xml.

Baliti:2018:MCS

[BHA18] Jamal Baliti, Mohamed Hssikou, and Mohammed Alaoui. Monte Carlo simulation of nonlinear gravity driven Poiseuille–Couette flow in a dilute gas. Monte Carlo Methods and Applications, 24(3):153– 163, September 2018. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL https://www.degruyter.com/view/
j/mcma.2018.24.issue-3/mcma-2018-0014/mcma-2018-0014.xml.

Bishwal:2009:BEI

[Bis09] Jaya P. N. Bishwal. Berry–Esseen inequalities for discretely observed diffusions. Monte Carlo Methods and Applications, 15(3): 229–239, November 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2009.15.issue-3/mcma.2009.013/mcma.2009.013.
xml.

Bossy:2001:SPM

[BJ01] Mireille Bossy and Benjamin Jourdain. A stochastic particle method for the solution of a 1D viscous scalar conservation law in a bounded interval. Monte Carlo Methods and Applications, 7(1–2): 45–53, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-1-2/mcma.2001.7.1-2.45/mcma.2001.7.

1-2.45.xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Ballesio:2022:UEG

[BJ22] Marco Ballesio and Ajay Jasra. Unbiased estimation of the gradient of the log-likelihood for a class of continuous-time state-space models. Monte Carlo Methods and Applications, 28(1):61–83, February 26, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2022-2105/html.

Buglanova:1995:CRS

[BK95] N. A. Buglanova and O. A. Kurbanmuradov. Convergence of the randomized spectral models of homogeneous Gaussian random fields. Monte Carlo Methods and Applications, 1(3):173–201, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1995.1.
issue-3/mcma.1995.1.3.173/mcma.1995.1.3.173.xml.

BenAlaya:2014:MMC

[BK14] Mohamed Ben Alaya and Ahmed Kebaier. Multilevel Monte Carlo for Asian options and limit theorems. Monte Carlo Methods and Applications, 20(3):181–??, September 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-3/mcma-2013-0025/mcma-2013-0025.xml.

Bender:2006:PIA

[BKS06] Christian Bender, Anastasia Kolodko, and John Schoenmakers. Policy iteration for American options: overview. Monte Carlo Methods and Applications, 12(5–6):347–362, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-5/156939606779329053/156939606779329053.xml.

Burch:2015:CET

[BL15] Nathanial Burch and R. B. Lehoucq. Computing the exittime for a finite-range symmetric jump process. Monte Carlo Methods and Applications, 21(2):139–??, June 2015. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-2/mcma-2014-0015/mcma-2014-0015.xml.

Bruti-Liberati:2006:FOS

[BLNSP06] Nicola Bruti-Liberati, Christina Nikitopoulos-Sklibosios, and Eckhard Platen. First order strong approximations of jump diffusions. Monte Carlo Methods and Applications, 12(3–4):191–209, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-3/156939606778705191/156939606778705191.xml.

Bayousef:2019:CIO

[BM19] Manal Bayousef and Michael Mascagni. A computational investigation of the optimal Halton sequence in QMC applications.

Monte Carlo Methods and Applications, 25(3):187–??, September 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-3/mcma-2019-2041/mcma-2019-2041.xml.

Bahlali:2001:SGP

[BMO01] Khaled Bahlali, Brahim Mezerdi, and Youssef Ouknine. Some generic properties in backward stochastic differential equations with continuous coefficient. Monte Carlo Methods and Applications, 7(1–2):15–19, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.15/mcma.2001.7.1-2.15.xml. Monte Carlo and probabilistic

methods for partial differential equations (Monte Carlo, 2000).

Belomestny:2009:MSV

[BMS09] Denis Belomestny, Stanley Mathew, and John Schoenmakers. Multiple stochastic volatility extension of the Libor market model and its implementation. Monte Carlo Methods and Applications, 15(4):285–310, December 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-4/mcma.2009.016/mcma.2009.016.xml.

Bounnite:2015:WCT

[BN15] Mohamed Yasser Bounnite and Abdelaziz Nasroallah. Widening and clustering techniques allowing the use of monotone CFTP algorithm. Monte Carlo Methods and Applications, 21(4):301–??, December 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2015.21.issue-4/mcma-2015-0111/mcma-2015-0111.xml.

Boubalou:2019:CMG

[BOTAZ19] Meriem Boubalou, Megdouda Ourbih-Tari, Abdelouhab Aloui, and Arezki Zioui. Comparing M/G/1 queue estimators in Monte Carlo simulation through the tested generator “getRDS” and the proposed “getLHS” using variance reduction. Monte Carlo Methods and Applications, 25(2):177–??, June 2019. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-2/mcma-2019-2033/mcma-2019-2033.xml.

Bouleau:1995:SPR

[Bou95] Nicolas Bouleau. The shift: properties and recommendations for practical use. Monte Carlo Methods and Applications, 1 (2):137–145, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1995.1.issue-2/mcma.1995.1.2.137/mcma.1995.
1.2.137.xml.

Bouleau:2005:DFS

[Bou05] Nicolas Bouleau. Dirichlet forms in simulation. Monte Carlo Methods and Applications, 11(4):385–395, ???? 2005. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-4/156939605777438541/156939605777438541.xml.

Borisov:1997:AIM

[BP97] N. M. Borisov and M. P. Panin. Adjoint importance Monte Carlo simulation for gamma ray deep penetration problem. Monte Carlo Methods and Applications, 3(3):241–250, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
3/mcma.1997.3.3.241/mcma.1997.3.3.241.xml.

Borisov:1998:AMC

[BP98a] N. M. Borisov and M. P. Panin. Adjoint Monte Carlo calculations of pulse-height-spectrum. Monte Carlo Methods and Applications, 4(3):273–284, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1998.4.issue-3/mcma.1998.4.3.273/mcma.1998.
4.3.273.xml.

Borisov:1998:GPC

[BP98b] N. M. Borisov and M. P. Panin. Generalized particle concept for adjoint Monte Carlo calculations of coupled gamma ray–electron transport. Monte Carlo Methods and Applications, 4(4):341–357, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
1998.4.issue-4/mcma.1998.4.4.341/mcma.1998.4.4.341.xml.

Bonvin:2002:MMV

[BP02] John Bonvin and Marco Picasso. Mesoscopic models for viscoelastic flows: Coupling finite element and Monte Carlo methods. Monte Carlo Methods and Applications, 8(1):73–81, ????

2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2002.8.
issue-1/mcma.2002.8.1.73/mcma.2002.8.1.73.xml.

Bally:2001:SQM

[BPP01] Vlad Bally, Gilles Pagès, and Jacques Printems. A stochastic quantization method for nonlinear problems. Monte Carlo Methods and Applications, 7(1–2):21–33, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.21/mcma.2001.7.1-2.21.xml. Monte Carlo and probabilistic

methods for partial differential equations (Monte Carlo, 2000).

Barghouthi:2003:MCS

[BQA03] Imad A. Barghouthi, Naji A. Qatanani, and Fathi M. Allan. Monte Carlo simulation of Boltzmann equation in space plasma at high latitudes. Monte Carlo Methods and Applications, 9(3):201–216, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2003.9.issue-3/156939603322728978/156939603322728978.xml.

Bouazza:2016:SSS

[BS16] Abdelkader Bouazza and Abderrahmane Settaouti. Study and simulation of the sputtering process of material layers in plasma. Monte Carlo Methods and Applications, 22(2):149–??, June 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-2/mcma-2016-0106/mcma-2016-0106.xml.

Bouazza:2018:UCE

[BS18] Abdelkader Bouazza and Abderrahmane Settaouti. Understanding the contribution of energy and angular distribution in the morphology of thin films using Monte Carlo simulation. Monte Carlo Methods and Applications, 24(3):215–224, September 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-3/mcma-2018-0019/mcma-2018-0019.xml.

Bayer:2010:AWA

[BST10] Christian Bayer, Anders Szepessy, and Raúl Tempone. Adaptive weak approximation of reflected and stopped diffusions. Monte Carlo Methods and Applications, 16(1):1–67, April 2010.

CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-1/mcma.2010.001/mcma.2010.001.xml.

Bally:1996:LES

[BT96] Vlad Bally and Denis Talay. The law of the Euler scheme for stochastic differential equations: II. Convergence rate of the density. Monte Carlo Methods and Applications, 2(2):93–128, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1996.2.
issue-2/mcma.1996.2.2.93/mcma.1996.2.2.93.xml.

Buckwar:2004:MSS

[Buc04] Evelyn Buckwar. The Θ-Maruyama scheme for stochastic functional differential equations with distributed memory term. Monte Carlo Methods and Applications, 10(3–4):235–244, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.235/mcma.2004.10.3-4.235.
xml.

Button:2003:PDP

[But03] Scott D. Button. Project duration prediction using a Monte Carlo simulation of the periodic output of the project resources. Monte Carlo Methods and Applications, 9(3):217–225, ???? 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
3/156939603322728987/156939603322728987.xml.

Bedouhene:2020:BPB

[BZ20] Kahina Bedouhene and Nabil Zougab. A Bayesian procedure for bandwidth selection in circular kernel density estimation. Monte Carlo Methods and Applications, 26(1):69–??, March 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2020.
26.issue-1/mcma-2020-2056/mcma-2020-2056.xml.

Chatterjee:2012:GFM

[CA12] Kausik Chatterjee and Akshay Anantapadmanabhan. A Green’s function Monte Carlo algorithm for the Helmholtz equation subject to Neumann and mixed boundary conditions: Validation with an 1D benchmark problem. Monte Carlo Methods and Applications, 18(3):265–??, September 2012. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-3/mcma-2012-0009/mcma-2012-0009.xml.

Cappe:2001:RCS

[Cap01] Olivier Cappé. Recursive computation of smoothed functionals of hidden Markovian processes using a particle approximation. Monte Carlo Methods and Applications, 7(1–2):81–92, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2001.7.issue-1-2/mcma.2001.7.1-2.81/mcma.2001.7.1-2.81.

xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Cepeda-Cuervo:2015:BBR

[CCG15] Edilberto Cepeda-Cuervo and Liliana Garrido. Bayesian beta regression models with joint mean and dispersion modeling. Monte Carlo Methods and Applications, 21(1):49–??, March 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-1/mcma-2014-0007/mcma-2014-0007.xml.

Cao:2008:ESD

[CCMZ08] Y. Cao, H. Chi, C. Milton, and W. Zhao. Exploitation of sensitivity derivatives via randomized quasi-Monte Carlo methods. Monte Carlo Methods and Applications, 14(3):269–279, September 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-3/mcma.2008.011/mcma.2008.011.xml.

Chraibi:2021:OPF

[CDGG21] Hassane Chraibi, Anne Dutfoy, Thomas Galtier, and Josselin Garnier. Optimal potential functions for the interacting particle system method. Monte Carlo Methods and Applications, 27(2):137–152, April 30, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/mcma-2021-2086/html.

Chi:2013:GPM

[Chi13] Hongmei Chi. Generation of parallel modified Kronecker sequences. Monte Carlo Methods and Applications, 19(4):261–??, December 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-4/mcma-2013-0008/mcma-2013-0008.xml.

Chorin:2001:NMO

[CHK01] Alexandre J. Chorin, Ole H. Hald, and Raz Kupferman. Non-Markovian optimal prediction. Monte Carlo Methods and Applications, 7(1–2):99–109, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.99/mcma.2001.7.1-2.99.xml. Monte Carlo and probabilistic

methods for partial differential equations (Monte Carlo, 2000).

Corcoran:2016:PPS

[CJV16] Jem N. Corcoran, Dale Jennings, and Paul VaughanMiller. Perfect and ϵ-perfect simulation methods for the one-dimensional Kac equation. Monte Carlo Methods and Applications, 22(4):291–??, December 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2016.22.issue-4/mcma-2016-0114/mcma-2016-0114.xml.

Castronovo:2004:SSL

[CK04] Emilio Castronovo and Peter R. Kramer. Subdiffusion and superdiffusion in Lagrangian stochastic models of oceanic transport. Monte Carlo Methods and Applications, 10(3–4):245– 256, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.245/
mcma.2004.10.3-4.245.xml.

Coskun:2018:PBO

[CK18] Sema Coskun and Ralf Korn. Pricing barrier options in the Heston model using the Heath–Platen estimator. Monte Carlo Methods and Applications, 24(1):29–41, March 2018. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-1/mcma-2018-0004/mcma-2018-0004.xml.

Campillo:2001:MCM

[CL01a] Fabien Campillo and Antoine Lejay. A Monte Carlo method to compute the exchange coefficient in the double porosity model. Monte Carlo Methods and Applications, 7(1–2):65–72, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2001.7.issue-1-2/mcma.2001.7.1-2.65/mcma.2001.7.1-2.65.

xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Cepa:2001:IBP

[CL01b] Emmanuel Cépa and Dominique Lépingle. Interacting Brownian particles with strong repulsion. Monte Carlo Methods and Applications, 7(1–2):93–98, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.93/mcma.2001.7.1-2.93.xml. Monte Carlo and probabilistic

methods for partial differential equations (Monte Carlo, 2000).

Campillo:2002:MCM

[CL02a] Fabien Campillo and Antoine Lejay. A Monte Carlo method without grid for a fractured porous domain model. Monte Carlo Methods and Applications, 8(2):129–147, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
2/mcma.2002.8.2.129/mcma.2002.8.2.129.xml.

Crisan:2002:MEA

[CL02b] Dan Crisan and Terry Lyons. Minimal entropy approximations and optimal algorithms. Monte Carlo Methods and Applications, 8(4):343–355, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-4/mcma.2002.8.4.343/mcma.2002.
8.4.343.xml.

Creasey:2018:FGI

[CL18] Peter E. Creasey and Annika Lang. Fast generation of isotropic Gaussian random fields on the sphere. Monte Carlo Methods and Applications, 24(1):1–11, March 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-1/
mcma-2018-0001/mcma-2018-0001.xml.

Cattiaux:2017:IDE

[CLP17] Patrick Cattiaux, José R. León, and Clémentine Prieur. Invariant density estimation for a reflected diffusion using an Euler scheme. Monte Carlo Methods and Applications, 23(2):71–??, June 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-2/mcma-2017-0104/mcma-2017-0104.xml.

Comman:2001:VNL

[Com01] Henri Comman. A view on noncommutative large deviations from a theory of noncommutative capacities. Monte Carlo Methods and Applications, 7(1–2):125–129, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.125/mcma.2001.7.1-2.125.xml. Monte

Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Costantini:2001:SVR

[Cos01] C. Costantini. A SIMPLE VARIANCE REDUCTION METHOD WITH APPLICATIONS TO FINANCE AND QUEUEING THE-ORY. Monte Carlo Methods and Applications, 7(1–2):131–139, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2001.7.issue-1-2/mcma.2001.7.1-2.131/mcma.2001.7.

1-2.131.xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Christensen:2001:MCI

[CP01] Bent Jesper Christensen and Rolf Poulsen. Monte Carlo improvement of estimates of the mean-reverting constant elasticity of variance interest rate diffusion. Monte Carlo Methods and Applications, 7(1–2):111–123, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.111/mcma.2001.7.1-2.111.xml. Monte

Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Caramellino:2002:SEH

[CP02] Lucia Caramellino and Barbara Pacchiarotti. Sharp estimates for the hitting probability on time-dependent barriers for a Brownian Motion. Weak approximation of a Brownian motion killed on time-dependent barriers. Monte Carlo Methods and Applications, 8(3):221–236, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
3/mcma.2002.8.3.221/mcma.2002.8.3.221.xml.

Corlay:2015:FQB

[CP15] Sylvain Corlay and Gilles Pagès. Functional quantization-based stratified sampling methods. Monte Carlo Methods and Applications, 21(1):1–??, March 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2015.21.issue-1/mcma-2014-0010/mcma-2014-0010.
xml.

Castro:2007:FGB

[CPSH07] Francesc Castro, Gustavo Patow, Mateu Sbert, and John H. Halton. Fast GPU-based reuse of paths in radiosity. Monte Carlo Methods and Applications, 13(4):253–273, November 20, 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-4/mcma.2007.014/mcma.2007.014.xml.

Chan:2018:SD

[CRGF18] Debora Chan, Andrea Rey, Juliana Gambini, and Alejandro C. Frery. Sampling from the G_{I}^{0} distribution. Monte Carlo Methods and Applications, 24(4):271–287, December 2018. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-4/mcma-2018-2023/mcma-2018-2023.xml.

Chatterjee:2014:NGF

[CRS14] Kausik Chatterjee, John R. Roadcap, and Surendra Singh. A new Green’s function Monte Carlo algorithm for the solution of the three-dimensional nonlinear Poisson–Boltzmann equation: Application to the modeling of plasma sheath layers. Monte Carlo Methods and Applications, 20(1):53–??, March 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-1/mcma-2013-0016/mcma-2013-0016.xml.

Cancela:2002:MEU

[CRT02] Héctor Cancela, Gerardo Rubino, and Bruno Tuffin. MTTF estimation using importance sampling on Markov models. Monte Carlo Methods and Applications, 8(4):321–341, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
4/mcma.2002.8.4.321/mcma.2002.8.4.321.xml.

Chauhan:1996:PAQ

[CS96] M. S. Chauhan and G. C. Sharma. Profit analysis of M/M/r queueing model with balking and reneging. Monte Carlo Methods and Applications, 2(2):139–144, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
2/mcma.1996.2.2.139/mcma.1996.2.2.139.xml.

deBouard:2001:TNA

[dBDD01] A. de Bouard, A. Debussche, and L. Di Menza. Theoretical and numerical aspects of stochastic nonlinear Schrödinger equations. Monte Carlo Methods and Applications, 7(1–2):55–63, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2001.7.issue-1-2/mcma.2001.7.1-2.55/mcma.2001.7.1-2.55.

xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Doerr:2008:CCC

[DGKP08] Benjamin Doerr, Michael Gnewuch, Peter Kritzer, and Friedrich Pillichshammer. Component-by-component construction of lowdiscrepancy point sets of small size. Monte Carlo Methods and Applications, 14(2):129–149, July 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2008.14.issue-2/mcma.2008.007/
mcma.2008.007.xml.

Dick:2006:TSM

[Dic06] Josef Dick. A Taylor space for multivariate integration. Monte Carlo Methods and Applications, 12(2):99–112, ???? 2006. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-2/156939606777488860/156939606777488860.xml.

Dimov:1998:PCE

[DK98a] Ivan Dimov and Aneta Karaivanova. Parallel computations of eigenvalues based on a Monte Carlo approach. Monte Carlo Methods and Applications, 4(1):33–52, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
1/mcma.1998.4.1.33/mcma.1998.4.1.33.xml.

Dreyer:1998:RES

[DK98b] Wolfgang Dreyer and Matthias Kunik. Reflections of Eulerian shock waves at moving adiabatic boundaries. Monte Carlo Methods and Applications, 4(3):231–252, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
3/mcma.1998.4.3.231/mcma.1998.4.3.231.xml.

Dick:2006:BPU

[DK06] Josef Dick and Peter Kritzer. A best possible upper bound on the star discrepancy of (t, m, 2)-nets. Monte Carlo Methods and Applications, 12(1):1–17, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2006.12.issue-1/156939606776886643/
156939606776886643.xml.

Dreyer:1998:IPM

[DKS^{+}98] W. Dreyer, M. Kunik, K. Sabelfeld, N. Simonov, and K. Wilmanski. Iterative procedure for multidimensional Euler equations. Monte Carlo Methods and Applications, 4(3):253–271, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
3/mcma.1998.4.3.253/mcma.1998.4.3.253.xml.

DelChicca:2014:HMC

[DL14] Lucia Del Chicca and Gerhard Larcher. Hybrid Monte Carlo methods in credit risk management. Monte Carlo Methods and Applications, 20(4):245–??, December 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-4/mcma-2014-0004/mcma-2014-0004.xml.

DeLuigi:2010:AIA

[DM10] Christophe De Luigi and Sylvain Maire. Adaptive integration and approximation over hyper-rectangular regions with applications to basket option pricing. Monte Carlo Methods and Applications, 16(3–4):265–282, December 2010. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.011/mcma.2010.011.xml.

Dereudre:2016:ERT

[DMR16] David Dereudre, Sara Mazzonetto, and Sylvie Roelly. An explicit representation of the transition densities of the skew Brownian motion with drift and two semipermeable barriers. Monte Carlo Methods and Applications, 22(1):1–??, March 2016. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-1/mcma-2016-0100/mcma-2016-0100.xml.

Ding:2003:QMC

[DMZ03] Jiu Ding, Dong Mao, and Aihui Zhou. A quasi Monte Carlo approach to piecewise linear Markov approximations of Markov operators. Monte Carlo Methods and Applications, 9(4):295–306, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2003.9.
issue-4/156939603322601932/156939603322601932.xml.

Dubus:2010:ESP

[DS10] Alain Dubus and Karl Sabelfeld. Editorial [selected papers from the Seventh IMACS Seminar on Monte Carlo methods]. Monte Carlo Methods and Applications, 16(3–4):195, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.021/mcma.2010.021.xml. Held at the

Université Libre de Bruxelles, Brussels, September 6–11, 2009.

Descombes:2001:RAO

[DSGZ01] Xavier Descombes, Radu Stoica, Laurent Garcin, and Josiane Zerubia. A RJMCMC algorithm for object processes in image processing. Monte Carlo Methods and Applications, 7(1–2):149– 156, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-1-2/mcma.2001.7.1-2.149/mcma.2001.

7.1-2.149.xml. Monte Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Deaconu:2001:GCB

[DT01] Madalina Deaconu and Etienne Tanré. A generalization of the connection between the additive and multiplicative solutions for the Smoluchowski’s coagulation equation. Monte Carlo Methods and Applications, 7(1–2):141–147, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.141/mcma.2001.7.1-2.141.xml. Monte

Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Dimov:2022:SHE

[DTS22] Ivan Dimov, Venelin Todorov, and Karl Sabelfeld. A study of highly efficient stochastic sequences for multidimensional sensitivity analysis. Monte Carlo Methods and Applications, 28(1):1–12, February 15, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2022-2101/html.

Evarest:2018:WDP

[EBSY18] Emmanuel Evarest, Fredrik Berntsson, Martin Singull, and Xiangfeng Yang. Weather derivatives pricing using regime switching model. Monte Carlo Methods and Applications, 24(1):13–27, March 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-1/mcma-2018-0002/mcma-2018-0002.xml.

Ech-Chafiq:2021:ACV

[ECLR21] Zineb El Filali Ech-Chafiq, Jérôme Lelong, and Adil Reghai. Automatic control variates for option pricing using neural networks. Monte Carlo Methods and Applications, 27(2):91–104, January 13, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2081/html.

Egecioglu:2009:UGA

[Eğe09] Ömer Eğecioğlu. Uniform generation of anonymous and neutral preference profiles for social choice rules. Monte Carlo Methods and Applications, 15(3):241–255, November 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-3/mcma.2009.014/mcma.2009.014.xml.

Egorov:1997:IGM

[Ego97] A. D. Egorov. On L^{2}-isomorphic Gaussian models for non-Gaussian distributions. Monte Carlo Methods and Applications, 3(2):141–154, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1997.3.issue-2/mcma.1997.3.2.141/mcma.1997.
3.2.141.xml.

Egorov:2007:AEF

[Ego07] A. D. Egorov. Approximations for expectations of functionals of solutions to stochastic differential equations. Monte Carlo Methods and Applications, 13(4):275–285, November 20, 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-4/mcma.2007.015/mcma.2007.015.xml.

Egorov:2020:AFC

[Ego20] Alexander Egorov. An approximate formula for calculating the expectations of functionals from random processes based on using the Wiener chaos expansion. Monte Carlo Methods and Applications, 26(4):285–292, October 7, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2020-2074/html.

Elkhechafi:2018:NHC

[EHE18] Mariam Elkhechafi, Hanaa Hachimi, and Youssfi Elkettani. A new hybrid cuckoo search and firefly optimization. Monte Carlo Methods and Applications, 24(1):71–77, March 2018. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-1/mcma-2018-0003/mcma-2018-0003.xml.

Ermakov:2018:RRQ

[EL18] Sergej M. Ermakov and Svetlana N. Leora. Remarks on randomization of quasi-random numbers. Monte Carlo Methods and Applications, 24(2):139–145, June 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-2/
mcma-2018-0012/mcma-2018-0012.xml.

Entacher:2004:NST

[ELRU04] Karl Entacher, Gerold Laimer, Harald Röck, and Andreas Uhl. Normalization of the spectral test in high dimensions. Monte Carlo Methods and Applications, 10(3–4):265–274, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.265/mcma.2004.10.3-4.265.
xml.

ElHaddad:2010:DNM

[ELV10] Rami El Haddad, Christian Lécot, and Gopalakrishnan Venkiteswaran. Diffusion in a nonhomogeneous medium: quasi-random walk on a lattice. Monte Carlo Methods and Applications, 16(3–4):211–230, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2010.16.issue-3-4/mcma.2010.009/mcma.2010.009.xml.

Eichler:2011:CFM

[ELZ11] Andreas Eichler, Gunther Leobacher, and Heidrun Zellinger. Calibration of financial models using quasi-Monte Carlo. Monte Carlo Methods and Applications, 17(2):99–131, June 2011. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-2/mcma.2011.004/mcma.2011.004.xml.

Egorov:2003:AEC

[EM03] A. D. Egorov and V. B. Malyutin. Approximate evaluation of a class of functional integrals over space of functions taking values ±1. Monte Carlo Methods and Applications, 9(4):307–314, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2003.9.
issue-4/156939603322601941/156939603322601941.xml.

Etore:2013:ESO

[ÉM13] Pierre Étoré and Miguel Martinez. Exact simulation of onedimensional stochastic differential equations involving the local time at zero of the unknown process. Monte Carlo Methods and Applications, 19(1):41–??, March 2013. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-1/mcma-2013-0002/mcma-2013-0002.xml.

Egorov:2017:MCC

[EM17] Alexander Egorov and Victor Malyutin. A method for the calculation of characteristics for the solution to stochastic differential equations. Monte Carlo Methods and Applications, 23(3):149–??, September 2017. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL https://www.degruyter.com/view/
j/mcma.2017.23.issue-3/mcma-2017-0110/mcma-2017-0110.xml.

Elkimakh:2020:HMM

[EN20] Karima Elkimakh and Abdelaziz Nasroallah. Hidden Markov model with Markovian emission. Monte Carlo Methods and Applications, 26(4):303–313, August 11, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2020-2072/html.

Ermakov:2019:SSD

[EP19] Sergej M. Ermakov and Anna A. Pogosian. On solving stochastic differential equations. Monte Carlo Methods and Applications, 25(2):155–??, June 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-2/mcma-2019-2038/mcma-2019-
2038.xml.

Ermakov:2006:QMC

[ER06] S. M. Ermakov and A. Rukavishnikova. Quasi-Monte Carlo algorithms for solving linear algebraic equations. Monte Carlo Methods and Applications, 12(5–6):363–384, ???? 2006. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-5/156939606779329071/156939606779329071.xml.

Ermakov:2011:VSM

[Erm11] Sergej M. Ermakov. Variance of the simplest Monte Carlo estimators in the sign-changing case. Monte Carlo Methods and Applications, 17(4):411–417, December 2011. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-4/mcma.2011.017/mcma.2011.017.xml.

Egorov:2010:AFE

[ES10] A. Egorov and K. Sabelfeld. Approximate formulas for expectations of functionals of solutions to stochastic differential equations. Monte Carlo Methods and Applications, 16(2):95–127, July 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-2/mcma.2010.003/mcma.2010.003.xml.

Englander:2011:CBR

[ES11] János Engländer and Nándor Sieben. Critical branching random walk in an IID environment. Monte Carlo Methods and Applications, 17(2):169–193, June 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2011.17.issue-2/mcma.2011.008/
mcma.2011.008.xml.

ElKhaldi:2017:TMD

[ES17] Khaldoun El Khaldi and Elias G. Saleeby. On the tangent model for the density of lines and a Monte Carlo method for computing hypersurface area. Monte Carlo Methods and Applications, 23(1):13–??, March 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2017.23.issue-1/mcma-2017-0100/mcma-2017-0100.xml.

ElKhaldi:2020:DLS

[ES20] Khaldoun El Khaldi and Elias G. Saleeby. On the density of lines and Santalo’s formula for computing geometric size measures. Monte Carlo Methods and Applications, 26(4):315–323, August 5, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2071/html.

Entacher:1998:LCG

[EUW98] K. Entacher, A. Uhl, and S. Wegenkittl. Linear congruential generators for parallel Monte Carlo: the leap-frog case. Monte Carlo Methods and Applications, 4(1):1–15, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
1/mcma.1998.4.1.1/mcma.1998.4.1.1.xml.

Eibeck:2001:SAS

[EW01] Andreas Eibeck and Wolfgang Wagner. Stochastic algorithms for studying coagulation dynamics and gelation phenomena. Monte Carlo Methods and Applications, 7(1–2):157–165, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.157/mcma.2001.7.1-2.157.xml. Monte

Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Ermakov:2002:MCD

[EW02] Sergej M. Ermakov and Wolfgang Wagner. Monte Carlo difference schemes for the wave equation. Monte Carlo Methods and Applications, 8(1):1–29, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-1/mcma.2002.8.1.1/mcma.2002.8.
1.1.xml.

Egorov:2004:AFI

[EZ04] A. D. Egorov and A. V. Zherelo. Approximations of functional integrals with respect to measures generated by solutions of stochastic differential equations. Monte Carlo Methods and Applications, 10(3–4):257–264, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.257/mcma.2004.10.3-4.257.xml.

Zakrad:2022:ESS

[eZN22] Az eddine Zakrad and Abdelaziz Nasroallah. Estimation of steadystate quantities of an HMM with some rarely generated emissions. Monte Carlo Methods and Applications, 28(1):27–44, February 15, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2022-2103/html.

Fournier:2004:ESN

[FG04] Nicolas Fournier and Jean-Sébastien Giet. Exact simulation of nonlinear coagulation processes. Monte Carlo Methods and Applications, 10(2):95–106, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2004.10.issue-2/156939604777303253/
156939604777303253.xml.

Field:2013:AFG

[FGD13] Richard V. Field, Jr., Mircea Grigoriu, and Clark R. Dohrmann. An algorithm for on-the-fly generation of samples of non-stationary Gaussian processes based on a sampling theorem. Monte Carlo Methods and Applications, 19(2):143–??, July 2013. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-2/mcma-2013-0004/mcma-2013-0004.xml.

Figotin:2001:GQS

[FGM^{+}01] A. Figotin, A. Gordon, S. Molchanov, J. Quinn, and N. Stavrakas. Generalized quantum statistics and testing of randomizers with and without asymptotic assumptions. Monte Carlo Methods and Applications, 7(1–2):167–175, ???? 2001. CODEN MCMAC6.

ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.167/mcma.2001.7.1-2.167.xml.

Fort:2017:MDB

[FGM17] Gersende Fort, Emmanuel Gobet, and Eric Moulines. MCMC design-based non-parametric regression for rare event. Application to nested risk computations. Monte Carlo Methods and Applications, 23(1):21–??, March 2017. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2017.23.issue-1/
mcma-2017-0101/mcma-2017-0101.xml.

Fernandez:2013:DBF

[FHS13] Lexuri Fernández, Peter Hieber, and Matthias Scherer. Doublebarrier first-passage times of jump-diffusion processes. Monte Carlo Methods and Applications, 19(2):107–??, July 2013. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-2/mcma-2013-0005/mcma-2013-0005.xml.

Fujita:2002:GVC

[FIN02] Takahiko Fujita, Shunji Ito, and Syoiti Ninomiya. The generalized van der Corput sequence and its application to numerical integration. Monte Carlo Methods and Applications, 8(2):149–158, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2002.8.
issue-2/mcma.2002.8.2.149/mcma.2002.8.2.149.xml.

Fujita:2008:UDS

[FKM08] Takahiko Fujita, Hiroshi Kaneko, and Shin Matsumoto. A uniformly distributed sequence on the ring of p-adic integers. Monte Carlo Methods and Applications, 14(4):303–310, November 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-4/mcma.2008.013/mcma.2008.013.xml.

Faure:2010:IHS

[FL10] Henri Faure and Christiane Lemieux. Improved Halton sequences and discrepancy bounds. Monte Carlo Methods and Applications, 16(3–4):231–250, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2010.16.issue-3-4/mcma.
2010.008/mcma.2010.008.xml.

Fournier:2001:MCA

[FM01] Nicolas Fournier and Sylvie Méléard. Monte-Carlo approximations for 2D homogeneous Boltzmann equations without cutoff and for non Maxwell molecules. Monte Carlo Methods and Applications, 7(1–2):177–192, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.177/mcma.2001.7.1-2.177.xml. Monte

Carlo and probabilistic methods for partial differential equations (Monte Carlo, 2000).

Fakhouri:2009:SMC

[FN09] H. Fakhouri and A. Nasroallah. On the simulation of Markov chain steady-state distribution using CFTP algorithm. Monte Carlo Methods and Applications, 15(2):91–105, August 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-2/mcma.2009.005/mcma.2009.005.xml.

Fischer:1999:ABM

[FP99] Paul Fischer and Eckhard Platen. Applications of the balanced method to stochastic differential equations in filtering. Monte Carlo Methods and Applications, 5(1):19–38, ???? 1999. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
1/mcma.1999.5.1.19/mcma.1999.5.1.19.xml.

Fort:2002:DSS

[FP02] Jean-Claude Fort and Gilles Pagès. Decreasing step stochastic algorithms: a.s. behaviour of weighted empirical measures. Monte Carlo Methods and Applications, 8(3):237–270, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
3/mcma.2002.8.3.237/mcma.2002.8.3.237.xml.

Frikha:2012:QBR

[FS12] Noufel Frikha and Abass Sagna. Quantization based recursive importance sampling. Monte Carlo Methods and Applications, 18 (4):287–??, December 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2012.18.issue-4/mcma-2012-0011/mcma-2012-0011.
xml.

Fukuyama:2000:PFA

[FT00] Katusi Fukuyama and Tetsuo Tomokuni. On pseudorandom functions and asymptotic distributions. Monte Carlo Methods and Applications, 6(3):167–174, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
3/mcma.2000.6.3.167/mcma.2000.6.3.167.xml.

Fukuyama:1996:RRS

[Fuk96] Katusi Fukuyama. Riesz–Raikov sums and Weyl transform. Monte Carlo Methods and Applications, 2(4):271–293, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
4/mcma.1996.2.4.271/mcma.1996.2.4.271.xml.

Fathi-Vajargah:2016:IMC

[FVK16] Behrouz Fathi-Vajargah and Mohadeseh Kanafchian. Improved Markov chain Monte Carlo method for cryptanalysis substitution–transposition cipher. Monte Carlo Methods and Applications, 22(4):323–??, December 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-4/
mcma-2016-0116/mcma-2016-0116.xml.

Fathi-Vajargah:2017:IMC

[FVK17] Behrouz Fathi-Vajargah and Mohadeseh Kanafchian. Improved Markov chain Monte Carlo method for cryptanalysis substitution-transposition cipher. Monte Carlo Methods and Applications, 23(2):147–??, June 2017. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2017.23.issue-2/
mcma-2017-0108/mcma-2017-0108.xml.

Grecksch:1999:ASH

[GA99] W. Grecksch and V. V. Anh. Approximation of stochastic Hammerstein integral equation with fractional Brownian motion input. Monte Carlo Methods and Applications, 5(4):311–323, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1999.5.
issue-4/mcma.1999.5.4.311/mcma.1999.5.4.311.xml.

Gabih:2005:OPS

[GG05] Abdelali Gabih and Wilfried Grecksch. An ϵ-optimal portfolio with stochastic volatility. Monte Carlo Methods and Applications, 11 (1):1–14, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2005.11.issue-1/1569396054027256/1569396054027256.
xml.

Gil:2006:RTP

[GGP06] Manuel Gil, Gaston H. Gonnet, and Wesley P. Petersen. A repetition test for pseudo-random number generators. Monte Carlo Methods and Applications, 12(5–6):385–393, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-5/156939606779329017/156939606779329017.xml;
http://www.inf.ethz.ch/personal/gonnet/RepetitionTest.
html.

Grecksch:2000:PPA

[GHT00] W. Grecksch, F. Heyde, and Chr. Tammer. Proximal point algorithm for an approximated stochastic optimal control problem. Monte Carlo Methods and Applications, 6(3):175–189, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2000.6.
issue-3/mcma.2000.6.3.175/mcma.2000.6.3.175.xml.

Gormin:2008:WVM

[GK08] A. A. Gormin and Y. N. Kashtanov. The weighted variance minimization for options pricing. Monte Carlo Methods and Applications, 13(5–6):333–351, January 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
13.issue-5-6/mcma.2007.018/mcma.2007.018.xml.

Giorgi:2017:LTW

[GLP17] Daphné Giorgi, Vincent Lemaire, and Gilles Pagès. Limit theorems for weighted and regular multilevel estimators. Monte Carlo Methods and Applications, 23(1):43–??, March 2017. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-1/mcma-2017-0102/mcma-2017-0102.xml.

Gobet:2004:SMC

[GM04] Emmanuel Gobet and Sylvain Maire. A spectral Monte Carlo method for the Poisson equation. Monte Carlo Methods and Applications, 10(3–4):275–285, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.275/mcma.2004.10.3-4.275.xml.

Golyandina:1999:HBE

[GN99] N. Golyandina and V. Nekrutkin. Homogeneous balance equations for measures: Errors of the stochastic solution. Monte Carlo Methods and Applications, 5(3):193–261, ???? 1999. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
3/mcma.1999.5.3.193/mcma.1999.5.3.193.xml.

Gerlovina:2005:ABL

[GN05] V. Gerlovina and V. Nekrutkin. Asymptotical behavior of linear congruential generators. Monte Carlo Methods and Applications, 11(2):135–162, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2005.11.issue-2/156939605777585971/156939605777585971.
xml.

Gobet:2001:ESW

[Gob01] Emmanuel Gobet. Efficient schemes for the weak approximation of reflected diffusions. Monte Carlo Methods and Applications, 7(1–2):193–202, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-1-2/mcma.2001.7.
1-2.193/mcma.2001.7.1-2.193.xml. Monte Carlo and probabilis-

tic methods for partial differential equations (Monte Carlo, 2000).

Golyandina:2003:CLT

[Gol03] N. Golyandina. Central Limit Theorem for (n, k)-particle processes solving balance equations. Monte Carlo Methods and Applications, 9(1):1–11, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
1/156939603322587425/156939603322587425.xml.

Golyandina:2004:CRS

[Gol04] N. Golyandina. Convergence rate for spherical processes with shifted centres. Monte Carlo Methods and Applications, 10(3–4): 287–296, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.287/
mcma.2004.10.3-4.287.xml.

Greslehner:2012:DHR

[GP12] Julia Greslehner and Friedrich Pillichshammer. Discrepancy of higher rank polynomial lattice point sets. Monte Carlo Methods and Applications, 18(1):79–108, March 2012. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-1/mcma-2012-0001/mcma-2012-0001.xml.

Gur:2019:SBC

[GP19] Sercan Gür and Klaus Pötzelberger. Sensitivity of boundary crossing probabilities of the Brownian motion. Monte Carlo Methods and Applications, 25(1):75–83, March 2019. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-1/mcma-2019-2031/mcma-2019-2031.xml.

Grecksch:2008:QSP

[GR08] Wilfried Grecksch and Christian Roth. A quasilinear stochastic partial differential equation driven by fractional white noise. Monte Carlo Methods and Applications, 13(5–6):353–367, January 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
13.issue-5-6/mcma.2007.019/mcma.2007.019.xml.

Grigoriu:2010:SBM

[Gri10] M. Grigoriu. A spectral-based Monte Carlo algorithm for generating samples of nonstationary Gaussian processes. Monte Carlo Methods and Applications, 16(2):143–165, July 2010. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-2/mcma.2010.006/mcma.2010.006.xml.

Grigoriu:2014:EMC

[Gri14] Mircea Grigoriu. An efficient Monte Carlo solution for problems with random matrices. Monte Carlo Methods and Applications, 20(2):121–??, June 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2014.20.issue-2/mcma-2013-0021/mcma-2013-0021.
xml.

Grigoriu:2017:MCA

[Gri17] Mircea Grigoriu. Monte Carlo algorithm for vector-valued Gaussian functions with preset component accuracies. Monte Carlo Methods and Applications, 23(3):165–??, September 2017. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-3/mcma-2017-0112/mcma-2017-0112.xml.

Guias:1997:MCA

[Gui97] Flavius Guiaş. A Monte Carlo approach to the Smoluchowski equations. Monte Carlo Methods and Applications, 3(4):313–326, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1997.3.
issue-4/mcma.1997.3.4.313/mcma.1997.3.4.313.xml.

Guias:1999:DSM

[Gui99] Flavius Guiaş. A direct simulation method for the coagulation– fragmentation. Equations with multiplicative coagulation kernels. Monte Carlo Methods and Applications, 5(4):287–309, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1999.5.
issue-4/mcma.1999.5.4.287/mcma.1999.5.4.287.xml.

Guias:2008:GBD

[Gui08] Flavius Guiaş. Generalized Becker–Döring equations modeling the time evolution of a process of preferential attachment with fitness. Monte Carlo Methods and Applications, 14(2):151–170, July 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-2/mcma.2008.008/mcma.2008.008.xml.

Gueron:2001:SEC

[GZ01] Shay Gueron and Or Zuk. On Smoluchowski equations for coagulation processes with multiple absorbing states. Monte Carlo Methods and Applications, 7(1–2):203–211, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
1-2/mcma.2001.7.1-2.203/mcma.2001.7.1-2.203.xml.

Habibi:2011:NAD

[Hab11] Reza Habibi. A note on approximating distribution functions of cusum and cusumsq tests. Monte Carlo Methods and Applications, 17(1):1–10, March 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2011.17.issue-1/mcma.2011.005/mcma.2011.005.
xml.

Habibi:2012:NNM

[Hab12] Reza Habibi. A note on Newton’s method for system of stochastic differential equations. Monte Carlo Methods and Applications, 18 (4):275–??, December 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2012.18.issue-4/mcma-2012-0010/mcma-2012-0010.
xml.

Halton:2004:OQP

[Hal04] John H. Halton. An outline of quasi-probability: Why quasi-Monte-Carlo methods are statistically valid and how their errors can be estimated statistically. Monte Carlo Methods and Applications, 10(3– 4):183–196, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.183/
mcma.2004.10.3-4.183.xml.

Halton:2005:CQP

[Hal05a] J. H. Halton. Corrigenda: “Quasi-probability. Why quasi-Monte–Carlo methods are statistically valid and how their errors can be estimated statistically” [Monte Carlo Methods Appl. 11 (2005), no. 3, 203–350; Cno. 2159755]. Monte Carlo Methods and Applications, 11(4):463, ???? 2005. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-4/156939605777438550/156939605777438550.xml. See

[Hal05b].

Halton:2005:QPW

[Hal05b] John H. Halton. Quasi-probability. Why quasi-Monte–Carlo methods are statistically valid and how their errors can be estimated statistically. Monte Carlo Methods and Applications, 11(3):203–350, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/

mcma.2005.11.issue-3/1569396054495130/1569396054495130.

xml. See corrigenda [Hal05a].

Halton:2006:SMC

[Hal06] John H. Halton. Sequential Monte Carlo techniques for solving nonlinear systems. Monte Carlo Methods and Applications, 12(2):113– 141, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2006.12.issue-2/156939606777488879/156939606777488879.
xml.

Halton:2008:SMC

[Hal08a] John H. Halton. Sequential Monte Carlo for linear systems — a practical summary. Monte Carlo Methods and Applications, 14(1): 1–27, May 2008. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2008.14.issue-1/mcma.2008.001/mcma.2008.001.xml.

Halton:2008:SAT

[Hal08b] John H. Halton. Sigma-algebra theorems. Monte Carlo Methods and Applications, 14(2):171–189, July 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-2/mcma.2008.009/mcma.2008.009.xml.

Halidias:2015:CPP

[Hal15a] Nikolaos Halidias. Constructing positivity preserving numerical schemes for the two-factor CIR model. Monte Carlo Methods and Applications, 21(4):313–??, December 2015. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2015.
21.issue-4/mcma-2015-0109/mcma-2015-0109.xml.

Halidias:2015:NNS

[Hal15b] Nikolaos Halidias. A new numerical scheme for the CIR process. Monte Carlo Methods and Applications, 21(3):245–??, September 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-3/mcma-2015-0101/mcma-2015-0101.xml.

Halidias:2016:CBP

[Hal16] Nikolaos Halidias. On the construction of boundary preserving numerical schemes. Monte Carlo Methods and Applications, 22 (4):277–??, December 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2016.22.issue-4/mcma-2016-0113/mcma-2016-
0113.xml.

Halidias:2021:APM

[Hal21] Nikolaos Halidias. On the absorption probabilities and mean time for absorption for discrete Markov chains. Monte Carlo Methods and Applications, 27(2):105–115, February 2, 2021. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2084/html.

Harase:2016:SEL

[Har16] Shin Harase. A search for extensible low-WAFOM point sets. Monte Carlo Methods and Applications, 22(4):349–??, December 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-4/mcma-2016-0119/mcma-2016-0119.xml.

Harase:2019:CSS

[Har19] Shin Harase. Comparison of Sobol’ sequences in financial applications. Monte Carlo Methods and Applications, 25(1):61–74, March 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-1/mcma-2019-2029/mcma-2019-2029.xml.

Hausenblas:2000:MCSb

[Hau00a] Erika Hausenblas. Monte Carlo simulation of killed diffusion. Monte Carlo Methods and Applications, 6(4):263–295, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2000.6.
issue-4/mcma.2000.6.4.263/mcma.2000.6.4.263.xml.

Hausenblas:2000:MCSa

[Hau00b] Erika Hausenblas. Monte Carlo simulation of reflected stochastic differential equations driven by Poisson random measures. Monte Carlo Methods and Applications, 6(1):1–14, ???? 2000. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
1/mcma.2000.6.1.1/mcma.2000.6.1.1.xml.

Hausenblas:2000:NSU

[Hau00c] Erika Hausenblas. A numerical scheme using excursion theory for simulating stochastic differential equations with reflection and local time at a boundary. Monte Carlo Methods and Applications, 6(2):81–103, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2000.6.issue-2/mcma.2000.6.2.
81/mcma.2000.6.2.81.xml.

Hssikou:2016:PCF

[HBA16] Mohamed Hssikou, Jamal Baliti, and Mohammed Alaoui. The planar Couette flow with slip and jump boundary conditions in a microchannel. Monte Carlo Methods and Applications, 22(4):337–??, December 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2016.22.issue-4/mcma-2016-0117/mcma-2016-0117.xml.

Hssikou:2015:DMT

[HBBA15] Mohamed Hssikou, Jamal Baliti, Yassir Bouzineb, and Mohammed Alaoui. DSMC method for a two-dimensional flow with a gravity field in a square cavity. Monte Carlo Methods and Applications, 21(1):59–??, March 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2015.21.issue-1/mcma-2014-0009/mcma-2014-0009.
xml.

Heinrich:1995:VRM

[Hei95] Stefan Heinrich. Variance reduction for Monte Carlo methods by means of deterministic numerical computation. Monte Carlo Methods and Applications, 1(4):251–277, ???? 1995. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
4/mcma.1995.1.4.251/mcma.1995.1.4.251.xml.

Heinrich:2004:PQA

[Hei04] Stefan Heinrich. On the power of quantum algorithms for vector valued mean computation. Monte Carlo Methods and Applications, 10(3–4):297–310, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2004.10.issue-3-4/mcma.
2004.10.3-4.297/mcma.2004.10.3-4.297.xml.

Heinz:2008:RDS

[Hei08] Stefan Heinz. Realizability of dynamic subgrid-scale stress models via stochastic analysis. Monte Carlo Methods and Applications, 14 (4):311–329, November 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2008.14.issue-4/mcma.2008.014/mcma.2008.014.
xml.

Heinz:2014:UQW

[Hei14] Stefan Heinz. Uncertainty quantification of world population growth: A self-similar PDF model. Monte Carlo Methods and Applications, 20(4):261–??, December 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-4/mcma-2014-0005/mcma-2014-0005.xml.

Hiderah:2020:AEM

[Hid20] Kamal Hiderah. Approximation of Euler–Maruyama for onedimensional stochastic differential equations involving the maximum process. Monte Carlo Methods and Applications, 26(1):33–??, March 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2020.26.issue-1/mcma-2020-2057/mcma-2020-2057.xml.

Hwang:2014:FIK

[HK14] Chi-Ok Hwang and Seung-Yeon Kim. Field-induced Kosterlitz– Thouless transition in critical triangular-lattice antiferromagnets. Monte Carlo Methods and Applications, 20(3):217–??, September 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-3/mcma-2013-0027/mcma-2013-0027.xml.

Hordosy:1998:IPK

[HKHV98] G. Hordósy, A. Keresztúri, Cs. Hegedûs, and P. Vértes. Influence of the photoneutrons on the kinetics of beryllium reflected core of the Budapest Research Reactor. Monte Carlo Methods and Applications, 4(2):163–176, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
2/mcma.1998.4.2.163/mcma.1998.4.2.163.xml.

Hanousek:2012:IPJ

[HKN12] Jan Hanousek, Evzen Kocenda, and Jan Novotný. The identification of price jumps. Monte Carlo Methods and Applications, 18(1): 53–77, March 2012. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2012.18.issue-1/mcma-2011-0019/mcma-2011-0019.xml.

Hwang:2001:RDM

[HMG01] Chi-Ok Hwang, Michael Mascagni, and James A. Given. Rapid diffusion Monte Carlo algorithms for fluid dynamic permeability. Monte Carlo Methods and Applications, 7(3–4):213–222, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2001.7.
issue-3-4/mcma.2001.7.3-4.213/mcma.2001.7.3-4.213.xml.

Horntrop:2002:MCS

[Hor02] David J. Horntrop. Monte Carlo simulation of a random field model for transport. Monte Carlo Methods and Applications, 8(1):31–49, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2002.8.issue-1/mcma.2002.8.1.31/mcma.2002.8.1.31.xml.

Huang:2007:ERQ

[HPY07] M. L. Huang, M. Pollanen, and W. K. Yuen. An efficient randomized quasi-Monte Carlo algorithm for the Pareto distribution. Monte Carlo Methods and Applications, 13(1):1–20, ???? 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-1/mcma.2007.001/mcma.2007.001.xml.

Heritage:2002:LIT

[HR02] J. P. Heritage and L. C. G. Rogers. Large investors, takeovers, and the rule of law. Monte Carlo Methods and Applications, 8 (4):357–370, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2002.8.issue-4/mcma.2002.8.4.357/mcma.2002.
8.4.357.xml.

Halidias:2022:NAS

[HS22] Nikolaos Halidias and Ioannis S. Stamatiou. A note on the asymptotic stability of the semi-discrete method for stochastic differential equations. Monte Carlo Methods and Applications, 28(1): 13–25, February 15, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2022-2102/html.

Hamlin:2019:GEW

[HTKM19] Preston Hamlin, W. John Thrasher, Walid Keyrouz, and Michael Mascagni. Geometry entrapment in walk-on-subdomains. Monte Carlo Methods and Applications, 25(4):329–??, December 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-4/mcma-2019-2052/mcma-2019-2052.xml.

Hoel:2014:IAA

[HvSST14] Håkon Hoel, Erik von Schwerin, Anders Szepessy, and Raúl Tempone. Implementation and analysis of an adaptive multilevel Monte Carlo algorithm. Monte Carlo Methods and Applications, 20(1):1– ??, March 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2014.20.issue-1/mcma-2013-0014/mcma-2013-0014.xml.

Imamura:2014:NSB

[IIO14] Yuri Imamura, Yuta Ishigaki, and Toshiki Okumura. A numerical scheme based on semi-static hedging strategy. Monte Carlo Methods and Applications, 20(4):223–??, December 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-4/mcma-2014-0002/mcma-2014-0002.xml.

Ivanov:2000:CSS

[IK00] Vladimir M. Ivanov and Maxim L. Korenevski. On convergence of semi-statistical and projection-statistical methods for integral equations. Monte Carlo Methods and Applications, 6(4):297–322, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2000.6.issue-4/mcma.2000.6.4.297/mcma.2000.6.4.297.xml.

Ichikawa:2004:DVC

[IM04] Yuko Ichikawa and Makoto Mori. Discrepancy of van der Corput sequences generated by piecewise linear transformations. Monte Carlo Methods and Applications, 10(2):107–116, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-2/156939604777303217/156939604777303217.xml.

Imai:2013:CRN

[Ima13] Junichi Imai. Comparison of random number generators via Fourier transform. Monte Carlo Methods and Applications, 19(3):237–259, September 2013. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2013.19.issue-3/mcma-2013-0012/mcma-2013-0012.xml.

Iddi:2017:ECM

[IN17] Samuel Iddi and Esther O. Nwoko. Effect of covariate misspecifications in the marginalized zero-inflated Poisson model. Monte Carlo Methods and Applications, 23(2):111–??, June 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-2/mcma-2017-0106/mcma-2017-0106.xml.

Izydorczyk:2021:FBR

[IOR21] Lucas Izydorczyk, Nadia Oudjane, and Francesco Russo. A fully backward representation of semilinear PDEs applied to the control of thermostatic loads in power systems. Monte Carlo Methods and Applications, 27(4):347–371, October 21, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2095/html.

Iglesias:2017:UBC

[IP17] Emma M. Iglesias and Garry D. A. Phillips. The use of bias correction versus the jackknife when testing the mean reversion and long term mean parameters in continuous time models. Monte Carlo Methods and Applications, 23(3):159–??, September 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-3/mcma-2017-0111/mcma-2017-0111.xml.

Jimenez:2010:SSE

[JLH10] Edwin Jimenez, Nathan Lay, and M. Yousuff Hussaini. A systematic study of efficient sampling methods to quantify uncertainty in crack propagation and the Burgers equation. Monte Carlo Methods and Applications, 16(1):69–93, April 2010. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-1/mcma.2010.002/mcma.2010.002.xml.

Ji:2020:GVF

[JML20] Hao Ji, Michael Mascagni, and Yaohang Li. Gaussian variant of Freivalds’ algorithm for efficient and reliable matrix product verification. Monte Carlo Methods and Applications, 26(4):273–284, October 8, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/mcma-2020-2076/html.

Jourdain:2007:ERM

[JS07] Benjamin Jourdain and Mohamed Sbai. Exact retrospective Monte Carlo computation of arithmetic average Asian options. Monte Carlo Methods and Applications, 13(2):135–171, ???? 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-2/mcma.2007.008/mcma.2007.008.xml. See erratum

[JS10].

Jourdain:2010:EER

[JS10] Benjamin Jourdain and Mohamed Sbai. Erratum: Exact retrospective Monte Carlo computation of arithmetic average Asian options [MR2338086]. Monte Carlo Methods and Applications, 16(2): 191–193, July 2010. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2010.16.issue-2/mcma.2010.005/mcma.2010.005.xml.

See [JS07].

Jang:2019:ELS

[JWK19] Hanbyeol Jang, Jian Wang, and Junseok Kim. Equity-linked security pricing and Greeks at arbitrary intermediate times using Brownian bridge. Monte Carlo Methods and Applications, 25(4):291–??, December 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2019.25.issue-4/mcma-2019-2048/mcma-2019-2048.xml.

Kablukova:2005:IMN

[Kab05] E. G. Kablukova. Investigation of methods of numerical integration with optimal convergence speed. Monte Carlo Methods and Applications, 11(4):397–406, ???? 2005. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-4/156939605777438587/156939605777438587.xml.

Kanagawa:1995:EEE

[Kan95] Shuya Kanagawa. Error estimations for the Euler–Maruyama approximate solutions of stochastic differential equations. Monte Carlo Methods and Applications, 1(3):165–171, ???? 1995. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
3/mcma.1995.1.3.165/mcma.1995.1.3.165.xml.

Kashtanov:2017:SMM

[Kas17] Yuri Kashtanov. Stochastic mesh method for optimal stopping problems. Monte Carlo Methods and Applications, 23(2):121–??, June 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/view/j/mcma.
2017.23.issue-2/mcma-2017-0107/mcma-2017-0107.xml.

Kawai:2006:ISM

[Kaw06] Reiichiro Kawai. An importance sampling method based on the density transformation of Lévy processes. Monte Carlo Methods and Applications, 12(2):171–186, ???? 2006. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-2/156939606777488833/156939606777488833.xml.

Kawai:2007:AMC

[Kaw07] Reiichiro Kawai. Adaptive Monte Carlo variance reduction with two-time-scale stochastic approximation. Monte Carlo Methods and Applications, 13(3):197–217, August 2007. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-3/mcma.2007.010/mcma.2007.010.xml.

Khazen:1999:NVR

[KD99] Michael Khazen and Arie Dubi. A note on variance reduction methods in Monte Carlo applications to systems engineering and reliability. Monte Carlo Methods and Applications, 5(4):345–374, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1999.5.
issue-4/mcma.1999.5.4.345/mcma.1999.5.4.345.xml.

Khazen:2004:MCV

[KD04] Michael Khazen and Arie Dubi. Monte Carlo variance reduction in applications to systems reliability using phase space splitting. Monte Carlo Methods and Applications, 10(2):117–128, ????

2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-2/156939604777303226/156939604777303226.xml.

Keller:2004:TSR

[Kel04] Alexander Keller. Trajectory splitting by restricted replication. Monte Carlo Methods and Applications, 10(3–4):321–329, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2004.10.issue-3-4/mcma.2004.10.3-4.321/mcma.2004.10.3-
4.321.xml.

Khisamutdinov:2000:CBC

[Khi00] A. I. Khisamutdinov. On connection between “continuous time” and “direct simulation” Monte Carlo methods for Boltzmann equation and on some new approximate methods. Monte Carlo Methods and Applications, 6(4):323–340, ???? 2000. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
4/mcma.2000.6.4.323/mcma.2000.6.4.323.xml.

Kohatsu-Higa:1997:WRC

[KHO97] Arturo Kohatsu-Higa and Shigeyoshi Ogawa. Weak rate of convergence for an Euler scheme of nonlinear SDE’s. Monte Carlo Methods and Applications, 3(4):327–345, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
4/mcma.1997.3.4.327/mcma.1997.3.4.327.xml.

Kirillov:2009:ENF

[KK09] A. A. Kirillov and I. A. Kirillov. The exponential-normal form and its application to ultra high energy cascades investigation. Monte Carlo Methods and Applications, 15(2):107–133, August 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-2/mcma.2009.006/mcma.2009.006.xml.

Kozachenko:2013:LBT

[KKS13] Yury Kozachenko, Oleksandr Kurchenko, and Olga Synyavska. Levy–Baxter theorems for one class of non-Gaussian random fields. Monte Carlo Methods and Applications, 19(3):171–??, September 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-3/mcma-2013-0007/mcma-2013-0007.xml.

Kharroubi:2014:NAF

[KLP14] Idris Kharroubi, Nicolas Langrené, and Huyên Pham. A numerical algorithm for fully nonlinear HJB equations: An approach by control randomization. Monte Carlo Methods and Applications, 20(2): 145–??, June 2014. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2014.20.issue-2/mcma-2013-0024/mcma-2013-0024.xml.

Kurbanmuradov:2003:SLF

[KLR^{+}03] O. Kurbanmuradov, A. Levykin, U. Rannik, K. Sabelfeld, and T. Vesala. Stochastic Lagrangian footprint calculations over a surface with an abrupt change of roughness height. Monte Carlo Methods and Applications, 9(2):167–188, ???? 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
2/156939603322663330/156939603322663330.xml.

Kharroubi:2021:DML

[KLW21] Idris Kharroubi, Thomas Lim, and Xavier Warin. Discretization and machine learning approximation of BSDEs with a constraint on the Gains-process. Monte Carlo Methods and Applications, 27(1): 27–55, January 15, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-2080/html.

Komori:1995:SRT

[KM95] Yoshio Komori and Taketomo Mitsui. Stable ROW-type weak scheme for stochastic differential equations. Monte Carlo Methods and Applications, 1(4):279–300, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
4/mcma.1995.1.4.279/mcma.1995.1.4.279.xml.

Konakov:2002:ETE

[KM02] Valentin Konakov and Enno Mammen. Edgeworth type expansions for Euler schemes for stochastic differential equations. Monte Carlo Methods and Applications, 8(3):271–285, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
3/mcma.2002.8.3.271/mcma.2002.8.3.271.xml.

Kawai:2011:EDS

[KM11a] Reiichiro Kawai and Hiroki Masuda. Exact discrete sampling of finite variation tempered stable Ornstein–Uhlenbeck processes. Monte Carlo Methods and Applications, 17(3):279–300, September 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-3/mcma.2011.012/mcma.2011.012.xml.

Kozachenko:2011:PLD

[KM11b] Yu. V. Kozachenko and Yu. Yu. Mlavets. Probability of large deviations of sums of random processes from Orlicz space. Monte Carlo Methods and Applications, 17(2):155–168, June 2011. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-2/mcma.2011.007/mcma.2011.007.xml.

Kozachenko:2015:RAS

[KM15] Yuriy V. Kozachenko and Yuriy Y. Mlavets. Reliability and accuracy in the space L_{p}(T) for the calculation of integrals depending on a parameter by the Monte Carlo method. Monte Carlo Methods and Applications, 21(3):233–??, September 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-3/mcma-2015-0104/mcma-2015-0104.xml.

Kolyukhin:2022:SGR

[KM22] Dmitriy Kolyukhin and Alexander Minakov. Simulation of Gaussian random field in a ball. Monte Carlo Methods and Applications, 28(1):85–95, February 26, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2022-2108/html.

Karaivanova:2004:PQW

[KMS04] Aneta Karaivanova, Michael Mascagni, and Nikolai A. Simonov. Parallel quasirandom walks on the boundary. Monte Carlo Methods and Applications, 10(3–4):311–319, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.311/mcma.2004.10.3-4.311.xml.

Kosina:2004:SSD

[KNS04] H. Kosina, M. Nedjalkov, and S. Selberherr. Solution of the space-dependent Wigner equation using a particle model. Monte Carlo Methods and Applications, 10(3–4):359–368, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.359/mcma.2004.10.3-4.359.
xml.

Kolyukhin:2018:GSA

[Kol18] Dmitriy Kolyukhin. Global sensitivity analysis for a stochastic flow problem. Monte Carlo Methods and Applications, 24(4):263–270, December 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2018.24.issue-4/mcma-2018-2022/mcma-2018-2022.xml.

Kolyukhin:2020:SMT

[Kol20] Dmitriy Kolyukhin. Statistical modeling of three-dimensional fractal point sets with a given spatial probability distribution. Monte Carlo Methods and Applications, 26(3):245–252, July 16, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2020-2066/html.

Kolyukhin:2021:GSA

[Kol21] Dmitriy Kolyukhin. Global sensitivity analysis of statistical models by double randomization method. Monte Carlo Methods and Applications, 27(4):341–346, October 27, 2021. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2096/html.

Kurbanmuradov:2001:ARD

[KOSY01] O. A. Kurbanmuradov, S. A. Orszag, K. K. Sabelfeld, and P. K. Yeung. Analysis of relative dispersion of two particles by Lagrangian stochastic models and DNS methods. Monte Carlo Methods and Applications, 7(3–4):245–263, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.245/mcma.2001.7.3-4.245.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Kubilius:2002:RWC

[KP02] Kestutis Kubilius and Eckhard Platen. Rate of weak convergence of the Euler approximation for diffusion processes with jumps.

Monte Carlo Methods and Applications, 8(1):83–96, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2002.8.
issue-1/mcma.2002.8.1.83/mcma.2002.8.1.83.xml.

Kopylov:1996:MCM

[KPSZ96] Yu. N. Kopylov, I. S. Postnova, V. A. Shpack, and G. S. Zinchenko. Monte Carlo modeling of radioactive fallout under land-based nuclear bursts. Monte Carlo Methods and Applications, 2(2):145–??, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
1996.2.issue-2/mcma.1996.2.2.145/mcma.1996.2.2.145.xml.

Kozachenko:2018:SGF

[KPV18] Yuriy Kozachenko, Anatolii Pashko, and Olga Vasylyk. Simulation of generalized fractional Brownian motion in C([0, T]). Monte Carlo Methods and Applications, 24(3):179–192, September 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-3/mcma-2018-0016/mcma-2018-0016.xml.

Kramer:2001:RSM

[Kra01] Peter R. Kramer. A review of some Monte Carlo simulation methods for turbulent systems. Monte Carlo Methods and Applications, 7(3–4):229–243, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.229/mcma.2001.7.3-4.229.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Khan:2017:CBC

[KRSJ17] Naushad Mamode Khan, Wasseem Rumjaun, Yuvraj Sunecher, and Vandna Jowaheer. Computing with bivariate COM-Poisson model under different copulas. Monte Carlo Methods and Applications, 23(2):131–??, June 2017. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2017.23.issue-2/
mcma-2017-0103/mcma-2017-0103.xml.

Kurbanmuradov:1999:DAM

[KRSV99] O. Kurbanmuradov, U. Rannik, K. Sabelfeld, and T. Vesala. Direct and adjoint Monte Carlo algorithms for the footprint problem. Monte Carlo Methods and Applications, 5(2):85–111, ????

1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1999.5.
issue-2/mcma.1999.5.2.85/mcma.1999.5.2.85.xml.

Khisamutdinov:1995:MCF

[KS95a] A. I. Khisamutdinov and L. L. Sidorenko. Monte Carlo fictitious collision algorithms for nonlinear Boltzmann equation. Monte Carlo Methods and Applications, 1(3):221–240, ???? 1995. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
3/mcma.1995.1.3.221/mcma.1995.1.3.221.xml.

Kurbanmuradov:1995:SLM

[KS95b] O. A. Kurbanmuradov and K. K. Sabel′fel′d. Stochastic Lagrangian models of relative dispersion of a pair of fluid particles in turbulent flows. Monte Carlo Methods and Applications, 1(2):101–136, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1995.1.
issue-2/mcma.1995.1.2.101/mcma.1995.1.2.101.xml.

Kanagawa:2000:SAR

[KS00] S. Kanagawa and Y. Saisho. Strong approximation of reflecting Brownian motion using penalty method and its application to computer simulation. Monte Carlo Methods and Applications, 6 (2):105–114, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2000.6.issue-2/mcma.2000.6.2.105/mcma.2000.
6.2.105.xml.

Kolodko:2001:SLM

[KS01] A. A. Kolodko and K. K. Sabelfeld. Stochastic Lagrangian model for spatially inhomogeneous Smoluchowski equation governing coagulating and diffusing particles. Monte Carlo Methods and Applications, 7(3–4):223–228, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.223/mcma.2001.7.3-4.223.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Kolodko:2003:SPM

[KS03] A. Kolodko and K. Sabelfeld. Stochastic particle methods for Smoluchowski coagulation equation: variance reduction and error estimations. Monte Carlo Methods and Applications, 9(4):315–339, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2003.9.issue-4/156939603322601950/156939603322601950.xml.

Kolodko:2004:UBB

[KS04a] A. Kolodko and J. Schoenmakers. Upper bounds for Bermudan style derivatives. Monte Carlo Methods and Applications, 10(3–4): 331–343, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.331/
mcma.2004.10.3-4.331.xml.

Kolyukhin:2004:SEM

[KS04b] Dmitry Kolyukhin and Karl Sabelfeld. Stochastic Eulerian model for the flow simulation in porous media. Unconfined aquifers. Monte Carlo Methods and Applications, 10(3–4):345–357, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.345/mcma.2004.10.3-4.345.
xml.

Kuzmin:2004:SMF

[KS04c] G. A. Kuzmin and O. N. Soboleva. Subgrid modeling of filtration in a porous medium with multiscale log-stable permeability. Monte Carlo Methods and Applications, 10(3–4):369–376, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.369/mcma.2004.10.3-4.369.
xml.

Kolyukhin:2005:SFS

[KS05] Dmitry Kolyukhin and Karl Sabelfeld. Stochastic flow simulation in 3D porous media. Monte Carlo Methods and Applications, 11 (1):15–37, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2005.11.issue-1/1569396054027292/1569396054027292.
xml.

Kahl:2006:BMM

[KS06a] Christian Kahl and Henri Schurz. Balanced Milstein methods for ordinary SDEs. Monte Carlo Methods and Applications, 12(2):143– 170, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/

mcma.2006.12.issue-2/156939606777488842/156939606777488842.
xml.

Kurbanmuradov:2006:EBP

[KS06b] O. Kurbanmuradov and K. Sabelfeld. Exponential bounds for the probability deviations of sums of random fields. Monte Carlo Methods and Applications, 12(3–4):211–229, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-3/156939606778705218/156939606778705218.xml.

Kurbanmuradov:2006:SSF

[KS06c] O. Kurbanmuradov and K. Sabelfeld. Stochastic spectral and Fourier-wavelet methods for vector Gaussian random fields. Monte Carlo Methods and Applications, 12(5–6):395–445, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-5/156939606779329080/156939606779329080.xml.

Kozachenko:2014:CHT

[KS14] Yuriy V. Kozachenko and Mykola P. Sergiienko. The criterion of hypothesis testing on the covariance function of a Gaussian stochastic process. Monte Carlo Methods and Applications, 20(2):137–??, June 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2014.20.issue-2/mcma-2013-0023/mcma-2013-0023.xml.

Kolyukhin:2015:SSP

[KS15] Dmitriy Kolyukhin and Karl K. Sabelfeld. Stochastic small perturbation based simulation technique for solving isotropic elastostatics equations. Monte Carlo Methods and Applications, 21(2):153–??, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2015.21.issue-2/mcma-2014-0016/mcma-2014-0016.xml.

Kong:2016:NPG

[KS16] Rong Kong and Jerome Spanier. A new proof of geometric convergence for the adaptive generalized weighted analog sampling (GWAS) method. Monte Carlo Methods and Applications, 22(3):161–??, September 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-3/
mcma-2016-0110/mcma-2016-0110.xml.

Kloeden:2011:EPS

[KSC11] Peter E. Kloeden and Carlos Sanz-Chacón. Efficient price sensitivity estimation of financial derivatives by weak derivatives. Monte Carlo Methods and Applications, 17(1):47–75, March 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-1/mcma.2011.001/mcma.2011.001.xml.

Kurbanmuradov:1997:SLMb

[KSK97] O. Kurbanmuradov, K. Sabelfeld, and D. Koluhin. Stochastic Lagrangian models for two-particle motion in turbulent flows. Numerical results. Monte Carlo Methods and Applications, 3(3):199–223, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
1997.3.issue-3/mcma.1997.3.3.199/mcma.1997.3.3.199.xml.

Kabanikhin:2015:NSI

[KSNS15] Sergey I. Kabanikhin, Karl K. Sabelfeld, Nikita S. Novikov, and Maxim A. Shishlenin. Numerical solution of an inverse problem of coefficient recovering for a wave equation by a stochastic projection methods. Monte Carlo Methods and Applications, 21(3):189–??, September 2015. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-3/mcma-2015-0103/mcma-2015-0103.xml.

Kablukova:2020:DVG

[KSPZ20] Evgenia Kablukova, Karl Sabelfeld, Dmitrii Y. Protasov, and Konstantin S. Zhuravlev. Drift velocity in GaN semiconductors: Monte Carlo simulation and comparison with experimental measurements. Monte Carlo Methods and Applications, 26(4):263–271, October 30, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2077/html.

Kurbanmuradov:2003:LSM

[KSSV03] O. Kurbanmuradov, K. Sabelfeld, O. F. Smidts, and H. Vereecken. A Lagrangian stochastic model for the transport in statistically homogeneous porous media. Monte Carlo Methods and Applications, 9(4):341–366, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2003.9.issue-4/156939603322601969/156939603322601969.
xml.

Karlsson:2011:TAG

[KT11a] Jesper Karlsson and Raúl Tempone. Towards automatic global error control: Computable weak error expansion for the tau-leap method. Monte Carlo Methods and Applications, 17(3):233–278, September 2011. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2011.17.issue-3/mcma.2011.011/mcma.2011.011.xml.

Kiessling:2011:DAL

[KT11b] Jonas Kiessling and Raúl Tempone. Diffusion approximation of Lévy processes with a view towards finance. Monte Carlo Methods and Applications, 17(1):11–45, March 2011. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2011.
17.issue-1/mcma.2011.003/mcma.2011.003.xml.

Kurbanmuradov:1995:SMR

[Kur95a] O. A. Kurbanmuradov. A 3D stochastic model of relative dispersion of particle pairs in local-isotropic turbulence. Monte Carlo Methods and Applications, 1(4):301–324, ???? 1995. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
4/mcma.1995.1.4.301/mcma.1995.1.4.301.xml.

Kurbanmuradov:1995:NLM

[Kur95b] O. A. Kurbanmuradov. A new Lagrangian model of two-particle relative turbulent dispersion. Monte Carlo Methods and Applications, 1(2):83–100, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1995.1.issue-2/mcma.1995.1.2.
83/mcma.1995.1.2.83.xml.

Kurbanmuradov:1997:SLMa

[Kur97] O. A. Kurbanmuradov. Stochastic Lagrangian models for twoparticle relative dispersion in high-Reynolds number turbulence. Monte Carlo Methods and Applications, 3(1):37–52, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1997.3.
issue-1/mcma.1997.3.1.37/mcma.1997.3.1.37.xml.

Kolodko:1997:CNT

[KW97] Anastasya A. Kolodko and Wolfgang Wagner. Convergence of a Nanbu type method for the Smoluchowski equation. Monte Carlo Methods and Applications, 3(4):255–273, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
4/mcma.1997.3.4.255/mcma.1997.3.4.255.xml.

Kraft:2002:ESC

[KW02] Markus Kraft and Wolfgang Wagner. An efficient stochastic chemistry approximation for the PDF transport equation. Monte Carlo Methods and Applications, 8(4):371–394, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
4/mcma.2002.8.4.371/mcma.2002.8.4.371.xml.

Lay:2018:IMM

[LCRK18] Harold A. Lay, Zane Colgin, Viktor Reshniak, and Abdul Q. M. Khaliq. On the implementation of multilevel Monte Carlo simulation of the stochastic volatility and interest rate model using multi-GPU clusters. Monte Carlo Methods and Applications, 24(4):309–321, December 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-4/
mcma-2018-2025/mcma-2018-2025.xml.

Lejay:2001:WSS

[Lej01] Antoine Lejay. Weak solution of semi-linear PDE, BSDE and homogenization. Monte Carlo Methods and Applications, 7(3–4):265– 272, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-3-4/mcma.2001.7.3-4.265/mcma.2001.

7.3-4.265.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Lejay:2003:SDG

[Lej03] Antoine Lejay. Simulating a diffusion on a graph. Application to reservoir engineering. Monte Carlo Methods and Applications, 9(3): 241–255, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2003.9.issue-3/156939603322729003/156939603322729003.
xml.

Lejay:2004:MCM

[Lej04] Antoine Lejay. Monte Carlo methods for fissured porous media: a gridless approach. Monte Carlo Methods and Applications, 10(3– 4):385–392, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.385/
mcma.2004.10.3-4.385.xml.

Leobacher:2006:SSQ

[Leo06] G. Leobacher. Stratified sampling and quasi-Monte Carlo simulation of Lévy processes. Monte Carlo Methods and Applications, 12 (3–4):231–238, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2006.12.issue-3/156939606778705155/156939606778705155.
xml.

Levada:2016:IGS

[Lev16] Alexandre L. Levada. Information geometry, simulation and complexity in Gaussian random fields. Monte Carlo Methods and Applications, 22(2):81–??, June 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-2/
mcma-2016-0107/mcma-2016-0107.xml.

Leydold:2004:STD

[LH04] Josef Leydold and Wolfgang Hörmann. Smoothed transformed density rejection. Monte Carlo Methods and Applications, 10(3–4): 393–401, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.393/
mcma.2004.10.3-4.393.xml.

Likhoded:1998:SAS

[Lik98] N. A. Likhoded. Systolic arrays for the solution of systems of linear algebraic equations by Monte Carlo method. Monte Carlo Methods and Applications, 4(1):17–32, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
1/mcma.1998.4.1.17/mcma.1998.4.1.17.xml.

Lin:2006:CMS

[Lin06] Ting Hsiang Lin. A comparison of model selection indices for nested latent class models. Monte Carlo Methods and Applications, 12 (3–4):239–259, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2006.12.issue-3/156939606778705164/156939606778705164.
xml.

Lukka:2002:UGO

[LK02] Tuomas J. Lukka and Janne V. Kujala. Using genetic operators to speed up Markov chain Monte Carlo integration. Monte Carlo Methods and Applications, 8(1):51–71, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
1/mcma.2002.8.1.51/mcma.2002.8.1.51.xml.

Lapeyre:2011:FAM

[LL11] Bernard Lapeyre and Jérôme Lelong. A framework for adaptive Monte Carlo procedures. Monte Carlo Methods and Applications, 17(1):77–98, March 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2011.17.issue-1/mcma.2011.002/mcma.2011.002.
xml.

Labart:2013:PAS

[LL13] Céline Labart and Jérôme Lelong. A parallel algorithm for solving BSDEs. Monte Carlo Methods and Applications, 19(1):11–??, March 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2013.19.issue-1/mcma-2013-0001/mcma-2013-0001.xml.

Lee:2014:MAE

[LL14] Seung-Hwan Lee and Eun-Joo Lee. A martingale approach to estimating confidence band with censored data. Monte Carlo Methods and Applications, 20(4):237–??, December 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-4/mcma-2014-0003/mcma-2014-0003.xml.

Lee:2020:WLR

[LL20] Seung-Hwan Lee and Eun-Joo Lee. A weighted log-rank test for comparing two survival curves. Monte Carlo Methods and Applications, 26(3):253–262, April 15, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2020-2064/html.

Lapeyre:2021:NNR

[LL21] Bernard Lapeyre and Jérôme Lelong. Neural network regression for Bermudan option pricing. Monte Carlo Methods and Applications, 27(3):227–247, July 1, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-2091/html.

Bris:2012:MFP

[LLLP12] Claude Le Bris, Tony Lelièvre, Mitchell Luskin, and Danny Perez. A mathematical formalization of the parallel replica dynamics. Monte Carlo Methods and Applications, 18(2):119–??, June 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-2/mcma-2012-0003/mcma-2012-0003.xml.

LeBris:2016:SQS

[LLM16] Claude Le Bris, Frédéric Legoll, and William Minvielle. Special quasirandom structures: A selection approach for stochastic homogenization. Monte Carlo Methods and Applications, 22(1):25–??, March 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2016.22.issue-1/mcma-2016-0101/mcma-2016-0101.xml.

Li:2005:GBQ

[LM05] Yaohang Li and Michael Mascagni. Grid-based quasi-Monte Carlo applications. Monte Carlo Methods and Applications, 11(1):39–55, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2005.11.issue-1/1569396054027265/1569396054027265.xml.

Lepingle:2004:ASM

[LN04] Dominique Lepingle and Thi Thao Nguyen. Approximating and simulating multivalued stochastic differential equations. Monte Carlo Methods and Applications, 10(2):129–152, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-2/156939604777303244/156939604777303244.xml.

Li:2015:OSP

[LNO15] Kai Li, Kaj Nyström, and Marcus Olofsson. Optimal switching problems under partial information. Monte Carlo Methods and Applications, 21(2):91–??, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2015.21.issue-2/mcma-2014-0013/
mcma-2014-0013.xml.

LeCavil:2018:MCA

[LOR18] Anthony Le Cavil, Nadia Oudjane, and Francesco Russo. Monte-Carlo algorithms for a forward Feynman–Kac-type representation for semilinear nonconservative partial differential equations. Monte Carlo Methods and Applications, 24(1):55–70, March 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-1/mcma-2018-0005/mcma-2018-0005.xml.

Lang:2011:FSG

[LP11] Annika Lang and Jürgen Potthoff. Fast simulation of Gaussian random fields. Monte Carlo Methods and Applications, 17(3):195– 214, September 2011. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2011.17.issue-3/mcma.2011.009/mcma.2011.009.xml.

See erratum [LP13].

Laruelle:2012:SAA

[LP12] Sophie Laruelle and Gilles Pagès. Stochastic approximation with averaging innovation applied to finance. Monte Carlo Methods and Applications, 18(1):1–51, March 2012. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-1/mcma-2011-0018/mcma-2011-0018.xml.

Lang:2013:EFS

[LP13] Annika Lang and Jürgen Potthoff. Erratum: Fast simulation of Gaussian random fields [Monte Carlo Methods Appl. 17 (2011), 195–214]. Monte Carlo Methods and Applications, 19(1):73–??, March 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2013.19.issue-1/mcma-2013-0003/mcma-2013-0003.xml.

See [LP11].

Larcher:2003:AAO

[LPT03] Gerhard Larcher, Martin Predota, and Robert F. Tichy. Arithmetic average options in the hyperbolic model. Monte Carlo Methods and Applications, 9(3):227–239, ???? 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
3/156939603322728996/156939603322728996.xml.

Li:1997:ATE

[LS97] Liming Li and Jerome Spanier. Approximation of transport equations by matrix equations and sequential sampling. Monte Carlo Methods and Applications, 3(3):171–198, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
3/mcma.1997.3.3.171/mcma.1997.3.3.171.xml.

Lecot:2004:CQM

[LT04] Christian Lécot and Bruno Tuffin. Comparison of quasi-Monte Carlo-based methods for simulation of Markov chains. Monte Carlo Methods and Applications, 10(3–4):377–384, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.377/mcma.2004.10.3-4.377.
xml.

Lecot:2008:QSS

[LT08] C. Lécot and A. Tarhini. A quasi-stochastic simulation of the general dynamics equation for aerosols. Monte Carlo Methods and Applications, 13(5–6):369–388, January 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
13.issue-5-6/mcma.2007.020/mcma.2007.020.xml.

Lesage:2001:CDM

[LTD01] Stéphane Lesage, Jean-Yves Tourneret, and Petar M. Djurić. Classification of digital modulations using MCMC methods. Monte Carlo Methods and Applications, 7(3–4):273–282, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.273/mcma.2001.7.3-4.273.xml.

Lambert:2010:GST

[LW10] Tiffany A. Lambert and Paula A. Whitlock. Generalizing Sudoku to three dimensions. Monte Carlo Methods and Applications, 16(3–4):251–263, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2010.16.issue-3-4/mcma.
2010.018/mcma.2010.018.xml.

Liu:2018:BEO

[LWC18] Baisen Liu, Liangliang Wang, and Jiguo Cao. Bayesian estimation of ordinary differential equation models when the likelihood has multiple local modes. Monte Carlo Methods and Applications, 24(2):117–127, June 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-2/
mcma-2018-0010/mcma-2018-0010.xml.

Makhotkin:2015:SRV

[Mak15] Oleg A. Makhotkin. Simulation of random variates with the Morgenstern distribution. Monte Carlo Methods and Applications, 21(4):325–??, December 2015. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2015.21.issue-4/
mcma-2015-0107/mcma-2015-0107.xml.

Malyutin:2007:AMV

[Mal07] V. Malyutin. On approximation of matrix valued functional integrals. Monte Carlo Methods and Applications, 13(4):287–298, November 20, 2007. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2007.13.issue-4/mcma.2007.016/mcma.2007.016.xml.

Mantalos:2003:BBG

[Man03] Panagiotis Mantalos. Bootstrapping the Breusch–Godfrey autocorrelation test for a single equation dynamic model: Bootstrapping the restricted vs. unrestricted model. Monte Carlo Methods and Applications, 9(3):257–269, ???? 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
3/156939603322729012/156939603322729012.xml.

Marxen:2010:MMC

[Mar10] Henning Marxen. The multilevel Monte Carlo method used on a Lévy driven SDE. Monte Carlo Methods and Applications, 16(2): 167–190, July 2010. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2010.16.issue-2/mcma.2010.007/mcma.2010.007.xml.

Mathe:1999:NIU

[Mat99] Peter Mathé. Numerical integration using Markov chains. Monte Carlo Methods and Applications, 5(4):325–343, ???? 1999. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
4/mcma.1999.5.4.325/mcma.1999.5.4.325.xml.

Mansor:2006:PRM

[MBK06] Siti Nurleena Abu Mansor, Adam Baharum, and Anton Abdulbasah Kamil. Portfolio resampling in Malaysian equity market. Monte Carlo Methods and Applications, 12(3–4):261–269, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-3/156939606778705146/156939606778705146.xml.

Mascagni:2004:SHS

[MC04] Michael Mascagni and Hongmei Chi. On the scrambled Halton sequence. Monte Carlo Methods and Applications, 10(3–4): 435–442, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.435/
mcma.2004.10.3-4.435.xml.

Marozzi:2020:ITP

[MC20] Marco Marozzi and Shovan Chowdhury. An index of teaching performance based on students’ feedback. Monte Carlo Methods and Applications, 26(2):83–91, April 15, 2020. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2020-2059/html.

McLeish:2011:GMD

[McL11] Don McLeish. A general method for debiasing a Monte Carlo estimator. Monte Carlo Methods and Applications, 17(4):301–315, December 2011. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2011.17.issue-4/mcma.2011.013/mcma.2011.013.xml.

Muller:2020:NNA

[MDMS20] Christian Müller, Holger Diedam, Thomas Mrziglod, and Andreas Schuppert. A neural network assisted Metropolis adjusted Langevin algorithm. Monte Carlo Methods and Applications, 26(2):93–111, April 15, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/mcma-2020-2060/html.

Missov:2009:ISP

[ME09] Trifon I. Missov and Sergey M. Ermakov. On importance sampling in the problem of global optimization. Monte Carlo Methods and Applications, 15(2):135–144, August 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-2/mcma.2009.007/mcma.2009.007.xml.

Makarov:2010:ESB

[MG10] Roman N. Makarov and Devin Glew. Exact simulation of Bessel diffusions. Monte Carlo Methods and Applications, 16(3–4):283–306, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2010.16.issue-3-4/mcma.2010.010/mcma.2010.010.xml.

Mascagni:2012:PRN

[MH12] Michael Mascagni and Lin-Yee Hin. Parallel random number generators in Monte Carlo derivative pricing: An application-based test. Monte Carlo Methods and Applications, 18(2):161–??, June 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-2/mcma-2012-0005/mcma-2012-0005.xml.

Mascagni:2013:PPR

[MH13] Michael Mascagni and Lin-Yee Hin. Parallel pseudo-random number generators: A derivative pricing perspective with the Heston stochastic volatility model. Monte Carlo Methods and Applications, 19(2):77–??, July 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2013.19.issue-2/mcma-2013-0006/mcma-2013-0006.
xml.

Minier:2001:PAT

[Min01] Jean-Pierre Minier. Probabilistic approach to turbulent twophase flows modelling and simulation: theoretical and numerical issues. Monte Carlo Methods and Applications, 7(3–4):295– 310, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-3-4/mcma.2001.7.3-4.295/mcma.2001.

7.3-4.295.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Missov:2007:IEU

[Mis07] Trifon I. Missov. Integral evaluation using the Δ^{2}-distribution. Simulation and illustration. Monte Carlo Methods and Applications, 13(3):219–225, August 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2007.13.issue-3/mcma.2007.011/mcma.2007.011.
xml.

Mosbach:2006:ESO

[MK06] S. Mosbach and M. Kraft. Explicit stochastic ODE solution methods applied to high-temperature combustion. Monte Carlo Methods and Applications, 12(1):19–45, ???? 2006. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-1/156939606776886670/156939606776886670.xml.

Mascagni:2001:QMC

[MKL01] Michael Mascagni, Aneta Karaivanova, and Yaohang Li. A quasi-Monte Carlo method for elliptic boundary value problems. Monte Carlo Methods and Applications, 7(3–4):283–293, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.283/mcma.2001.7.3-4.283.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Meglinsky:2000:AMC

[MM00] I. V. Meglinsky and S. J. Matcher. The application of the Monte Carlo technique for estimation of the detector depth sensitivity for the skin oxygenation measurements. Monte Carlo Methods and Applications, 6(1):15–25, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
1/mcma.2000.6.1.15/mcma.2000.6.1.15.xml.

Mori:2012:DSG

[MM12] Makoto Mori and Masaki Mori. Dynamical system generated by algebraic method and low discrepancy sequences. Monte Carlo Methods and Applications, 18(4):327–??, December 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-4/mcma-2012-0012/mcma-2012-0012.xml.

Malekzadeh:2020:CCI

[MM20] Ahad Malekzadeh and Seyed Mahdi Mahmoudi. Constructing a confidence interval for the ratio of normal distribution quantiles. Monte Carlo Methods and Applications, 26(4):325–334, August 5, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2070/html.

mongaMade:2004:AAM

[mMSD04] M. Magolu monga Made, O. F. Smidts, and A. Dubus. Adaptive adjoint Monte Carlo simulation for the uncertainty. Monte Carlo Methods and Applications, 10(3–4):403–413, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.403/mcma.2004.10.3-4.403.
xml.

Mori:1998:LDS

[Mor98] Makoto Mori. Low discrepancy sequences generated by piecewise linear maps. Monte Carlo Methods and Applications, 4(2):141–162, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
1998.4.issue-2/mcma.1998.4.2.141/mcma.1998.4.2.141.xml.

Mori:1999:DSG

[Mor99] Makoto Mori. Discrepancy of sequences generated by piecewise monotone maps. Monte Carlo Methods and Applications, 5(1):55– 68, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.1999.5.issue-1/mcma.1999.5.1.55/mcma.1999.5.1.55.
xml.

Mori:2002:CTD

[Mor02] Makoto Mori. Construction of two dimensional low discrepancy sequences. Monte Carlo Methods and Applications, 8(2):159–169, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2002.8.issue-2/mcma.2002.8.2.159/mcma.2002.8.2.159.xml.

Mori:2004:DSG

[Mor04] Makoto Mori. Discrepancy of sequences generated by dynamical system. Monte Carlo Methods and Applications, 10(3–4): 455–459, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.455/
mcma.2004.10.3-4.455.xml.

Mori:2005:CTD

[Mor05] Makoto Mori. Construction of three-dimensional low discrepancy sequences. Monte Carlo Methods and Applications, 11(2):163–174, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2005.11.issue-2/156939605777585962/156939605777585962.
xml.

Mori:2008:SPF

[Mor08] Makoto Mori. Spectra of Perron–Frobenius operator and new construction of two dimensional low discrepancy sequences. Monte Carlo Methods and Applications, 14(1):53–74, May 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-1/mcma.2008.003/mcma.2008.003.xml.

Moskvin:2002:MTR

[MP02] Vadim Moskvin and Lech Papiez. Method of trajectory rotation as a Monte Carlo variance reduction technique. Monte Carlo Methods and Applications, 8(3):287–298, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
3/mcma.2002.8.3.287/mcma.2002.8.3.287.xml.

Maire:2012:RED

[MP12] Sylvain Maire and Cyril Prissette. A restarted estimation of distribution algorithm for solving sudoku puzzles. Monte Carlo Methods and Applications, 18(2):147–??, June 2012. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-2/mcma-2012-0004/mcma-2012-0004.xml.

Minier:2003:WFS

[MPC03] Jean-Pierre Minier, Eric Peirano, and Sergio Chibbaro. Weak first- and second-order numerical schemes for stochastic differential equations appearing in Lagrangian two-phase flow modeling. Monte Carlo Methods and Applications, 9(2):93–133, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2003.9.
issue-2/156939603322663312/156939603322663312.xml.

Marseguerra:2004:NMT

[MPZP04] M. Marseguerra, E. Padovani, E. Zio, and F. Giacobbo E. Patelli. A nuclear measurement technique of water penetration in concrete barriers. Monte Carlo Methods and Applications, 10(3–4): 415–422, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.415/
mcma.2004.10.3-4.415.xml.

Mascagni:2014:HPC

[MQH14] Michael Mascagni, Yue Qiu, and Lin-Yee Hin. High performance computing in quantitative finance: A review from the pseudo-random number generator perspective. Monte Carlo Methods and Applications, 20(2):101–120, June 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-2/mcma-2013-0020/mcma-2013-0020.xml.

Mazzolo:2004:MCS

[MR04] Alain Mazzolo and Benoît Roesslinger. Monte-Carlo simulation of the chord length distribution function across convex bodies, non-convex bodies and random media. Monte Carlo Methods and Applications, 10(3–4):443–454, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2004.10.issue-3-4/mcma.
2004.10.3-4.443/mcma.2004.10.3-4.443.xml.

Metzler:2014:RES

[MS14] Adam Metzler and Alexandre Scott. Rare event simulation for diffusion processes via two-stage importance sampling. Monte Carlo Methods and Applications, 20(2):77–??, June 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-2/mcma-2013-0019/mcma-2013-0019.xml.

Matoussi:2016:NCB

[MS16] Anis Matoussi and Wissal Sabbagh. Numerical computation for backward doubly SDEs with random terminal time. Monte Carlo Methods and Applications, 22(3):229–??, September 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-3/mcma-2016-0111/mcma-2016-0111.xml.

Maire:2008:SNS

[MT08] Sylvain Maire and Etienne Tanré. Some new simulations schemes for the evaluation of Feynman–Kac representations. Monte Carlo Methods and Applications, 14(1):29–51, May 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-1/mcma.2008.002/mcma.2008.002.xml.

Maire:2013:MCA

[MT13] Sylvain Maire and Etienne Tanré. Monte Carlo approximations of the Neumann problem. Monte Carlo Methods and Applications, 19(3):201–??, September 2013. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-3/mcma-2013-0010/mcma-2013-0010.xml.

Muller:2018:MCM

[MWMS18] Christian Müller, Fabian Weysser, Thomas Mrziglod, and Andreas Schuppert. Markov-Chain Monte-Carlo methods and nonidentifiabilities. Monte Carlo Methods and Applications, 24(3):203– 214, September 2018. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL https://www.degruyter.com/view/
j/mcma.2018.24.issue-3/mcma-2018-0018/mcma-2018-0018.xml.

Mascagni:2009:SSS

[MY09] Michael Mascagni and Haohai Yu. Scrambled Sobol′ sequences via permutation. Monte Carlo Methods and Applications, 15(4):311– 332, December 2009. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2009.15.issue-4/mcma.2009.017/mcma.2009.017.xml.

Marseguerra:1998:WUF

[MZ98] M. Marseguerra and E. Zio. Weight updating in forced Monte Carlo approach to dynamic PSA. Monte Carlo Methods and Applications, 4(4):359–373, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1998.4.issue-4/mcma.1998.4.4.359/mcma.1998.
4.4.359.xml.

Marseguerra:2004:SAR

[MZB04] Mario Marseguerra, Enrico Zio, and Francesco Bosi. System availability and reliability analysis by direct Monte Carlo with biasing. Monte Carlo Methods and Applications, 10(3–4):423– 434, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.423/
mcma.2004.10.3-4.423.xml.

Nadarajah:2007:CTE

[Nad07] Saralees Nadarajah. Comparison of time-to-event data for clinical trials. Monte Carlo Methods and Applications, 13(1):21–35, ???? 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-1/mcma.2007.002/mcma.2007.002.xml.

Nadarajah:2008:CPP

[Nad08a] Saralees Nadarajah. Computing percentage points of the largest among Student’s t random variables. Monte Carlo Methods and Applications, 14(1):75–84, May 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2008.14.issue-1/mcma.2008.004/
mcma.2008.004.xml.

Nadarajah:2008:SDG

[Nad08b] Saralees Nadarajah. Skewed distributions generated by the Student’s t kernel. Monte Carlo Methods and Applications, 13(5– 6):389–404, January 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2008.13.issue-5-6/mcma.2007.021/mcma.2007.
021.xml.

Nakao:1997:EEL

[Nak97] Hajime Nakao. Exact expression of Lagrangian velocity autocorrelation function in isotropic homogeneous turbulence. Monte Carlo Methods and Applications, 3(3):225–240, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
3/mcma.1997.3.3.225/mcma.1997.3.3.225.xml.

Nakao:1998:TDE

[Nak98] Hajime Nakao. Time dependence of Eulerian velocity correlation tensor spectrum in isotropic homogeneous turbulence. Monte Carlo Methods and Applications, 4(2):113–125, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
2/mcma.1998.4.2.113/mcma.1998.4.2.113.xml.

Nedjalkov:2004:OSM

[NAKS04] M. Nedjalkov, E. Atanassov, H. Kosina, and S. Selberherr. Operator-split method for variance reduction in stochastic solutions of the Wigner equation. Monte Carlo Methods and Applications, 10(3–4):461–468, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2004.10.issue-3-4/mcma.
2004.10.3-4.461/mcma.2004.10.3-4.461.xml.

Naono:1995:CCN

[Nao95] Ken Naono. Comparative computations of non-parametric density estimation between some kernel method and the wavelet method. Monte Carlo Methods and Applications, 1(2):147–163, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1995.1.
issue-2/mcma.1995.1.2.147/mcma.1995.1.2.147.xml.

Nasroallah:2019:KDF

[NB19] Abdelaziz Nasroallah and Mohamed Yasser Bounnite. A kind of dual form for coupling from the past algorithm, to sample from Markov chain steady-state probability. Monte Carlo Methods and Applications, 25(4):317–??, December 2019. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-4/mcma-2019-2050/mcma-2019-2050.xml.

Nagy:2020:MMC

[NEBW20] Shady Ahmed Nagy, Mohamed A. El-Beltagy, and Mohamed Wafa. Multilevel Monte Carlo by using the Halton sequence. Monte Carlo Methods and Applications, 26(3):193–203, April 17, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2020-2065/html.

Nekrutkin:2003:KPS

[Nek03] V. Nekrutkin. Kac particle systems with free motion and equations of Boltzmann type. Monte Carlo Methods and Applications, 9(1): 13–25, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2003.9.issue-1/156939603322587434/156939603322587434.
xml.

Nekrutkin:2016:CBF

[Nek16] Vladimir Nekrutkin. On the complexity of binary floating point pseudorandom generation. Monte Carlo Methods and Applications, 22(2):109–??, June 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-2/
mcma-2016-0105/mcma-2016-0105.xml.

Nekrutkin:2020:BDP

[Nek20] Vladimir Nekrutkin. Binary decompositions of probability densities and random-bit simulation. Monte Carlo Methods and Applications, 26(2):163–169, April 17, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2020-2063/html.

Newton:2001:ANF

[New01] Nigel J. Newton. Approximations for nonlinear filters based on quantisation. Monte Carlo Methods and Applications, 7(3–4):311– 320, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-3-4/mcma.2001.7.3-4.311/mcma.2001.

7.3-4.311.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Ngnepieba:2006:OCS

[NHD06] Pierre Ngnepieba, M. Y. Hussaini, and Laurent Debreu. Optimal control and stochastic parameter estimation. Monte Carlo Methods and Applications, 12(5–6):461–476, ???? 2006. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-5/156939606779329062/156939606779329062.xml.

Ninomiya:2003:PSM

[Nin03] Syoiti Ninomiya. A partial sampling method applied to the Kusuoka approximation. Monte Carlo Methods and Applications, 9(1):27–38, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2003.9.issue-1/156939603322587443/156939603322587443.
xml.

Nadarajah:2006:DSR

[NK06] Saralees Nadarajah and Samuel Kotz. Determination of software reliability based on multivariate exponential, Lomax and Weibull models. Monte Carlo Methods and Applications, 12(5–6):447–459, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2006.12.issue-5/156939606779329035/156939606779329035.
xml.

Nguyen:2016:ARM

[Nk16] Nguyet Nguyen and Giray kten. The acceptance–rejection method for low-discrepancy sequences. Monte Carlo Methods and Applications, 22(2):133–??, June 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-2/
mcma-2016-0104/mcma-2016-0104.xml.

Nilsson:2004:FBM

[NMH04] Hans-Erik Nilsson, Antonio Martinez, and Mats Hjelm. Full band Monte Carlo simulation — beyond the semiclassical approach. Monte Carlo Methods and Applications, 10(3–4):481– 490, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.481/
mcma.2004.10.3-4.481.xml.

Ngo:2009:CLT

[NO09a] Hoang-Long Ngo and Shigeyoshi Ogawa. A central limit theorem for the functional estimation of the spot volatility. Monte Carlo Methods and Applications, 15(4):353–380, December 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-4/mcma.2009.019/mcma.2009.019.xml.

Nystrom:2009:MCA

[NÖ09b] Kaj Nyström and Thomas Önskog. On Monte Carlo algorithms applied to Dirichlet problems for parabolic operators in the setting of time-dependent domains. Monte Carlo Methods and Applications, 15(1):11–47, May 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2009.15.issue-1/mcma.2009.002/mcma.2009.002.
xml.

Nekrutkin:2004:TVS

[NP04a] V. Nekrutkin and P. Potapov. Two variants of a stochastic Euler method for homogeneous balance differential equations. Monte Carlo Methods and Applications, 10(3–4):469–479, December 2004.

CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.469/mcma.2004.10.3-4.469.
xml.

Novak:2004:CTA

[NP04b] Erich Novak and Harald Pfeiffer. Coin tossing algorithms for integral equations and tractability. Monte Carlo Methods and Applications, 10(3–4):491–498, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.491/mcma.2004.10.3-4.491.xml.

Nieto:2006:MCS

[NPM^{+}06] F. J. Rodríguez Nieto, M. A. Pasquale, M. E. Martins, F. A. Bareilles, and A. J. Arvia. Monte Carlo simulations of solid 2D phase growth on 1D solid substrates with square-wave surface profiles. Influence of hole design and depositing particle surface diffusion. Monte Carlo Methods and Applications, 12(3–4):271–289, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2006.12.issue-3/156939606778705209/156939606778705209.
xml.

Nekrutkin:2002:PCK

[NR02] V. Nekrutkin and K. Romkin. Propagation of chaos for Kac particle systems. Monte Carlo Methods and Applications, 8(3):299–315, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2002.8.issue-3/mcma.2002.8.3.299/mcma.2002.8.3.299.xml.

Nekrutkin:2007:AAO

[NS07] V. Nekrutkin and M. Samakhova. Admissible and asymptotically optimal linear congruential generators. Monte Carlo Methods and Applications, 13(3):227–244, August 2007. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-3/mcma.2007.012/mcma.2007.012.xml.

Nekrutkin:2009:STS

[NS09] V. Nekrutkin and R. Sabitov. Spectral test and spectral distance for multiplicative generators with moduli 2^{p}. Monte Carlo Methods and Applications, 15(1):1–10, May 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-1/mcma.2009.001/mcma.2009.001.xml.

Nekrutkin:1997:AEE

[NT97] V. Nekrutkin and N. Tur. Asymptotic expansions and estimators with small bias for Nanbu processes. Monte Carlo Methods and Applications, 3(1):1–35, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
1/mcma.1997.3.1.1/mcma.1997.3.1.1.xml.

Nakagawa:2021:MCS

[NT21] Takuya Nakagawa and Akihiro Tanaka. On a Monte Carlo scheme for some linear stochastic partial differential equations. Monte Carlo Methods and Applications, 27(2):169–193, April 24, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2088/html.

Nedjalkov:2007:MIB

[NVDA07] M. Nedjalkov, D. Vasileska, I. Dimov, and G. Arsov. Mixed initialboundary value problem in particle modeling of microelectronic devices. Monte Carlo Methods and Applications, 13(4):299–331, November 20, 2007. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2007.13.issue-4/mcma.2007.017/mcma.2007.017.xml.

Nguyen:2018:QMC

[NXÖ18] Nguyet Nguyen, Linlin Xu, and Giray Ökten. A quasi-Monte Carlo implementation of the ziggurat method. Monte Carlo Methods and Applications, 24(2):93–99, June 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-2/
mcma-2018-0008/mcma-2018-0008.xml.

Naito:2019:SOD

[NY19a] Riu Naito and Toshihiro Yamada. A second-order discretization for forward-backward SDEs using local approximations with Malliavin calculus. Monte Carlo Methods and Applications, 25(4):341–??, December 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2019.25.issue-4/mcma-2019-2053/mcma-2019-2053.xml.

Naito:2019:TOW

[NY19b] Riu Naito and Toshihiro Yamada. A third-order weak approximation of multidimensional Itô stochastic differential equations. Monte Carlo Methods and Applications, 25(2):97–??, June 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-2/mcma-2019-2036/mcma-2019-2036.xml.

Neuenkirch:2009:AED

[NZ09] Andreas Neuenkirch and Henryk Zähle. Asymptotic error distribution of the Euler method for SDEs with non-Lipschitz coefficients. Monte Carlo Methods and Applications, 15(4):333–351, December 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-4/mcma.2009.018/mcma.2009.018.xml.

Okten:2009:CPP

[ÖG09] Giray Ökten and Michael Gnewuch. Correction of a proof in “A probabilistic result on the discrepancy of a hybrid-Monte Carlo sequence and applications” [Cno. 1434421]. Monte Carlo Methods and Applications, 15(2):169–172, August 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-2/mcma.2009.010/mcma.2009.010.xml. See [Ökt96].

Ogawa:1996:RRP

[Oga96] Shigeyoshi Ogawa. On a robustness of the random particle method. Monte Carlo Methods and Applications, 2(3):175–189, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1996.2.
issue-3/mcma.1996.2.3.175/mcma.1996.2.3.175.xml. See erra-

tum [Oga97].

Ogawa:1997:EAR

[Oga97] Shigeyoshi Ogawa. Erratum to the article: “On a robustness of the random particle method” [Monte Carlo Methods Appl. 2 (1996), no. 3, 175–189; MR1414863 (97j:65008)]. Monte Carlo Methods and Applications, 3(1):83–??, ???? 1997. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
1/mcma.1997.3.1.83/mcma.1997.3.1.83.xml. See [Oga96].

Ogawa:2001:CSC

[Oga01] Shigeyoshi Ogawa. On a class of SPDEs called Brownian particle equation — model for nonlinear diffusions. Monte Carlo Methods and Applications, 7(3–4):321–328, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.321/mcma.2001.7.3-4.321.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Ogawa:2008:RTS

[Oga08] Shigeyoshi Ogawa. Real-time scheme for the volatility estimation in the presence of microstructure noise. Monte Carlo Methods and Applications, 14(4):331–342, November 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-4/mcma.2008.015/mcma.2008.015.xml.

Okten:1996:PRD

[Ökt96] Giray Ökten. A probabilistic result on the discrepancy of a hybrid-Monte Carlo sequence and applications. Monte Carlo Methods and Applications, 2(4):255–270, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
4/mcma.1996.2.4.255/mcma.1996.2.4.255.xml. See correction

[ÖG09].

Olla:2001:DBI

[Oll01] Stefano Olla. Diffusive behavior of interacting particle systems. Monte Carlo Methods and Applications, 7(3–4):329–338, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2001.7.issue-3-4/mcma.2001.7.3-4.329/mcma.2001.7.

3-4.329.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Ongaro:1999:MCS

[ONZ99] C. Ongaro, U. Nastasi, and A. Zanini. Monte Carlo simulation of the photo-neutron production in the high-Z components of radiotherapy linear accelerators. Monte Carlo Methods and Applications, 5(1):69–79, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1999.5.issue-1/mcma.1999.5.1.
69/mcma.1999.5.1.69.xml.

Ogihara:2003:DSA

[OO03] Shuhei Ogihara and Shigeyoshi Ogawa. On a discrete stochastic approximation and its application to data analysis. Monte Carlo Methods and Applications, 9(1):39–50, January 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
1/156939603322587452/156939603322587452.xml.

Osada:2001:TPI

[Osa01] Hirofumi Osada. Tagged particles of interacting Brownian motions with skew symmetric drifts. Monte Carlo Methods and Applications, 7(3–4):339–348, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2001.7.issue-3-4/mcma.2001.7.
3-4.339/mcma.2001.7.3-4.339.xml. Monte Carlo and probabilis-

tic methods for partial differential equations, Part II (Monte Carlo, 2000).

Ogawa:2007:RTS

[OW07] Shigeyoshi Ogawa and Koji Wakayama. On a real-time scheme for the estimation of volatility. Monte Carlo Methods and Applications, 13(2):99–116, ???? 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2007.13.issue-2/mcma.2007.006/mcma.2007.006.
xml.

Owen:2006:WHQ

[Owe06] Art B. Owen. On the Warnock–Halton quasi-standard error. Monte Carlo Methods and Applications, 12(1):47–54, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-1/156939606776886652/156939606776886652.xml.

Okano:2019:CVM

[OY19] Yusuke Okano and Toshihiro Yamada. A control variate method for weak approximation of SDEs via discretization of numerical error of asymptotic expansion. Monte Carlo Methods and Applications, 25(3):239–??, September 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-3/
mcma-2019-2044/mcma-2019-2044.xml.

Pages:2007:MSR

[Pag07] Gilles Pagès. Multi-step Richardson–Romberg extrapolation: Remarks on variance control and complexity. Monte Carlo Methods and Applications, 13(1):37–70, ???? 2007. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-1/mcma.2007.003/mcma.2007.003.xml.

Pantsulaia:2015:IDM

[Pan15] Gogi R. Pantsulaia. Infinite-dimensional Monte-Carlo integration. Monte Carlo Methods and Applications, 21(4):283–??, December 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2015.
21.issue-4/mcma-2015-0108/mcma-2015-0108.xml.

Papadopoulos:1998:NTM

[Pap98] C. Papadopoulos. A new technique for MTTF estimation in highly reliable Markovian systems. Monte Carlo Methods and Applications, 4(2):95–111, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1998.4.issue-2/mcma.1998.4.2.
95/mcma.1998.4.2.95.xml.

Papancheva:2004:OKC

[Pap04] R. Y. Papancheva. Optimal Korobov coefficients for good lattice points in quasi Monte Carlo algorithms. Monte Carlo Methods and Applications, 10(3–4):499–509, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.499/mcma.2004.10.3-4.499.xml.

Pillards:2004:TVT

[PC04] Tim Pillards and Ronald Cools. A theoretical view on transforming low-discrepancy sequences from a cube to a simplex. Monte Carlo Methods and Applications, 10(3–4):511–529, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.511/mcma.2004.10.3-4.511.
xml.

Pelsser:2019:MCM

[PG19] Antoon Pelsser and Kossi Gnameho. A Monte Carlo method for backward stochastic differential equations with Hermite martingales. Monte Carlo Methods and Applications, 25(1):37–60, March 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-1/mcma-2019-2028/mcma-2019-2028.xml.

Pohl:1998:PML

[PGB98] Thomas Pohl, Wilfried Grecksch, and Holger Blaar. A parallel modified Lagrangian method for an optimal control problem of a linear distributed stochastic system. Monte Carlo Methods and Applications, 4(4):319–340, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
4/mcma.1998.4.4.319/mcma.1998.4.4.319.xml.

Pletnev:2009:CMS

[PGS09] Leonid Pletnev, Maxim Gvozdev, and Kiryl Samartsau. Computer modeling of stationary particles transport in open cylindrical nanosystems by Monte Carlo method. Monte Carlo Methods and Applications, 15(1):49–62, May 2009. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2009.15.issue-1/mcma.2009.003/
mcma.2009.003.xml.

Portolan:2004:WSS

[PIR04] Stefano Portolan, Rita C. Iotti, and Fausto Rossi. Weighted simulation of steady-state transport within the standard Monte Carlo paradigm. Monte Carlo Methods and Applications, 10(3–4): 531–539, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.531/
mcma.2004.10.3-4.531.xml.

Piterbarg:2006:PEM

[Pit06] Leonid I. Piterbarg. Parameter estimation in multi particle Lagrangian stochastic models. Monte Carlo Methods and Applications, 12(5–6):477–493, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2006.12.issue-5/156939606779329044/
156939606779329044.xml.

Pletnev:2000:MCS

[Ple00] Leonid Pletnev. Monte Carlo simulation of evaporation process into the vacuum. Monte Carlo Methods and Applications, 6(3):191–203, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2000.6.issue-3/mcma.2000.6.3.191/mcma.2000.6.3.191.xml.

Prigarin:2010:SBR

[PMW10] Sergei M. Prigarin, Andreas Martin, and Gerhard Winkler. Simulation of binary random fields with Gaussian numerical models. Monte Carlo Methods and Applications, 16(2):129–142, July 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-2/mcma.2010.004/mcma.2010.004.xml.

Protasov:2004:DPM

[PO04] A. V. Protasov and V. A. Ogorodnikov. Dynamic probabilistic method of numerical modeling of multidimensional hydrometeorological fields. Monte Carlo Methods and Applications, 10(3–4): 541–549, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.541/
mcma.2004.10.3-4.541.xml.

Polala:2020:IBE

[PÖ20] Arun Kumar Polala and Giray Ökten. Implementing de-biased estimators using mixed sequences. Monte Carlo Methods and Applications, 26(4):293–301, October 2, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2020-2075/html.

Potzelberger:2012:IMC

[Pöt12] Klaus Pötzelberger. Improving the Monte Carlo estimation of boundary crossing probabilities by control variables. Monte Carlo Methods and Applications, 18(4):353–??, December 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-4/mcma-2012-0013/mcma-2012-0013.xml.

Pages:2003:OQQ

[PP03] Gilles Pagès and Jacques Printems. Optimal quadratic quantization for numerics: the Gaussian case. Monte Carlo Methods and Applications, 9(2):135–165, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
2/156939603322663321/156939603322663321.xml.

Puig:2004:WNS

[PP04] Bénédicte Puig and Fabrice Poirion. White noise and simulation of ordinary Gaussian processes. Monte Carlo Methods and Applications, 10(1):69–89, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2004.10.issue-1/156939604323091216/156939604323091216.
xml.

Pages:2005:FQN

[PP05] Gilles Pagès and Jacques Printems. Functional quantization for numerics with an application to option pricing. Monte Carlo Methods and Applications, 11(4):407–446, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-4/156939605777438578/156939605777438578.xml.

Pham:2019:BAM

[PP19] Hoa Pham and Huong T. T. Pham. A Bayesian approach for multi-stage models with linear time-dependent hazard rate. Monte Carlo Methods and Applications, 25(4):307–??, December 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-4/mcma-2019-2051/mcma-2019-2051.xml.

Pham:2021:EPM

[PP21] Huong T. T. Pham and Hoa Pham. On the existence of posterior mean for Bayesian logistic regression. Monte Carlo Methods and Applications, 27(3):277–288, May 18, 2021. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2089/html.

Pham:2020:BIP

[PPN20] Huong Thi Thu Pham, Hoa Pham, and Darfiana Nur. A Bayesian inference for the penalized spline joint models of longitudinal and time-to-event data: A prior sensitivity analysis. Monte Carlo Methods and Applications, 26(1):49–??, March 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2020.
26.issue-1/mcma-2020-2058/mcma-2020-2058.xml.

Pages:2019:RCI

[PR19] Gilles Pagès and Clément Rey. Recursive computation of the invariant distributions of Feller processes: Revisited examples and new applications. Monte Carlo Methods and Applications, 25(1):1–36, March 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2019.25.issue-1/mcma-2018-2027/mcma-2018-2027.xml.

Printems:2001:DTP

[Pri01] Jacques Printems. On the discretization in time of parabolic stochastic partial differential equations. Monte Carlo Methods and Applications, 7(3–4):359–368, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.359/mcma.2001.7.3-4.359.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Pham:2005:AQF

[PRS05] Huyên Pham, Wolfgang Runggaldier, and Afef Sellami. Approximation by quantization of the filter process and applications to optimal stopping problems under partial observation. Monte Carlo Methods and Applications, 11(1):57–81, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-1/1569396054027283/1569396054027283.xml.

Plotnikov:1998:EEO

[PS98] M. Yu. Plotnikov and E. V. Shkarupa. Error estimation and optimization in C-space of Monte Carlo iterative solution of nonlinear integral equations. Monte Carlo Methods and Applications, 4(1):53–71, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1998.4.issue-1/mcma.1998.4.1.
53/mcma.1998.4.1.53.xml.

Plotnikov:2005:DSA

[PS05] Mikhail Plotnikov and Elena Shkarupa. The discrete-stochastic approaches to solving the linearized Boltzmann equation. Monte Carlo Methods and Applications, 11(4):447–462, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-4/156939605777438532/156939605777438532.xml.

Pausinger:2010:GPO

[PS10] Florian Pausinger and Wolfgang Ch. Schmid. A good permutation for one-dimensional diaphony. Monte Carlo Methods and Applications, 16(3–4):307–322, December 2010. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.015/mcma.2010.015.xml.

Pareschi:2001:RMC

[PW01] Lorenzo Pareschi and Bernt Wennberg. A recursive Monte Carlo method for the Boltzmann equation in the Maxwellian case. Monte Carlo Methods and Applications, 7(3–4):349–357, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.349/mcma.2001.7.3-4.349.xml. Monte

Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Perel:1999:MCC

[PWY99] R. L. Perel, J. J. Wagschal, and Y. Yeivin. Monte Carlo calculation of deep penetration benchmark sensitivities. Monte Carlo Methods and Applications, 5(2):169–187, ???? 1999. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
2/mcma.1999.5.2.169/mcma.1999.5.2.169.xml.

Rajput:2019:GAD

[Raj19] Nikhil Kumar Rajput. Gillespie algorithm and diffusion approximation based on Monte Carlo simulation for innovation diffusion: a comparative study. Monte Carlo Methods and Applications, 25(3):209–??, September 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-3/
mcma-2019-2040/mcma-2019-2040.xml.

Refas:2021:SSC

[RBB21] Salah Eddine Chouaib Refas, Abdelkader Bouazza, and Youcef Belhadji. 3D sputtering simulations of the CZTS, Si and CIGS thin films using Monte-Carlo method. Monte Carlo Methods and Applications, 27(4):373–382, October 21, 2021. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2094/html.

Reichl:2020:EML

[Rei20] Johannes Reichl. Estimating marginal likelihoods from the posterior draws through a geometric identity. Monte Carlo Methods and Applications, 26(3):205–221, August 5, 2020. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2020-2068/html.

Rey:2017:CTV

[Rey17] Clément Rey. Convergence in total variation distance of a third order scheme for one-dimensional diffusion processes. Monte Carlo Methods and Applications, 23(1):1–??, March 2017. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-1/mcma-2016-0120/mcma-2016-0120.xml.

Rief:1999:TZV

[Rie99] H. Rief. Touching on a zero-variance scheme in solving linear equations by random walk processes. Monte Carlo Methods and Applications, 5(2):135–148, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1999.5.issue-
2/mcma.1999.5.2.135/mcma.1999.5.2.135.xml.

Ruzayqat:2020:UES

[RJ20] Hamza M. Ruzayqat and Ajay Jasra. Unbiased estimation of the solution to Zakai’s equation. Monte Carlo Methods and Applications, 26(2):113–129, April 15, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-2061/html.

Rostamy:2013:WCE

[RJG13] Davoud Rostamy, Mohammad Jabbari, and Mahshid Gadirian. Worst case error for integro-differential equations by a lattice– Nyström method. Monte Carlo Methods and Applications, 19(4): 281–??, December 2013. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2013.19.issue-4/mcma-2013-0013/mcma-2013-0013.
xml.

Rasulov:2004:QSB

[RKM04] Abdujabor Rasulov, Aneta Karaivanova, and Michael Mascagni. Quasirandom sequences in branching random walks. Monte Carlo Methods and Applications, 10(3–4):551–558, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.551/mcma.2004.10.3-4.551.
xml.

Rozora:2018:MLS

[RL18] Iryna Rozora and Mariia Lyzhechko. On the modeling of linear system input stochastic processes with given accuracy and reliability. Monte Carlo Methods and Applications, 24(2):129–137, June 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-2/mcma-2018-0011/mcma-2018-0011.xml.

Rogasinsky:1996:PCP

[Rog96] S. V. Rogasinsky. On the pair correlations of particle evolution in the direct statistical simulation. Monte Carlo Methods and Applications, 2(1):25–40, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1996.2.issue-1/mcma.1996.2.1.
25/mcma.1996.2.1.25.xml.

Rogasinsky:1999:SSB

[Rog99] S. V. Rogasinsky. Solution of stationary boundary value problems for the Boltzmann equation by the Monte Carlo method. Monte Carlo Methods and Applications, 5(3):263–280, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1999.5.
issue-3/mcma.1999.5.3.263/mcma.1999.5.3.263.xml.

Roth:2007:WAS

[Rot07] Christian Roth. Weak approximations of solutions of a first order hyperbolic stochastic partial differential equation. Monte Carlo Methods and Applications, 13(2):117–133, ???? 2007. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-2/mcma.2007.007/mcma.2007.007.xml.

Rowe:2000:FSP

[Row00] Daniel B. Rowe. Factorization of separable and patterned covariance matrices for Gibbs sampling. Monte Carlo Methods and Applications, 6(3):205–210, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
3/mcma.2000.6.3.205/mcma.2000.6.3.205.xml.

Rowe:2002:JDM

[Row02] Daniel B. Rowe. Jointly distributed mean and mixing coefficients for Bayesian source separation using MCMC and ICM. Monte Carlo Methods and Applications, 8(4):395–403, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
4/mcma.2002.8.4.395/mcma.2002.8.4.395.xml.

Rowe:2003:SFN

[Row03] Daniel B. Rowe. Significant FMRI neurologic synchrony using Monte Carlo methods. Monte Carlo Methods and Applications, 9(4):367–385, ???? 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2003.9.issue-4/156939603322601978/156939603322601978.
xml.

Rabiei:2019:SMM

[RS19] Nima Rabiei and Elias G. Saleeby. On the sample-mean method for computing hyper-volumes. Monte Carlo Methods and Applications, 25(2):163–??, June 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-2/
mcma-2019-2034/mcma-2019-2034.xml.

Rabiei:2021:IVC

[RS21] Nima Rabiei and Elias G. Saleeby. On intersection volumes of confidence hyper-ellipsoids and two geometric Monte Carlo methods. Monte Carlo Methods and Applications, 27(2):153–167, April 30, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2021-2087/html.

Radovic:1996:QMC

[RST96] Igor Radović, Ilya M. Sobol′, and Robert F. Tichy. Quasi-Monte Carlo methods for numerical integration: comparison of different low discrepancy sequences. Monte Carlo Methods and Applications, 2(1):1–14, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1996.2.issue-1/mcma.1996.2.1.1/mcma.1996.2.
1.1.xml.

Rudolf:2010:EBC

[Rud10] Daniel Rudolf. Error bounds for computing the expectation by Markov chain Monte Carlo. Monte Carlo Methods and Applications, 16(3–4):323–342, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2010.16.issue-3-4/mcma.
2010.012/mcma.2010.012.xml.

Rouzankin:1999:CAR

[RV99] P. S. Rouzankin and A. V. Voytishek. On the cost of algorithms for random selection. Monte Carlo Methods and Applications, 5(1):39–54, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1999.5.issue-1/mcma.1999.5.1.
39/mcma.1999.5.1.39.xml.

Sabelfeld:1996:MCS

[SA96] K. K. Sabelfeld and T. A. Averina. Monte Carlo simulation of particle’s dispersion in convective boundary layer of the atmosphere. Monte Carlo Methods and Applications, 2(2):159–169, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1996.2.
issue-2/mcma.1996.2.2.159/mcma.1996.2.2.159.xml.

Sabelfeld:2004:F

[Sab04a] Karl Sabelfeld. Foreword. Monte Carlo Methods and Applications, 10(3–4):i–??, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.i/mcma.
2004.10.3-4.i.xml.

Sabelfeld:2004:FSP

[Sab04b] Karl Sabelfeld. Foreword: [selection of papers presented at the IV IMACS Seminar on Monte Carlo methods]. Monte Carlo Methods and Applications, 10(3–4):181–182, 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). Held in Berlin, September 15–19, 2003.

Sabelfeld:2008:ERB

[Sab08] Karl Sabelfeld. Expansion of random boundary excitations for elliptic PDEs. Monte Carlo Methods and Applications, 13(5–6): 405–453, January 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2008.13.issue-5-6/mcma.2007.022/mcma.2007.
022.xml.

Sabelfeld:2016:RWSa

[Sab16a] Karl K. Sabelfeld. Random walk on semi-cylinders for diffusion problems with mixed Dirichlet–Robin boundary conditions. Monte Carlo Methods and Applications, 22(2):117–??, June 2016. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-2/mcma-2016-0108/mcma-2016-0108.xml.

Sabelfeld:2016:RWSb

[Sab16b] Karl K. Sabelfeld. Random walk on spheres method for solving drift–diffusion problems. Monte Carlo Methods and Applications, 22(4):265–??, December 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-4/
mcma-2016-0118/mcma-2016-0118.xml.

Sabelfeld:2016:SSP

[Sab16c] Karl K. Sabelfeld. Splitting and survival probabilities in stochastic random walk methods and applications. Monte Carlo Methods and Applications, 22(1):55–??, March 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-1/
mcma-2016-0103/mcma-2016-0103.xml.

Sabelfeld:2016:VMC

[Sab16d] Karl K. Sabelfeld. Vector Monte Carlo stochastic matrixbased algorithms for large linear systems. Monte Carlo Methods and Applications, 22(3):259–??, September 2016. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-3/mcma-2016-0112/mcma-2016-0112.xml.

Sabelfeld:2017:RWS

[Sab17] Karl K. Sabelfeld. Random walk on spheres algorithm for solving transient drift-diffusion-reaction problems. Monte Carlo Methods and Applications, 23(3):189–??, September 2017. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-3/mcma-2017-0113/mcma-2017-0113.xml.

Sabelfeld:2019:GRW

[Sab19a] Karl K. Sabelfeld. A global random walk on spheres algorithm for transient heat equation and some extensions. Monte Carlo Methods and Applications, 25(1):85–96, March 2019. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-1/mcma-2019-2032/mcma-2019-2032.xml.

Sabelfeld:2019:RWR

[Sab19b] Karl K. Sabelfeld. Random walk on rectangles and parallelepipeds algorithm for solving transient anisotropic drift–diffusion–reaction problems. Monte Carlo Methods and Applications, 25(2):131–??, June 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/view/j/mcma.
2019.25.issue-2/mcma-2019-2039/mcma-2019-2039.xml.

Sagna:2011:PBO

[Sag11] Abass Sagna. Pricing of barrier options by marginal functional quantization. Monte Carlo Methods and Applications, 17(4):371– 398, December 2011. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2011.17.issue-4/mcma.2011.015/mcma.2011.015.xml.

Sak:2010:INI

[Sak10] Halis Sak. Increasing the number of inner replications of multifactor portfolio credit risk simulation in the t-copula model. Monte Carlo Methods and Applications, 16(3–4):361–377, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.013/mcma.2010.013.xml.

Sharifzadeh:2015:MBM

[SAKG15] Mohsen Sharifzadeh, Hosein Afarideh, Hosein Khalafi, and Reza Gholipour. A Matlab-based Monte Carlo algorithm for transport of gamma-rays in matter. Monte Carlo Methods and Applications, 21(1):77–??, March 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2015.21.issue-1/mcma-2014-0011/mcma-2014-0011.
xml.

Sbert:2004:RPR

[SBH04] Mateu Sbert, Philippe Bekaert, and John Halton. Reusing paths in radiosity and global illumination. Monte Carlo Methods and Applications, 10(3–4):575–585, December 2004. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.575/mcma.2004.10.3-4.575.xml.

Sharma:1996:PMQ

[SC96] G. C. Sharma and M. S. Chauhan. Preventive maintenance of an M^{X}/G/1 queue-like production system. Monte Carlo Methods and Applications, 2(2):129–137, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
2/mcma.1996.2.2.129/mcma.1996.2.2.129.xml.

Sahoo:1996:MCC

[SD96] L. N. Sahoo and M. Dalabehera. A Monte Carlo comparison of six almost unbiased ratio estimators. Monte Carlo Methods and Applications, 2(3):237–249, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
3/mcma.1996.2.3.237/mcma.1996.2.3.237.xml.

Sabelfeld:2018:HKT

[SE18] Karl K. Sabelfeld and Georgy Eremeev. A hybrid kineticthermodynamic Monte Carlo model for simulation of homogeneous burst nucleation. Monte Carlo Methods and Applications, 24(3):193–202, September 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-3/
mcma-2018-0017/mcma-2018-0017.xml.

Seibold:2004:OPM

[Sei04] Benjamin Seibold. Optimal prediction in molecular dynamics. Monte Carlo Methods and Applications, 10(1):25–50, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-1/156939604323091199/156939604323091199.xml.

Sentis:2001:MCM

[Sen01] R. Sentis. Monte Carlo methods in neutron and photon transport problems. Monte Carlo Methods and Applications, 7(3–4): 383–395, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-3-4/mcma.2001.7.3-4.383/mcma.2001.

7.3-4.383.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Stollinger:2008:PMS

[SH08] Michael Stöllinger and Stefan Heinz. PDF modeling and simulation of premixed turbulent combustion. Monte Carlo Methods and Applications, 14(4):343–377, November 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
14.issue-4/mcma.2008.016/mcma.2008.016.xml.

Slaoui:2022:RRE

[SH22] Yousri Slaoui and Salima Helali. Recursive regression estimation based on the two-time-scale stochastic approximation method and Bernstein polynomials. Monte Carlo Methods and Applications, 28 (1):45–59, February 15, 2022. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2022-2104/html.

Shah:2010:GAA

[Sha10] Manan Shah. A genetic algorithm approach to estimate lower bounds of the star discrepancy. Monte Carlo Methods and Applications, 16(3–4):379–398, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2010.16.issue-3-4/mcma.
2010.014/mcma.2010.014.xml.

Shvets:2003:ABM

[Shv03] V. V. Shvets. On asymptotic behaviour of modelling time in the importance sampling method. Monte Carlo Methods and Applications, 9(1):77–85, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
1/156939603322587489/156939603322587489.xml.

Simonov:1995:BVP

[Sim95] N. A. Simonov. Boundary value problem and stochastic algorithm for two-dimensional Navier–Stokes equations. Monte Carlo Methods and Applications, 1(1):59–70, ???? 1995. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
1/mcma.1995.1.1.59/mcma.1995.1.1.59.xml.

Simonov:2018:RWA

[Sim18] Nikolai A. Simonov. Random walk algorithms for elliptic equations and boundary singularities. Monte Carlo Methods and Applications, 24(4):323–327, December 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-4/
mcma-2018-2026/mcma-2018-2026.xml.

Singh:2014:CMC

[Sin14] Vipul Kumar Singh. Competency of Monte Carlo and Black– Scholes in pricing Nifty index options: A vis-à-vis study. Monte Carlo Methods and Applications, 20(1):61–??, March 2014. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-1/mcma-2013-0017/mcma-2013-0017.xml.

Sabelfeld:1997:MCS

[SK97a] K. K. Sabelfeld and A. A. Kolodko. Monte Carlo simulation of the coagulation processes governed by Smoluchowski equation with random coefficients. Monte Carlo Methods and Applications, 3 (4):275–311, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1997.3.issue-4/mcma.1997.3.4.275/mcma.1997.
3.4.275.xml.

Sabelfeld:1997:SLM

[SK97b] K. K. Sabelfeld and O. Kurbanmuradov. Stochastic Lagrangian models for two-particle motion in turbulent flows. Monte Carlo Methods and Applications, 3(1):53–72, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1997.3.issue-
1/mcma.1997.3.1.53/mcma.1997.3.1.53.xml.

Sabelfeld:1998:OPS

[SK98] K. Sabelfeld and O. Kurbanmuradov. One-particle stochastic Lagrangian model for turbulent dispersion in horizontally homogeneous turbulence. Monte Carlo Methods and Applications, 4 (2):127–140, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1998.4.issue-2/mcma.1998.4.2.127/mcma.1998.
4.2.127.xml.

Sabelfeld:2000:CAP

[SK00] K. Sabelfeld and O. Kurbanmuradov. Coagulation of aerosol particles in intermittent turbulent flows. Monte Carlo Methods and Applications, 6(3):211–253, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
3/mcma.2000.6.3.211/mcma.2000.6.3.211.xml.

Sabelfeld:2003:SEM

[SK03] Karl Sabelfeld and Dmitry Kolyukhin. Stochastic Eulerian model for the flow simulation in porous media. Monte Carlo Methods and Applications, 9(3):271–290, ???? 2003. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2003.9.issue-
3/156939603322729021/156939603322729021.xml.

Sobol:2005:GSA

[SK05] I. M. Sobol′ and S. S. Kucherenko. On global sensitivity analysis of quasi-Monte Carlo algorithms. Monte Carlo Methods and Applications, 11(1):83–92, ???? 2005. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-1/1569396054027274/1569396054027274.xml.

Stallinga:2015:CS

[SK15] Peter Stallinga and Igor Khmelinskii. Consensus in science. Monte Carlo Methods and Applications, 21(1):69–??, March 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-1/mcma-2014-0008/mcma-2014-0008.xml.

Sabelfeld:2018:PDL

[SK18] Karl K. Sabelfeld and Anastasiya Kireeva. Probability distribution of the life time of a drift–diffusion-reaction process inside a sphere with applications to transient cathodoluminescence imaging. Monte Carlo Methods and Applications, 24(2):79–92, June 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-2/mcma-2018-0007/mcma-2018-0007.xml.

Sabelfeld:2009:SSP

[SKL09] K. Sabelfeld, O. Kurbanmuradov, and A. Levykin. Stochastic simulation of particle transport by a random Darcy flow through a porous cylinder. Monte Carlo Methods and Applications, 15(1): 63–90, May 2009. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2009.15.issue-1/mcma.2009.004/mcma.2009.004.xml.

Sabelfeld:2010:SIP

[SL10] Karl Sabelfeld and Nadja Loshchina. Stochastic iterative projection methods for large linear systems. Monte Carlo Methods and Applications, 16(3–4):343–359, December 2010. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.020/mcma.2010.020.xml.

Sabelfeld:2014:SMI

[SL14] Karl K. Sabelfeld and Alexander I. Levykin. A spectral method for isotropic diffusion equation with random concentration fluctuations of incoming flux of particles through circular-shaped boundaries. Monte Carlo Methods and Applications, 20(3):173–??, September 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-3/mcma-2014-0001/mcma-2014-0001.xml.

Sabelfeld:2015:SSF

[SLK15] Karl K. Sabelfeld, Alexander I. Levykin, and Anastasiya E. Kireeva. Stochastic simulation of fluctuation-induced reactiondiffusion kinetics governed by Smoluchowski equations. Monte Carlo Methods and Applications, 21(1):33–??, March 2015. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2015.
21.issue-1/mcma-2014-0012/mcma-2014-0012.xml.

Sabelfeld:2007:FSS

[SLP07] K. Sabelfeld, A. Levykin, and T. Privalova. A fast stratified sampling simulation of coagulation processes. Monte Carlo Methods and Applications, 13(1):71–88, ???? 2007. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-1/mcma.2007.004/mcma.2007.004.xml.

Sobol:2003:MCR

[SM03] I. M. Sobol and E. E. Myshetskaya. Modelling correlated random variables. Monte Carlo Methods and Applications, 9(1):67–76, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2003.9.issue-1/156939603322587470/156939603322587470.xml.

Simonov:2004:RWA

[SM04] N. A. Simonov and M. Mascagni. Random walk algorithms for estimating effective properties of digitized porous media. Monte Carlo Methods and Applications, 10(3–4):599–608, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.599/mcma.2004.10.3-4.599.
xml.

Sobol:2008:MCE

[SM08] I. M. Sobol′ and E. E. Myshetskaya. Monte Carlo estimators for small sensitivity indices. Monte Carlo Methods and Applications, 13(5–6):455–465, January 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2008.13.issue-5-6/mcma.
2007.023/mcma.2007.023.xml.

Sabelfeld:2009:SRA

[SM09] K. Sabelfeld and N. Mozartova. Sparsified randomization algorithms for large systems of linear equations and a new version of the random walk on boundary method. Monte Carlo Methods and Applications, 15(3):257–284, November 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-3/mcma.2009.015/mcma.2009.015.xml.

Sabelfeld:2012:SBC

[SM12] Karl Sabelfeld and Nadezhda Mozartova. Stochastic boundary collocation and spectral methods for solving PDEs. Monte Carlo Methods and Applications, 18(3):217–??, September 2012. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2012.
18.issue-3/mcma-2012-0008/mcma-2012-0008.xml.

Smidts:1998:PSE

[Smi98] O. F. Smidts. Point and surface estimations by a non-analog Monte Carlo simulation for the transport of radionuclide chains in porous media. Monte Carlo Methods and Applications, 4(4):289–318, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1998.4.
issue-4/mcma.1998.4.4.289/mcma.1998.4.4.289.xml.

Schretter:2013:DIM

[SN13] Colas Schretter and Harald Niederreiter. A direct inversion method for non-uniform quasi-random point sequences. Monte Carlo Methods and Applications, 19(1):1–??, March 2013. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-1/mcma-2012-0014/mcma-2012-0014.xml.

Sellier:2014:BSW

[SNDS14] Jean Michel Sellier, Mihail Nedjalkov, Ivan Dimov, and Siegfried Selberherr. A benchmark study of the Wigner Monte Carlo method. Monte Carlo Methods and Applications, 20(1):43–??, March 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-1/mcma-2013-0018/mcma-2013-0018.xml.

Sabelfeld:2020:MCT

[SP20] Karl K. Sabelfeld and Nikita Popov. Monte Carlo tracking driftdiffusion trajectories algorithm for solving narrow escape problems. Monte Carlo Methods and Applications, 26(3):177–191, August 6, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2073/html.

Spade:2021:EDM

[Spa21] David A. Spade. Estimating drift and minorization coefficients for Gibbs sampling algorithms. Monte Carlo Methods and Applications, 27(3):195–209, August 8, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2021-2093/html.

Sabelfeld:1996:SAS

[SRKL96] K. K. Sabelfeld, S. V. Rogasinsky, A. A. Kolodko, and A. I. Levykin. Stochastic algorithms for solving Smolouchovsky coagulation equation and applications to aerosol growth simulation. Monte Carlo Methods and Applications, 2(1):41–87, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1996.2.
issue-1/mcma.1996.2.1.41/mcma.1996.2.1.41.xml.

Sabelfeld:1995:RWS

[SS95] K. K. Sabelfeld and I. A. Shalimova. Random walk on spheres process for exterior Dirichlet problem. Monte Carlo Methods and Applications, 1(4):325–331, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
4/mcma.1995.1.4.325/mcma.1995.1.4.325.xml.

Sahoo:1997:TUS

[SS97] L. N. Sahoo and J. Sahoo. On three unbiased strategies in sample surveys. Monte Carlo Methods and Applications, 3(1):73–81, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1997.3.
issue-1/mcma.1997.3.1.73/mcma.1997.3.1.73.xml.

Sabelfeld:2001:FBS

[SS01] K. Sabelfeld and I. Shalimova. Forward and backward stochastic Lagrangian models for turbulent transport and the well-mixed condition. Monte Carlo Methods and Applications, 7(3–4):369– 381, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-3-4/mcma.2001.7.3-4.369/mcma.2001.

7.3-4.369.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Sabelfeld:2002:RWS

[SS02] Karl K. Sabelfeld and Irina A. Shalimova. Random walk on spheres methods for iterative solution of elasticity problems. Monte Carlo Methods and Applications, 8(2):171–202, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
2/mcma.2002.8.2.171/mcma.2002.8.2.171.xml.

Sabelfeld:2003:FRW

[SS03] Karl Sabelfeld and Elena Shkarupa. Functional random walk on spheres algorithm for biharmonic equation: optimization and error estimation. Monte Carlo Methods and Applications, 9(1):51–65, January 2003. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2003.9.issue-1/156939603322587461/156939603322587461.
xml.

Sobol:2007:GSI

[SS07] Ilya M. Sobol′ and Boris V. Shukhman. On global sensitivity indices: Monte Carlo estimates affected by random errors. Monte Carlo Methods and Applications, 13(1):89–97, ???? 2007.

CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2007.
13.issue-1/mcma.2007.005/mcma.2007.005.xml.

Shalimova:2014:SPC

[SS14a] Irina A. Shalimova and Karl K. Sabelfeld. Stochastic polynomial chaos based algorithm for solving PDEs with random coefficients. Monte Carlo Methods and Applications, 20(4):279–??, December 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2014.
20.issue-4/mcma-2014-0006/mcma-2014-0006.xml.

Sobol:2014:QMC

[SS14b] Ilya M. Sobol and Boris V. Shukhman. Quasi-Monte Carlo: A high-dimensional experiment. Monte Carlo Methods and Applications, 20(3):167–171, September 2014. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2014.20.issue-3/mcma-2013-0022/
mcma-2013-0022.xml.

Shukhman:2015:LTA

[SS15] Boris V. Shukhman and Ilya M. Sobol. A limit theorem for average dimensions. Monte Carlo Methods and Applications, 21(2):175–??, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2015.21.issue-2/mcma-2014-0018/mcma-2014-0018.xml.

Shalimova:2017:SPC

[SS17] Irina A. Shalimova and Karl K. Sabelfeld. Stochastic polynomial chaos expansion method for random Darcy equation. Monte Carlo Methods and Applications, 23(2):101–??, June 2017. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2017.
23.issue-2/mcma-2017-0109/mcma-2017-0109.xml.

Shalimova:2018:RWS

[SS18a] Irina Shalimova and Karl K. Sabelfeld. Random walk on spheres method for solving anisotropic drift-diffusion problems. Monte Carlo Methods and Applications, 24(1):43–54, March 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-1/mcma-2018-0006/mcma-2018-0006.xml.

Sobol:2018:ADP

[SS18b] Ilya M. Sobol and Boris V. Shukhman. On average dimensions of particle transport estimators. Monte Carlo Methods and Applications, 24(2):147–151, June 2018. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2018.24.issue-2/
mcma-2018-0013/mcma-2018-0013.xml.

Shalimova:2019:RWS

[SS19a] Irina Shalimova and Karl K. Sabelfeld. A random walk on small spheres method for solving transient anisotropic diffusion problems. Monte Carlo Methods and Applications, 25(3):271–??, September 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-3/mcma-2019-2047/mcma-2019-2047.xml.

Sobol:2019:QMC

[SS19b] I. M. Sobol and B. V. Shukhman. Quasi-Monte Carlo method for solving Fredholm equations. Monte Carlo Methods and Applications, 25(3):253–??, September 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-3/
mcma-2019-2045/mcma-2019-2045.xml.

Shalimova:2020:RWE

[SS20a] Irina Shalimova and Karl K. Sabelfeld. Random walk on ellipsoids method for solving elliptic and parabolic equations. Monte Carlo Methods and Applications, 26(4):335–353, November 20, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2078/html.

Sobol:2020:QIE

[SS20b] Ilya M. Sobol and Boris V. Shukhman. QMC integration errors and quasi-asymptotics. Monte Carlo Methods and Applications, 26 (3):171–176, July 16, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-2067/html.

Sabelfeld:2021:GRWa

[SS21a] Karl K. Sabelfeld and Dmitrii Smirnov. A global random walk on grid algorithm for second order elliptic equations. Monte Carlo Methods and Applications, 27(3):211–225, August 8, 2021.

CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2092/html.

Shalimova:2021:RWS

[SS21b] Irina Shalimova and Karl K. Sabelfeld. Random walk on spheres algorithm for solving steady-state and transient diffusionrecombination problems. Monte Carlo Methods and Applications, 27(4):301–313, November 4, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2021-2099/html.

Sakouvogui:2021:SAS

[SSDM21] Kekoura Sakouvogui, Saleem Shaik, Curt Doetkott, and Rhonda Magel. Sensitivity analysis of stochastic frontier analysis models. Monte Carlo Methods and Applications, 27(1):71–90, February 2, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2021-2083/html.

Sabelfeld:2021:GRWb

[SSDT21] Karl K. Sabelfeld, Dmitry Smirnov, Ivan Dimov, and Venelin Todorov. A global random walk on grid algorithm for second order elliptic equations. Monte Carlo Methods and Applications, 27(4): 325–339, October 27, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-2097/html.

Sobol:1999:DRR

[SSG99] I. M. Sobol, B. V. Shukhman, and A. Guinzbourg. On the distribution of random ranges. Monte Carlo Methods and Applications, 5(2):113–134, ???? 1999. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1999.5.issue-2/mcma.1999.5.2.113/mcma.1999.
5.2.113.xml.

Sabelfeld:2004:DRW

[SSL04] K. K. Sabelfeld, I. A. Shalimova, and A. I. Levykin. Discrete random walk on large spherical grids generated by spherical means for PDEs. Monte Carlo Methods and Applications, 10(3–4): 559–574, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.559/
mcma.2004.10.3-4.559.xml.

Sabelfeld:2006:RWF

[SSL06] K. K. Sabelfeld, I. A. Shalimova, and A. I. Levykin. Random walk on fixed spheres for Laplace and Lamé equations. Monte Carlo Methods and Applications, 12(1):55–93, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-1/156939606776886634/156939606776886634.xml.

Sahoo:2006:ESA

[SSS06] L. N. Sahoo, R. K. Sahoo, and S. C. Senapati. An empirical study on the accuracy of ratio and regression estimators in the presence of measurement errors. Monte Carlo Methods and Applications, 12(5–6):495–501, ???? 2006. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2006.
12.issue-5/156939606779329026/156939606779329026.xml.

Sabelfeld:1995:IFB

[ST95] K. K. Sabelfeld and D. Talay. Integral formulation of the boundary value problems and the method of random walk on spheres. Monte Carlo Methods and Applications, 1(1):1–34, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1995.1.
issue-1/mcma.1995.1.1.1/mcma.1995.1.1.1.xml.

Sugita:2000:RWS

[ST00] Hiroshi Sugita and Satoshi Takanobu. Random Weyl sampling for robust numerical integration of complicated functions. Monte Carlo Methods and Applications, 6(1):27–48, ???? 2000. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
1/mcma.2000.6.1.27/mcma.2000.6.1.27.xml.

Starkov:1995:MCS

[Sta95] A. V. Starkov. Monte Carlo splitting importance sampling. Monte Carlo Methods and Applications, 1(3):241–250, ???? 1995. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1995.1.issue-
3/mcma.1995.1.3.241/mcma.1995.1.3.241.xml.

Stefanescu:2000:GUR

[Ste00] Stefan V. Stefanescu. Generating uniform random points inside a cone. Monte Carlo Methods and Applications, 6(2):115–130, ????

2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2000.6.
issue-2/mcma.2000.6.2.115/mcma.2000.6.2.115.xml.

Sugita:1995:PRN

[Sug95] Hiroshi Sugita. Pseudo-random number generator by means of irrational rotation. Monte Carlo Methods and Applications, 1(1):35–57, ???? 1995. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
1995.1.issue-1/mcma.1995.1.1.35/mcma.1995.1.1.35.xml.

Sugita:2004:SPR

[Sug04] Hiroshi Sugita. Security of pseudo-random generator and Monte Carlo method. Monte Carlo Methods and Applications, 10(3–4): 609–615, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.

com/view/j/mcma.2004.10.issue-3-4/mcma.2004.10.3-4.609/
mcma.2004.10.3-4.609.xml.

Schell:2004:MUD

[SUZ04] Thomas Schell, Andreas Uhl, and Peter Zinterhof. Measures of uniform distribution in wavelet based image compression. Monte Carlo Methods and Applications, 10(3–4):587–597, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.587/mcma.2004.10.3-4.587.
xml.

Suciu:2004:NML

[SVH^{+}04] N. Suciu, C. Vamoş, H. Hardelauf, J. Vanderborght, and H. Vereecken. Numerical modeling of large scale transport of contaminant solutes using the global random walk algorithm. Monte Carlo Methods and Applications, 10(2):153–177, ???? 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-2/156939604777303235/156939604777303235.xml.

Svit:2021:SMS

[SZKS21] Kirill Svit, Konstantin Zhuravlev, Sergey Kireev, and Karl K. Sabelfeld. A stochastic model, simulation, and application to aggregation of cadmium sulfide nanocrystals upon evaporation of the Langmuir–Blodgett matrix. Monte Carlo Methods and Applications, 27(4):289–299, November 6, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.

degruyter.com/document/doi/10.1515/mcma-2021-2100/html.

Takashima:1996:HWT

[Tak96a] Keizo Takashima. On Hamming weight test and sojourn time test of m-sequences. Monte Carlo Methods and Applications, 2 (4):331–340, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1996.2.issue-4/mcma.1996.2.4.331/mcma.1996.
2.4.331.xml.

Takashima:1996:NMC

[Tak96b] Keizo Takashima. On the number of multiples of certain primitive polynomials over GF(2). Monte Carlo Methods and Applications, 2(1):15–24, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1996.2.issue-1/mcma.1996.2.1.
15/mcma.1996.2.1.15.xml.

Takashima:1997:RWT

[Tak97] Keizo Takashima. Random walk tests of reciprocal m-sequences. Monte Carlo Methods and Applications, 3(2):155–166, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1997.3.
issue-2/mcma.1997.3.2.155/mcma.1997.3.2.155.xml.

Takashima:2000:HPR

[Tak00] K. Takashima. Hybrid pseudo-random number generation. Monte Carlo Methods and Applications, 6(1):49–59, ???? 2000. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
1/mcma.2000.6.1.49/mcma.2000.6.1.49.xml.

Thrasher:2020:ESR

[TM20] W. John Thrasher and Michael Mascagni. Examining sharp restart in a Monte Carlo method for the linearized Poisson–Boltzmann equation. Monte Carlo Methods and Applications, 26(3):223– 244, August 11, 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2020-2069/html.

Torres:2020:DPD

[Tor20] David J. Torres. Describing the Pearson R distribution of aggregate data. Monte Carlo Methods and Applications, 26(1):17–??, March 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2020.
26.issue-1/mcma-2020-2054/mcma-2020-2054.xml.

Tamiti:2018:UVR

[TOTAI18] Kenza Tamiti, Megdouda Ourbih-Tari, Abdelouhab Aloui, and Khelidja Idjis. The use of variance reduction, relative error and bias in testing the performance of M/G/1 retrial queues estimators in Monte Carlo simulation. Monte Carlo Methods and Applications, 24(3):165–178, September 2018. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2018.
24.issue-3/mcma-2018-0015/mcma-2018-0015.xml.

Tinet:2001:RTT

[TTEA01] E. Tinet, J. M. Tualle, D. Ettori, and S. Avrillier. Real time transformation of pre-computed Monte Carlo results for fitting optical measurements in biomedical applications. Monte Carlo Methods and Applications, 7(3–4):397–409, ???? 2001. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2001.7.issue-
3-4/mcma.2001.7.3-4.397/mcma.2001.7.3-4.397.xml.

Tuffin:1996:ULD

[Tuf96] Bruno Tuffin. On the use of low discrepancy sequences in Monte Carlo methods. Monte Carlo Methods and Applications, 2(4): 295–320, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.1996.2.issue-4/mcma.1996.2.4.295/mcma.1996.

2.4.295.xml. See comments [Tuf98].

Tuffin:1998:CUL

[Tuf98] Bruno Tuffin. Comments on: “On the use of low discrepancy sequences in Monte Carlo methods” [Monte Carlo Methods Appl. 2 (1996), no. 4, 295–320; MR1434423 (97m:65018)]. Monte Carlo Methods and Applications, 4(1):87–90, ???? 1998. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1998.4.issue-
1/mcma.1998.4.1.87/mcma.1998.4.1.87.xml. See [Tuf96].

Tuffin:2004:RQM

[Tuf04] Bruno Tuffin. Randomization of quasi-Monte Carlo methods for error estimation: Survey and normal approximation. Monte Carlo Methods and Applications, 10(3–4):617–628, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2004.
10.issue-3-4/mcma.2004.10.3-4.617/mcma.2004.10.3-4.617.
xml.

Turchyn:2011:SSG

[Tur11] Yevgen V. Turchyn. Simulation of sub-Gaussian processes using wavelets. Monte Carlo Methods and Applications, 17(3):215–231, September 2011. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2011.17.issue-3/mcma.2011.010/mcma.2011.010.xml.

Turchyn:2019:WBS

[Tur19] Ievgen Turchyn. Wavelet-based simulation of random processes from certain classes with given accuracy and reliability. Monte Carlo Methods and Applications, 25(3):217–??, September 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2019.
25.issue-3/mcma-2019-2042/mcma-2019-2042.xml.

Ugrin-Sparac:1996:NAG

[UŠ96] Dimitrije Ugrin-Šparac. A natural algorithm for generation of pseudo-random numbers and its applications. Monte Carlo Methods and Applications, 2(3):191–217, ???? 1996. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.1996.2.issue-
3/mcma.1996.2.3.191/mcma.1996.2.3.191.xml.

Uhinov:2000:UIS

[UV00] S. A. Uhinov and A. V. Voytishek. Usage of the importance sample in Monte Carlo methods. Monte Carlo Methods and Applications, 6(4):341–348, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
4/mcma.2000.6.4.341/mcma.2000.6.4.341.xml.

Vasileska:2004:MCS

[VA04] Dragica Vasileska and Shaikh S. Ahmed. Monte Carlo simulation of narrow-width SOI devices: Incorporation of the short range Coulomb interaction. Monte Carlo Methods and Applications, 10(3–4):629–640, December 2004. CODEN MCMAC6.

ISSN 0929-9629 (print), 1569-3961 (electronic). URL http:

//www.degruyter.com/view/j/mcma.2004.10.issue-3-4/mcma.
2004.10.3-4.629/mcma.2004.10.3-4.629.xml.

Vanslette:2020:WSQ

[VAYT20] Kevin Vanslette, Abdullatif Al Alsheikh, and Kamal Youcef-Toumi. Why simple quadrature is just as good as Monte Carlo. Monte Carlo Methods and Applications, 26(1):1–??, March 2020. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2020.
26.issue-1/mcma-2020-2055/mcma-2020-2055.xml.

Voytishek:2000:GMC

[VDM00] A. V. Voytishek, E. G. Dyatlova, and T. E. Mezentseva. Geometrical Monte Carlo method and its modifications. Monte Carlo Methods and Applications, 6(2):131–139, ???? 2000. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
2/mcma.2000.6.2.131/mcma.2000.6.2.131.xml.

Vidya:2007:PRA

[Vid07] K. P. Vidya. Pollard’s rho attack on ECDLP and threshold schemes. Monte Carlo Methods and Applications, 13(3):245– 252, August 2007. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2007.13.issue-3/mcma.2007.013/mcma.2007.013.xml.

Voytishek:2008:UAN

[VMS08] Anton Voytishek, Alexandr Myasnikov, and Leonid Saneev. A use of algorithms for numerical modeling of order statistics. Monte Carlo Methods and Applications, 13(5–6):467–483, January 2008. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2008.
13.issue-5-6/mcma.2007.024/mcma.2007.024.xml.

Voytishek:1997:USF

[Voy97] A. V. Voytishek. Using the Strang–Fix approximation in discretestochastic numerical procedures. Monte Carlo Methods and Applications, 3(2):89–112, ???? 1997. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.1997.3.issue-2/mcma.1997.3.2.
89/mcma.1997.3.2.89.xml.

Voytishek:1998:RMM

[Voy98] A. V. Voytishek. Rejection methods for modelling of betadistribution. Monte Carlo Methods and Applications, 4(1):73–85, ???? 1998. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
1998.4.issue-1/mcma.1998.4.1.73/mcma.1998.4.1.73.xml.

Wagner:2008:DPM

[Wag08] Wolfgang Wagner. Deviational particle Monte Carlo for the Boltzmann equation. Monte Carlo Methods and Applications, 14(3):191– 268, September 2008. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2008.14.issue-3/mcma.2008.010/mcma.2008.010.xml.

Wagner:2010:RDF

[Wag10] Wolfgang Wagner. Random and deterministic fragmentation models. Monte Carlo Methods and Applications, 16(3–4):399–420, December 2010. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/
mcma.2010.16.issue-3-4/mcma.2010.016/mcma.2010.016.xml.

Wagner:2015:CPM

[Wag15] Wolfgang Wagner. A class of probabilistic models for the Schrödinger equation. Monte Carlo Methods and Applications, 21(2):121–??, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.degruyter.com/

view/j/mcma.2015.21.issue-2/mcma-2014-0014/mcma-2014-0014.
xml.

Warin:2018:NMC

[War18] Xavier Warin. Nesting Monte Carlo for high-dimensional non-linear PDEs. Monte Carlo Methods and Applications, 24(4):225–247, December 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2018.24.issue-4/mcma-2018-2020/mcma-2018-2020.xml.

Wells:2006:SAS

[Wel06] C. G. Wells. A stochastic approximation scheme and convergence theorem for particle interactions with perfectly reflecting boundary conditions. Monte Carlo Methods and Applications, 12(3–4):291– 342, ???? 2006. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/j/

mcma.2006.12.issue-3/156939606778705182/156939606778705182.
xml.

Wang:2009:BNB

[WENG09] Yi Wang, Kent M. Eskridge, S. Nadarajah, and Andrzey T. Galecki. Bayesian and non-Bayesian analysis of mixed-effects PK/PD models based on differential equations. Monte Carlo Methods and Applications, 15(2):145–167, August 2009. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2009.
15.issue-2/mcma.2009.009/mcma.2009.009.xml.

Wihstutz:2001:CSD

[Wih01] Volker Wihstutz. Communication structure of discretized degenerate diffusion processes and approximation of Lyapunov exponents. Monte Carlo Methods and Applications, 7(3–4):411– 419, ???? 2001. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2001.7.issue-3-4/mcma.2001.7.3-4.411/mcma.2001.

7.3-4.411.xml. Monte Carlo and probabilistic methods for partial differential equations, Part II (Monte Carlo, 2000).

Wells:2005:DSM

[WK05] Clive G. Wells and Markus Kraft. Direct simulation and mass flow stochastic algorithms to solve a sintering–coagulation equation. Monte Carlo Methods and Applications, 11(2):175–197, ???? 2005. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2005.
11.issue-2/156939605777585980/156939605777585980.xml.

Wiart:2021:DSQ

[WLD21] Jaspar Wiart, Christiane Lemieux, and Gracia Y. Dong. On the dependence structure and quality of scrambled (t, m, s)-nets. Monte Carlo Methods and Applications, 27(1):1–26, January 10, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2020-2079/html.

Weiss:2019:AOR

[WN19] Christian Weiß and Zoran Nikolić. An aspect of optimal regression design for LSMC. Monte Carlo Methods and Applications, 25 (4):283–??, December 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-4/mcma-2019-2049/mcma-2019-
2049.xml.

Xiao:1996:VPF

[Xia96] Yi-Jun Xiao. Variation of product function and numerical solution of some partial differential equations by low-discrepancy sequences. Monte Carlo Methods and Applications, 2(4):321–330, ???? 1996. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.1996.2.
issue-4/mcma.1996.2.4.321/mcma.1996.2.4.321.xml.

Xiao:2002:SGP

[Xia02] Yi-Jun Xiao. Some geometric properties of (0, m, 2)-nets in base b ≥ 2. Monte Carlo Methods and Applications, 8(1):97–106, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2002.8.
issue-1/mcma.2002.8.1.97/mcma.2002.8.1.97.xml.

Yaguchi:2000:RHR

[Yag00] Hirotake Yaguchi. Randomness of Horner’s rule and a new method of generating random numbers. Monte Carlo Methods and Applications, 6(1):61–76, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://www.

degruyter.com/view/j/mcma.2000.6.issue-1/mcma.2000.6.1.
61/mcma.2000.6.1.61.xml.

Yaguchi:2002:CLP

[Yag02] Hirotake Yaguchi. Construction of a long-period nonalgebraic and nonrecursive pseudorandom number generator. Monte Carlo Methods and Applications, 8(2):203–213, ???? 2002. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2002.8.issue-
2/mcma.2002.8.2.203/mcma.2002.8.2.203.xml.

Yamada:2021:HOW

[Yam21] Toshihiro Yamada. High order weak approximation for irregular functionals of time-inhomogeneous SDEs. Monte Carlo Methods and Applications, 27(2):117–136, February 20, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/mcma-2021-2085/html.

Yang:2013:NNS

[Yan13] Xuewei Yang. A new numerical scheme for a class of reflected stochastic differential equations. Monte Carlo Methods and Applications, 19(4):273–??, December 2013. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2013.
19.issue-4/mcma-2013-0011/mcma-2013-0011.xml.

Yu:2021:DMC

[YJH21] Unjong Yu, Hoseung Jang, and Chi-Ok Hwang. A diffusion Monte Carlo method for charge density on a conducting surface at nonconstant potentials. Monte Carlo Methods and Applications, 27(4): 315–324, October 28, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/mcma-2021-2098/html.

Yaguchi:2008:NNP

[YK08] Hirotake Yaguchi and Izumi Kubo. A new nonrecursive pseudorandom number generator based on chaos mappings. Monte Carlo Methods and Applications, 14(1):85–98, May 2008. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://hdl.handle.net/10076/9251;   http://
www.degruyter.com/view/j/mcma.2008.14.issue-1/mcma.2008.
005/mcma.2008.005.xml.

Yamada:2018:SOW

[YY18] Toshihiro Yamada and Kenta Yamamoto. A second-order weak approximation of SDEs using a Markov chain without Lévy area simulation. Monte Carlo Methods and Applications, 24(4):289–308, December 2018. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL https://www.degruyter.com/view/j/
mcma.2018.24.issue-4/mcma-2018-2024/mcma-2018-2024.xml.

Zalesky:2000:SRB

[Zal00] B. A. Zalesky. Stochastic relaxation for building some classes of piecewise linear regression functions. Monte Carlo Methods and Applications, 6(2):141–157, ???? 2000. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic).

URL http://www.degruyter.com/view/j/mcma.2000.6.issue-
2/mcma.2000.6.2.141/mcma.2000.6.2.141.xml.

Zarezadeh:2019:PDM

[ZC19] Zakarya Zarezadeh and Giovanni Costantini. Particle diffusion Monte Carlo (PDMC). Monte Carlo Methods and Applications, 25(2):121–??, June 2019. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL https://www.degruyter.

com/view/j/mcma.2019.25.issue-2/mcma-2019-2037/mcma-2019-
2037.xml.

Zio:2004:DSV

[ZCC04] E. Zio, A. Cammi, and A. Cioncolini. Dagger-sampling variance reduction in Monte Carlo reliability analysis. Monte Carlo Methods and Applications, 10(3–4):641–652, December 2004. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

http://www.degruyter.com/view/j/mcma.2004.10.issue-3-4/
mcma.2004.10.3-4.641/mcma.2004.10.3-4.641.xml.

Zherelo:2013:CMB

[Zhe13] Anatoly Zherelo. On convergence of the method based on approximately exact formulas for functional polynomials for calculation of expectations of functionals to solutions of stochastic differential equations. Monte Carlo Methods and Applications, 19(3):183–??, September 2013. CODEN MCMAC6. ISSN 0929-9629 (print),

1569-3961 (electronic). URL http://www.degruyter.com/view/
j/mcma.2013.19.issue-3/mcma-2013-0009/mcma-2013-0009.xml.

Zalewska:2010:MIA

[ZNS10] Marta Zalewska, Wojciech Niemiro, and Bolesław Samoliński. MCMC imputation in autologistic model. Monte Carlo Methods and Applications, 16(3–4):421–438, December 2010. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL http://www.degruyter.com/view/j/mcma.2010.
16.issue-3-4/mcma.2010.017/mcma.2010.017.xml.

Zapadinsky:2002:ECC

[ZPK02] Evgeni Zapadinsky, Liisa Pirjola, and Markku Kulmala. Effect of cross-correlated fluctuations on the aerosol dynamics: Monte Carlo simulations. Monte Carlo Methods and Applications, 8(4):405–419, ???? 2002. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2002.8.issue-4/mcma.2002.8.4.405/mcma.2002.8.4.405.xml.

Zhang:2019:PMM

[ZYD19] Lihao Zhang, Zeyang Ye, and Yuefan Deng. Parallel MCMC methods for global optimization. Monte Carlo Methods and Applications, 25(3):227–??, September 2019. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2019.25.issue-3/
mcma-2019-2043/mcma-2019-2043.xml.

Ziane:2021:BTA

[ZZA21] Yasmina Ziane, Nabil Zougab, and Smail Adjabi. Body tail adaptive kernel density estimation for nonnegative heavy-tailed data. Monte Carlo Methods and Applications, 27(1):57–69, February 2, 2021. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/mcma-2021-2082/html.